A Bibliography of Publications in Computer Graphics World

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

21 June 2011 Version 1.18

Title word cross-reference

1- [Tuc96]. 1.1 [Ano98-48]. 16-bit [Ano95-28].

2.0 [Gra96]. 21-inch [Ano95-29]. 2D [Coc95f, Mac90a]. 2nd [Ano96n].

3.0 [Ano97-40]. 3.1 [Ano98-49]. 3.5 [Ano97-44]. 3000 [Ano91-32]. 3D [Ano95-48, Coc95b, Ano90f, Rob90f]. 3D. [Ano98q]. 3D/Eye [Ano95-38].

4.1 [Ano97-47]. 4.5 [Ano98-45]. 4D [Ano97-37].

5 [Ano97-57]. 5.0 [Ano96-53].

76 [Ano91b].

’90s [Rob93g]. ’92 [Ano92-38]. ’95 [Doy95]. ’96 [Ano95-60]. 97 [Ano97-45]. 97/ FastSOLID [Ano97-37]. 98 [Ano98-47].

A-h-h-h-h [Rob98h]. A/V [Har96a]. ABC [Spr94]. ABCs [Emm91c]. ability [Rob94a]. able [Coh96]. Accelerating [Sch95a, Sch95b]. acceleration [Rob96d]. accelerator [Cla96]. accelerators [Coc95a, Ano95-34]. acceptance [Rob94a]. accommodate [Pot97a]. According [Mah95c]. achieve [Emm94]. Achieving [Pot96a]. ACIS [Por95b]. Across [Sor90a]. Act [Rob94a]. Action [Mor93]. action-adventure [Rob96c]. activity [RR97]. actual [Phi96e]. ADAMS [Ano95-52]. ADAMS/ [Ano95-52]. Adding [Ano98q]. addresses [Ano96-69]. Adds [Ano90n]. adept [Pot96d]. adhered [Phi95e]. Adobe [Ano96-50]. advance [Phi97e]. advanced [Ano95-52]. Advancements [Sor91a]. advances [Oli94]. advantage [Cra94]. advantages [Gar95]. adventure [Rob95b, Gra93a]. Adventures [Nov96]. AEC [Rob90d, Ano91e]. aerospace [Pot96d]. Affordable [Mae97]. after [Cla96, Ano96-50]. AG [Sha95]. again [Rob96c]. Age [Fly92]. ago [Pot96e]. Ahead [Ros90]. aided [Tho95b, Sen90a]. aids [Ano95-58]. aircraft [Ano96-81]. airplane [Ano96s]. Airport [Ano98-29]. Airwaves [Sor90a]. Albert [Ano96-79]. Algorithmic [Sor91a]. algorithms [Kin97]. Alias [Ano92-37, Ano95l]. Alias/ Wavefront [Ano98-45]. alien [Ano96p]. Alive [Mac90a]. all-digital [Coc95e]. Alliant [Ano90b]. Allplan [Ano97-45]. Allure [Pot95f]. alone [Emm96a]. already [Pot96e]. also [Mah95c]. Although [Coh96]. Altschul [Mah95c]. always [Rob97b]. Amazing [Rob92f, Ano96k]. Amblin [Ano96-27]. America [Ano95-56]. American [Ano96-81]. Americans [Har96c]. Amiga [Ano91-32, Ano90c]. AmigaVision [Ano90c]. analyses [Ano95-57, Ano95-61]. Analysis [Pot92a]. analysts [Pot97a]. analyze [Coh96]. Analyzing [Ano95-63]. Anastasia [Ano97w]. Anatomy [Sim95b, Sim95a]. Ancient [Pfi92a]. Angeles [Doy95]. ani [Rob95c]. Animals [Rob96c]. animated [Emm95a]. AnimaTek [Ano95-43]. animating [Phi96c]. Animation [Mac90a, Vas90b]. Animation/ Photo [Ano90f]. Animation/ Photo-Realism [Ano90f]. animations [Ano95-60, Phi95e]. animator [Mah95c, Ano96-28]. animators [Cra94]. Anime [Sav98]. announced [Pot96e]. Annual [Ano96n]. answer [Mae96]. Antz [Rob98a, Rob98f]. Antz-piration [Rob98a]. anyone [Pot97a]. appeal [Phi96g]. appearing [For97]. Apple [Ano90e]. application [Emm96c]. Applications [Ano92d]. Approach [Kee86]. approaches [Emm94]. arcade [Ano96j]. ArchiDesign [Ano98-46]. architect [Row96]. architected [Pot96c]. Architects [Mah94c, Phi95a]. Architectural [Mah94f, Pot95b]. Architecture [Nov96, Phi95a]. Architectures [Pot96c]. archive [Ano96-69]. area [Phi96c]. aren’t [Coc97d]. argue [Smi96a]. Arizona [Ano96-75]. array [Emm95a]. arrival [Coc97d]. Ars [Ano98-43]. Art [Ano90h]. artful [Rob94a]. Artificial [Mah93a]. Artisan [Ano97-39]. artist [Ano96-36]. Artistic [Coc98c, LB90a]. artists [Doy94, LB91b]. Arts [Pfi93b]. Assemblies [Pot96d, Coo90a]. assumptions [Coh96]. Athlete [Ano98-67]. Atlanta [Vas90a]. attempt [Mae96]. attitude [Ano98-68]. Attraction [Ano98-27]. Author [SH90]. auto.des.sys [Ano95-45]. Autocad [Ano91h]. Autodesk [Ano90f]. Autofact [Ano95-35]. automaker [Ano96-49]. automotive [Pot96d]. available [Vac96, Doy94]. avalanches [Phi96a]. Avant [Mah94c]. Avant-Garde [Mah94c]. average [Phi97a]. Averting [Ano98-61]. Aviation [Ano98-61]. Avid [Ano96-50]. avionics [Phi96d]. Awakening [Nov98a]. award [Rob94f, Ano96-31]. award-winning [Rob94f].

Awards [Ano95w, Ano96n].

babies [Ano98y]. Babylon [Ano97-57]. Back [BZJ94]. Bag [Rob94b]. Balto [Ano96-27]. Bananas [Ano98l]. Barbie [Ano98s]. Bargain [Pot96g]. Barrier [Han90]. Barriers [Fic86]. Bart [Ano95-59]. based [Ano92b, Ano90g]. basic [Ash96]. battle [Rob96a]. Battling [Rob96a]. Battlinq [Pot98a]. be [Tuc95a]. beating [Pot97b]. Beauty [Ano98-42]. Because [Pot96a]. become [Rob94d]. becomes [Pot95f]. becoming [Vac96]. been [Cra94]. before [Doy94]. begin [Hel97, Mah92c]. Beginner [Woo93]. beginning [Pot96a]. begins [Pot95f]. behind [Spr94]. behind-the-scenes [Spr94]. Beholder [FK90]. BeInfinite [Ano95-46]. being [Mae96]. beings [Rob94a]. believable [Ano95m]. Believe [Emm86]. Bell [Ano95-52]. benchmark [Kur97]. Benchmarks [Coh96]. benefits [Ano95r]. Bent [Gra93a]. Bentley [Ano95-39]. Best [Pot95d]. best-suited [Coc95f]. better [Phi95f]. Beyond [Pop94]. bid [Ano95-60]. Big [Coc95e, Dav92]. big-screen [Coc95e]. Big-Time [Dav92]. bike [Ano98-66]. Bin [Pot96g]. Bingo [Ano98c]. biological [Ano97-56]. bit [Ano95-28]. Biz [Rob91c]. Blitz [Ano96-55]. Blue [Ano98-60]. blueprint [Pot96f]. Blues [Ano91-28]. Board [Rob93a]. boards [Cla96]. boasting [Coc96a]. bold [Ano96-71, LB90b]. Bolsters [McM90b]. Bolt [Gra96]. Bones [Ano98-62]. Boolean [Ano96w]. Boolean-shy [Ano96w]. booming [Mah97]. Boss [Rob95b]. both [Coc95a]. Boundaries [KPV91]. brain [Ano97-50]. Brainstorm [Mah98b]. Breaking [Han90]. breathing [Rob96c, Coc95c]. Breed [FP91]. bridge [Ano97-51, Ano95-63]. Bridging [Rob90a]. brighten [Ano95-63]. bring [Sch95b]. bringing [Coc96b]. brings [Coc95e]. Broad [Ano95a]. broadcast [Rob95h]. Broadening [KPV91]. broader [Tuc96]. Bros. [Rob95a]. Brushing [Rob98b]. Bryce [Ano96-54]. Bucci [Ano96-36]. Bug [Ano98d]. buggy [Ano98d]. Builder [Ano97-47, Ano95-43]. Building [Lon98, Min94]. built [Ano97-64]. bureaus [Vin97]. burgeoning [Phi95a]. business [Ano96b, Nov93b]. butterflies [Ano96x]. buy [Tuc95a]. buyers [Ano96i]. Buying [Vin97].

CAD [Ros90, Fre90a]. CAD/ CAM [Ano95-47, Ano95f]. Cadam [Ano90g]. Cadkey [Ano97-37]. CAD’s [Pot96f]. CAE [Ano98g]. CAID [Ano95g]. Caligan [Ano95-40]. Caligari [Ano97-40]. camera [Ano98-60]. cameras [Rob95h, Ano95-31]. campaign [Ano96-64]. Campus [Mah92a]. can [Doy94]. Canadian [Ano95-55]. candidates [Ano96-64]. capabilities [Cra94]. CAPS [Rob94b]. capture [Rob94a]. captured [Oli94]. Capturing [Mae95b, Mae95c]. car [App93]. car-collision [Ano95-61]. cards [Ano96-43, Ano95-28]. Care [McM90b, Mah92b]. Career [Emm95c]. careful [Emm96c]. Carnegie [Ano97-54]. Carpentry [Pfi92b]. cartoon [Ano95-63, Vas90b]. Casino [Ano96-74]. Casper [Rob95e]. Cast [Rob98f]. Catalyst [Ano98-48]. Caught [Rob94a]. Cave [Ano95-54]. CD [Doy94]. CD-ROM [Ano95-63]. CD-ROM-based [Spr94]. CD-ROMs [Doy94]. CDI [Ano90h]. cell [Ano97t]. center [Ano95-62]. centralized [For97]. ceramic [Ano95-54]. cereal [Ano95-63]. CFD [Ano95-56]. CG [Rob95h]. CG-enhanced [Ano97-50]. challenge [Doy95, LB90a]. challenges [Ano95e]. chance [Rob94f]. chances [Row96]. change [Rob97b, Por91c]. changes [Hir96, Dav92]. Changing [Rob90d]. Chaos [Mah93d]. Character [Ano96-60, Ano96p]. character-animation [Ano98b]. Characters [Rob94g, Sim95a]. Chase [Ano98-70]. cheaper [Pot97c]. Check [Dav91]. Chemists [Ano97-54]. children [Rob94d, Ano95-53]. China [Ano96-73]. chip [Ano96j, Sto90]. chips [Ano96j]. Choice [Yar95]. choices [Tuc96]. choose [Mac95]. Choosing [Emm96c]. chops [Pot97b]. Chrysler [Ano97-58]. Cinema [Ano97-37]. circles [Pot95f]. Citadines [Ano96-49]. cited [Sim95a]. Cities [Nov93c]. City [Phi96e]. Cityscapes [Mah94d]. claim [Smi96b]. claiming [Rob96a]. clever [Rob97b]. clinch [Ano95-60]. Clones [Ano98-62]. clothing [Phi96c]. clues [Kur97]. collaborates [Ano98d]. collaboration [Rom97]. collaborative [Pot95c]. collision [Ano95-61]. Color [McM90a, Ano91j]. Colorful [Hol93]. colors [Phi95f, Ano91-33]. combine [Doy94]. combines [Rob96b]. come [Coc95e]. comeback [Pot97b]. comedy [Rob94f]. Comes [Lan92a]. comfort [Phi96d]. coming [Cla96]. commercial [Cra94]. common [Pot96c]. Common-sense [Ano97b]. Communication [Nov93a]. communications [Pot95a]. community [Pot95f]. companies [Pot96b, Gar95]. company [Ano96-74]. compare [Kur97]. Comparison [Kur97]. Competition [Ano95t, Ano95w]. competitors [Pot97b]. complex [Phi95f]. complexities [Ano98-60]. Composer [Ano98-45]. composite [Ano96q]. compositing [Emm95a, Coc95d]. Compression [Coo90b]. Computer [Vas90d]. computer-aided [Tho95b, Sen90a]. computer-generated [Phi96f]. computing [Tuc95a, Mah92a]. Concept [Pot97c]. concepts [App93]. conductive [Ano95-51]. conference [Doy95]. conjunction [Phi96b]. conjuring [Emm95a]. connection [Ano95-51]. conquers [Coc95b]. Consciousness [LS92]. consider [Yar95]. constraints [Sim95a]. Constructing [Ano97l]. Construction [Phi98]. Constructive [Nov93a]. consuming [Nov96]. containing [Pot96d]. Contender [Ano91m]. contest [Ano96-30]. continue [Pot97d]. continues [Row96]. Contracts [Pot96b]. contribute [Ano95-56]. control [Phi96d, Doy91a]. Convention [Ano95-60]. conventional [Pot97c]. converters [Ano96-40]. convincing [Ano96q]. cool [Ano95k]. Coping [Alo94]. Copy [Fly92]. Copyright [Alo94]. Core [Pot93a]. Corel [Ano95-42]. Corp [Ano95-42]. corporate [Har96d]. correctly [Coh96]. Cosmos [Sor91b]. cost [Gar95, Ano95-48]. cost-effective [Coc95a]. costs [Sim95a]. could [Ano97-64]. Country [Hir98]. Coverage [Ano98-67]. covert [Woh97]. Craft [Ano96x]. crash [Ano98-68]. create [Spr94]. created [Doy94]. creates [Ano96-75]. Creating [Ano95-63]. creativity [Ano96-30, Ano92i]. creators [Phi95e]. creature [Rob95b]. creatures [Rob96c]. Creeps [Coc98b]. critical [Mac95]. criticized [Pot97b]. critter [Ano98b]. cross [Ano97n]. cross-cultural [Ano97n]. crowd [Ano97p]. CRT [Ano96-38]. Cruise [Doy91a]. cube [Ano97-54]. cultural [Ano97n]. Cup [Ano98-67, Ano95-56]. custom [Rob95c]. Cutting [Rob94e]. Cyber [Ano98-67].

D [Ano95-38]. Dancers [Ano98-72]. Darkness [Ano96k]. DarkStar [Ano96-81]. DARPA [Ano98j]. data [Ano95a, Emi91]. data-exchange [Ano95a]. database [Ano95q]. databases [Ano96-64]. DataCAD [Ano98-46]. Datavision [Ano95-41]. Dawson [Ano96-35]. Deal [McC96]. Dealing [Ano96r]. DEAS [Ano97-39]. DeBabelizer [Ano97-38]. debate [Pot97d]. debuts [Ano96-67]. decide [Pyr96]. deciding [Tuc95a]. DECstation [Ano90i]. Deep [Ano98-69]. defenses [Ano97-56]. Defining [Ano96a]. degree [Coc97d]. deliver [Emm96a]. delivering [Phi97e]. demand [Pot95a, CP92]. demands [Pot96d]. Departments [Ano98o]. depends [Sha95]. Design [Ano90h, Pfi90b]. design-to-manufacturing [Tho95b]. DesignBase [Pot97d]. designed [Rob94d]. designer [RR97, Lan90, Ano95k]. designers [Pot96a]. Designing [App93, Pfi90a]. designs [Ano96-33, Emm92b]. Desire [LeW93]. Desktop [Ano90m, McM90a]. Despite [Row96]. details [Ano96r]. determine [Hir96]. Detroit [Ano95-60]. develop [Phi96g]. Developer [Coc94a]. developers [Sha95]. developing [Phi95e, Coc94b]. development [Ano95-45, Ano92m]. Developments [Tho95b]. device [Yar95]. devices [Emm96a]. devising [Pot96c]. Dial [Gra94]. Diego [Ano95-60]. differ [Pot96c]. different [Coc95f]. differs [Rob96d]. difficult [Pot95d]. Digimation [Ano95-42]. digital [Ano92-30, Ano90i]. digital-video [Coc96a]. digitizer [Emm96c]. digitizers [Woh97, Woh92]. Digitizing [Ano95-47, Ano95e]. Dilbert [Ano97l]. Dimension [Ano91l]. dinnerware [Ano95-54]. dinosaur [Ano95n, Rob93b]. dinosaurs [Rob96c]. Direct [Ano96-39, Rob95b]. Direct-imaging [Ano96-39]. direction [Sha95]. directly [Rob94c]. director [Rob95b]. Dirty [War98a]. disappear [Pot96a]. Disasters [Ano98-61]. Disc [Rob92b]. Disc-O-Mania [Rob92b]. Disclosure [Rob95a, Rob95a]. Discovery [Mah93b]. disk [Ano97-32]. Disney [Rob94b]. Dispatching [Sen90a]. displacement [Ano96o]. display [Ano96-68]. Displaying [Ano97c]. Displays [Hel97, Win90]. Dissection [Ano98p]. distort [Coc96a]. diverse [Coc96a]. Do [Emm86, Sen90b]. Do-it-Yourself [Sen90b]. Documents [Bea91]. does [Smi96b]. doesn’t [Coc97d]. Dominance [Rob96a]. dominate [Yar95]. done [Rob96c]. don’t [Vin97, Gra94]. Doom [Ano96l]. Doors [Nov95]. down [Fic86]. downtown [Ano95-59]. Drafting [Ano98q]. dramatic [Hir96]. drawn [Ano96-27]. Dream [Ano97-36, Ano95u]. Drive [Har96a]. drives [Ano97-32, Sch95c]. Driving [Mah95b, Mah94e]. drops [Pot95f]. Drug [Mah93b]. Due [Phi96c]. DVD [Gla98]. dye [Gar95]. Dye-sub [Emm96a]. dye-sublimation [Gar95]. Dynamic [Ano95-60]. Dynamism [Emm92a].

each [Rob95h]. Earth [Ano96-69]. ease [Phi96d]. eases [Ano96-64]. easier [Smi96a]. easy [Phi96a, Pfi90a].

Edge [Pot94, Rob94e]. edit [Oli94]. Editing [Mac92]. editor [Ano95-42]. Educating [Por91b]. education [Ano96-70, McM90b]. edutainment [Ano97n]. Effect [Ano97-43]. effective [Coc95a, Mac93b]. effectively [Sch97e]. Effects [Sor91d, Coc95d]. Efficiency [Haa94]. effort [Emm94]. Efforts [Boy98a]. either [Phi96f]. ElectricImage [Ano96-59]. Electrify [Rob95d]. Electro [Pot93b]. Electro-Mechanical [Pot93b]. Electronic [Rob91e]. Electronica [Ano98-43]. element [Pot96a]. elements [Rob96b]. elevates [Ano96-78]. Elysium [Ano95-45]. Embrace [Coc97d]. embraces [Ano97-58]. Emergence [Rob90g]. emergency [Ano96-65]. emotional [Ano98-40]. Emphasis [Ano91f]. Emphasizing [Ano95b]. employ [Kin97]. enable [Kin97]. enabled [Rob95b]. enables [Rob94c]. encompasses [Ash96]. encountered [Fan96]. End [Fic86, Mac90b]. engineer [Row96]. Engineered [Row97c]. Engineering [Emi91, For92]. Engineers [Pot95a, Por91b]. Engines [Pot97d]. enhance [Pot96a]. enhanced [Ano95-52]. enhancements [Phi96g]. Enhances [Ano90i]. enlighten [Doy95]. enough [Rob94d]. enriched [Oli94, Oli94]. entering [Mae95c]. Enterprise [Ano96i]. Enterprise-Wide [Ano96i]. Enters [Ano90m]. entertainment [Phi95a, Coo92]. entertains [Rob94d]. entirely [Coh96]. Envelope [Rob97e]. environment [Rob95a, Rob95b]. Environmental [Doy91b]. environments [Rob95h]. equipment [Nov95]. Erasable [Sho93a]. ergonomics [Emm96c]. ESCOM [Sha95]. established [Pot96c]. Europe [Har96c]. Europeans [Har96c]. Evaluating [For93a]. Even [Tuc95a]. Events [Ano98-44]. ever [Rob94c]. ever-increasing [Rob97b]. every [Kin97]. Everybody [Emm92c]. everything [Mae95c]. Everywhere [Sor91e]. evidence [Ano96-79]. ex [Ano98-38]. Example [Haa94]. Excellence [Ano92i]. exception [Cra94]. exchange [Ano95a]. excitement [Rob95h]. exciting [Phi95a]. excluded [Cra94]. exhibit [App93]. exhibition [Doy95]. Exhibits [Ano98-69]. Expand [Har96a]. expensive [Nov96]. experience [Rob96b]. Experiencing [Ano96l]. experimenting [Nov96]. expertise [Ano96-49]. explode [FA95]. explodes [Sch97e]. Exploration [Ano98-61]. Exploring [Ano96-77, Sor91b]. explosion [Emm96b]. expressing [Emm94]. expressions [Mah97]. expressive [Emm94]. Extend [Ano96-63]. extraordinarily [Phi96c]. Extreme [Ano96-58, Ano97-45]. exude [Ano96t]. Eye [FK90].

Fabric [Phi96c]. Face [Rob90d].

face-tracking [Mah97]. facelift [Ano96-79]. Faces [Por90a]. facial [Mah95c, Mah95c]. facilities [Phi96c]. Facing [Bri92]. fact [For97]. Factor [Pot91b]. Factories [Boy98b]. factors [Pot96b]. fairly [Emm95a]. familiar [Tuc95a]. Family [Ano90i]. fancy [Ano95c]. fans [Ano95-60]. fascinate [Doy95]. Fascinating [Rob94i]. fashioned [Kin97]. Fast [Ano90e]. fast-action [Rob96b]. faster [Har96b]. favorite [Doy94]. FEA [Ano97-59, Por90a]. feature [Rob94f]. Feature-based [Ano97e]. featured [Sch97e]. features [Emm95a]. feel [Har96c]. Feeling [FA95]. Feels [Hod98a]. fetus [Ano96-71]. fever [For97]. Few [Pot94]. fi [Rob95b]. fiction [Rob96b]. field [RR97]. Fight [Pot97b]. Fighters [Ano98-67]. figure [Ano96-75]. Film [Sor91d, Mac90c]. filmmakers [Ano96-76]. films [Coc95e]. Filters [Rob94i]. Final [Coc98c]. finally [Pot96d]. find [Nov96]. Finding [Hol94]. Finish [Ive92c]. finishes [DeY95]. finite [Pot96a]. finite-element [Pot96a]. fire [Ano96-75]. Fired [Phi97a]. firm [Ano96-73]. First [Ano90j]. five [Phi97a]. Flat [Sch97a]. flawed [Coh96]. floating [Ano97-51]. flow [Har96a]. flowing [Emm94]. Flubber [Ano98-40]. fly [Ano95o]. Flying [Rob92f, Mil93]. flythrough [Ano96-77]. Focus [Ano95-60]. Focusing [Ano96b]. Foe [Rob90b]. follow [Phi95e]. font [Emm95a]. fooled [Coh96]. footage [Phi96e]. Force [Fic86]. Fore [Coo90b]. Forest [Pfi92c]. forests [Phi96a]. form [Ano95-45]. Format [Emm96d]. former [McC96]. Forms [Ano91-31]. formulated [Pot96e]. Forum [Ano96-69]. Fractal [Ano95-41]. Fractals [Sor90d]. France [Ano95-54]. frantic [For97]. Free [Ano96j]. Freeform [Wol93]. freelance [Mah95c]. Freightliner [Phi97a]. French [Ano96-49]. Fresh [Rob94c]. Friend [Rob90b]. friendly [Phi96g]. front [Mae95c, Ano91n]. Frontier [Coc98c, Fre90a]. Fruits [Sor90d]. frustrated [FA95]. FT [Ano97-45]. full [Sch97e]. full-featured [Sch97e]. Fun [Phi95a, Ano92j]. functionality [Coc96a]. Fusion [Lan92b]. future [Emm94, CGW92a]. futuristic [Phi96e].

gaining [Yar95]. Galactic [BS92]. Game [Coc94a]. Gamers [Coc96b]. Games [Mah92c, Rob93a]. gamut [Hel97]. Gap [Rob90a]. Garde [Mah94c].

Gate [Ano95-63]. geared [Coc96b].

generated [Ano95-56].

generating [Rob95h]. generation [Sim95a]. generator [Ano95-42]. geological [Ano98-65]. Geri [Rob98e]. German [Sha95]. get [Sch95b]. Gets [Vas90a]. Getting [Ano97f, Fre90b]. giant [Ano98y, Ano96k]. GIS [Mah98a, LS90]. give [Ano96-27]. Given [Mae96]. gives [Ano95-54]. gladly [Vin97]. gloomy [Pot97b]. glove [Ano95-51]. go [Mae95c]. Gods [Rev94]. Godzilla [Ano98y]. Goes [Ano90h]. Going [Ano96-74]. Gold [Vas90a]. Golden [Ano95w, Ano95-63]. Gone [Tuc96]. good [McC96]. got [Rob94f]. grade [Ano95-62]. Gradient [Ano95k]. Grand [Ano97-31]. Graphic [KPV91]. graphical [Pot96e, Tes94]. Graphics [Sor90c]. Graphisoft [Ano96-30]. great [McC96, Ano96-81]. greater [Phi97e]. Ground [Ano91a]. Group [Ano96-66, Ano95-46]. growing [Pot95a]. Grows [Ano98-32]. Guardians [Emm91a]. Guide [Woo93, Ano92y].

h [Rob98h]. hair [Rob95c, Sor90c].

Hair-Raising [Sor90c]. Hall [Ano96-79]. Halley [Tru86]. hand [Ano96-27]. hand-drawn [Ano96-27]. handcrafted [Ano95-54]. handful [Pot95a]. handling [Pot96d]. hands [Pot97a]. Hard [Fly92]. Hard-Copy [Fly92]. hardware [Oli94]. Harmony [Ano98-61]. Hash [Ano95j]. haven’t [Pot96a]. having [Rob94c]. HDTV [Rob90e]. Head [Ano95-32]. Head-mounted [Ano95-32]. headache [Tuc95a]. Heads [Kin93]. healing [Ano96-67]. Health [McM90b]. Healthy [Ano91c]. heart [Ano95-57, Ano96k]. heart-valve [Ano95-57]. Heat [Ano90c]. Heights [Nov98b]. held [Doy95]. Hell [Gra93a]. Hell-Bent [Gra93a]. help [Ano95-63]. helped [Ano98b]. helping [Tho95b]. helps [Ano95-60]. Henk [Ano96-35]. Here [Hir96]. High [Mac90b, Sch92]. high-end [McC96]. high-flying [Pot97b]. High-Res [Hel97]. High-speed [Ano97-54]. highlight [Ano95-59]. Highlighting [Tru86]. HiJaak [Ano96-52]. history [Ano97-53]. hit [Rob95a]. Hits [Mac92]. hobbyist [Row96]. Hollywood [Row97e, Sho91]. Hologram [Emm91b]. holograms [Ano96-66]. Home [Ano96-81]. Homer [Ano95-59]. honest [Rob96a]. honor [Ano96n]. hot [Rob95a]. Hottest [Ano96n]. Houdini [Ano97-43]. house [Har96d]. houses [Rob95c]. HP [Ano90j]. HP/Apollo [Ano90j]. HSC [Ano95-40]. huge [Ano96-64]. Human [Pot91b]. humans [Rob95c, Coc95c]. Hunchback [Ano96-29]. hundreds [Gar95]. Hungry [Ano98m]. hurdles [Sim95a]. Hurrah [Ano97-57]. hurricanes [Phi96a]. Hybrid [Smi96b]. Hypermatter [Ano97-42].

I-DEAS [Ano97-39]. IBM [Ano90k, Ano91o]. ideas [Rom97]. If [Tuc96]. II [Ano95-62]. ill [Ano95-53]. Illusion [Rob98d]. Illustration [KPV91]. illustrations [Ano95-63]. illustrator [Ano95s]. ILM [Rob96c]. Image [Fre90b, Coo90b]. image-editing [Rob94c]. imagery [Phi96f]. imagery-provided [Phi96e]. images [Coc96a, CGW92b]. Imageworks [Ano96x]. imagine [Rob94c]. Imaging [Ano90n, Ano91p]. Imax [Phi96e]. immersive [Rob96b]. impact [Cad93, Pfi92a]. implementing [Mae95c]. important [Rom97]. impossible [Rob94c]. impractical [Phi96f]. impressive [Ano97p]. improved [Oli94]. improvements [Emm96d]. improving [Rob95b]. inch [Ano95-29]. including [Phi95f]. incorporation [Pot96c]. increased [Har96a]. increasing [Emm96d]. increasingly [Phi96f]. Indeo [Ano95-52]. Industrial [Ano90h, Rob94f]. industries [Pot96d]. industry [Hir96, Ano91q]. Inexpensive [Mac93b]. infertility [Ano95-58]. Infini [Ano95-38]. Infini-D [Ano95-38]. InfiniteFX [Ano95-46]. ingredient [Pot96c]. ink [Ano97-33]. inkjet [Gar95]. Innovation [Ano97p]. innovative [Ano95b]. Input [Ano91r]. Inquisitor [Ano97-31]. Ins [For97, Rob95d]. insight [Ano97-63]. install [Ano96i]. installation [Ano97-55]. installment [Rob96b]. IntelliCAD98 [Ano98-48]. Intelligent [Coc97c]. Intel’s [Ano95-52]. interact [Rob95h]. Interacting [Coc97e, Rev94]. Interactive [Coc94b, McM90b]. interactively [Rob94c]. Interactivity [Coc95e]. interest [Coc96b]. interested [Tuc96]. interesting [Rob94d]. Interface [Bri92]. interfaces [Phi96g]. Interfilm [Coc95e]. intergalactic [Ano97k]. Intergraph [Ano95w]. Interiors [Lan90]. International [Ano98-29]. Internet [Ano98j, Tes94]. Internet-related [Pot96e]. Interoperability [Row97b]. interplay [Mah97]. Interview [Por92]. introduced [Ano96n]. Invasion [Sav98]. invest [Mae96]. invisible [Ano97n]. involves [Mah95c]. Iron [Ano96-28]. Island [Ano97l]. issues [Tuc95a, Bri92]. itself [Ano97-61, Rob92a]. jaded [Emm95a]. James [Ano96k]. Japanese [Ano97-60]. Jaron [Por92]. Java [Ano96-63]. jet [Ano97-33]. job [RR97]. Journeys [Emm98]. Juiciest [Ano91g]. Jumanji [Rob96c]. Jurassic [Rob96c]. just [Mah95c].

Kai [Ano96-56]. Kernel [Ano98f]. kernels [Pot97d]. key [Ano96-79]. Kicking [Bri93]. Kid [Rob94d]. Kids [Ano95-63]. kind [Ano96-80]. Kinetix [Ano96-54]. Kingsley [Ano98t]. Kit [Spr94]. Knowledge [Ano95-44]. KPT [Ano95k].

L5 [Phi96e]. lab [Ano97-51, Mah92e]. Landscape [Hir96]. landscapes [Ano97x]. language [Ano95f, Vac96]. Lanier [Por92, Ano95x]. Laptop [Ano91-29]. Large [Mac93c]. Large-Format [Emm96d]. Large-Screen [Mac93c]. Lascaux [Ano95-54]. Laser [Emm95b]. lasers [Ano97-54]. last [Cla96]. latest [Ano96-49]. launch [Ano98b]. Launches [Ano90g]. Launching [Phi98]. Law [Alo94]. Lawnmower [Pfi92d, Ano95-62]. lawsuit [Ano96-79]. layout [Rob94c]. LBE [Ano97m]. LCD [Ano95-33]. LCDs [Emm91c]. leaders [Rob96a]. leading [Ano97p]. leaky [Ano96-79]. leaky-roof [Ano96-79]. Learn [Ano98-58]. LED [Ano96-41]. Lee [Ano95x]. Lego [Ano97l]. LenZFX [Ano95-42]. Lessons [Ano98-37]. Let [McC96, Mah92c]. Lets [Rob94b]. Letters [Ano91k]. level [Sch95b]. libraries [Ano95-30]. licking [Pot97b]. Life [Ano91-31, Lan92a]. Light [Rob94f]. Lighting [Mac94]. Lightscape [Ano98-49]. LightWave [Mae96, Ano95-44, Ano95-41]. Like [Rob94c]. limitations [Ano96i]. Line [Ano90k, Ano90j]. lines [Mae95c].

Linker [Ano95-38]. lip [Mah95c]. Lips [Rob97f]. little [Pot96g]. Live [Ano91s, Ano95-40]. live-action [Phi96e]. living [Rob96c]. lizard [Ano98y]. local [Ano96i]. locale [Ano96-77]. long [Sha95]. Look [Coo93]. Looking [Pot96g, BZJ94]. Looks [Ano98s]. Lost [Nov93c]. lot [Cla96]. Love [Cra91]. Love/Hate [Cra91]. Low [Fic86, Ano95-48]. low-cost [Gar95, Ano95-48]. Low-End [Fic86]. Lower [Ano96-32]. Lower-end [Ano96-32].

M [Rob95a]. Mac [Ano90a, Ano91t]. Mach [Ano96-51]. Machina [Ano98-38]. Machine [Gro91]. machines [Rob96a, Rob92f]. Macintosh [Ano90e]. Macromedia [Ano96-63]. MacVisualization [Ano90d]. Made [Pfi90a, Por90b]. Magazine [Ano95-63]. Magazines [Mag91]. Magic [Gra93b, Sor92]. Magical [Rob94k]. magically [Woh97]. Magiera [Ano95s]. Magnet [Ano96-75]. Magnetic [Ano98-27]. maintaining [McC96]. major [Tuc95a]. make [Ano95-62]. makes [Pot95d]. Making [Spr94, Sor91c]. Man [Gro91, Pfi92d]. managed [Phi97a]. management [Ano96i]. Managers [Emm86]. Managing [Emi91]. Mania [Rob92b]. Mannequins [Ano91v]. manufacturer [Sha95]. manufacturing [Tho95b]. many [Rob94a]. Map [Ano98-70]. mapping [Rob96d, Sen90b]. maps [Rob94c]. Market [Fic86, Ano90c]. marketing [Gar95]. Marketplace [Mac90c]. Marks [Sch92]. married [Sch97e]. Mars [Ano96l]. mascots [Har96d]. Mask [Rob94f]. Master [Ano96-58, Ano95j]. Masters [Ano95u]. Match [Vas91]. matches [Mil95]. Math [Sta90]. mathematical [Phi97e]. Matra [Ano95-41]. Matrox [Ano95-43]. Matte [Rob94k]. Max [Ano98-45, Mae96]. maybe [Rob97b]. MCAD [Ano91c]. MCAD’s [Ano98-31]. Me [Gra96]. Mean [Nov93b]. meaning [Coc95e]. Meat [Ano92l]. Meat-and-Potatoes [Ano92l]. Mechanical [Pot93b]. Mechanism [Pot92a]. Media [Ano95-46]. Medialab [Har96d]. Medical [Phi95d]. medicine [Ano97-55, Mah90]. medium [Rom97, Pot95f]. Meet [Rob98e]. Meets [Mac91b]. Mellon [Ano97-54]. Merging [Mah98a, Gro91]. MeshPaint [Ano95-39]. Metaballs [Gra93b]. MetaReyes [Ano96-55]. MetaTools [Ano96-56]. method [Pyr96]. methods [Pyr96]. MGA [Ano95-43]. MGM [Ano91u]. Michanicol [Ano98-49]. Microcadam [Ano95-43]. MicroScribe [Ano97-41]. Microsoft [Mae96]. MicroStation [Ano95-39]. Microvisualization [Fri91]. Mid [Kur97]. Mid-Range [Kur97]. midrange [Rob96a]. Midranqe [Pot98a]. might [Pot95d]. military [Ano96-81]. Millennium [Ano95-43]. mimic [Phi96a]. mind [Ano95-51]. Minds [Mah91b]. miniseries [Spr94]. MIPS [Ano91b]. Miracle [Emm91b]. misconceptions [Woh97]. misleading [Coh96]. mission [Ano97u]. Mix [Vas91]. Mix-and-Match [Vas91]. Mixing [Ano98-28]. Mode [Mah94e]. Model [Ano94g, Ano94h]. Modeler [Ano95-39]. modelers [Phi95e]. Modeling [Rob91a, Pot92b]. Modelling [Ano92-37]. models [Rob94c]. models-is [Rob94c]. Modern [Pfi92a]. modest [Nov95]. modular [Ano96y]. Module [Ano96-55]. mold [Tho95a]. money [Pot96g, Ano95-63]. monitor [Tuc96]. monitors [Ano96-44, Mac93c]. month [Doy95]. months [Phi97a]. Moonstruck [BM98]. morphing [Coc96a, Sor92]. Most [Ano92-45]. Mother [Phi96a]. motion [Rob94a, Pfi91]. Motion-Based [Phi96f]. motion-capture [Rob95b]. Motion-captured [Ano96m]. Motion-rapture [Mae95c]. mountains [Phi96a]. mounted [Ano95-32]. Mouse [Ano95w]. Movable [Ano91v]. Move [Mah93c]. movement [Mah95c]. Moves [Pot92a]. movie [Ano95y]. Moviemaking [Mah92e]. Movies [Phi96f, Mag92]. moving [Rob95b, Rob92c]. MTV [Ano96-71]. much [Cra94]. Multi [Ano97-53]. Multi-threaded [Ano97-53]. Multimedia [Ano90c, McM90c]. multiplatform [Cra94]. multiplexing [Ano96-68]. muscles [Mah95c]. Music [Doy94, Pfi91]. must [Phi95e]. My [Rob97f]. myriad [Mah95c]. Mysteries [Phi95f]. Myths [Ano98r, Mac90c].

NAB [Ano95-37]. NASA [Ano96-64].

National [Ano95-60]. nationwide [Coc95e]. Natural [Kee86]. Nature [Phi96a, Ano97-42]. navigate [Nov95]. nearly [Smi96a]. necessary [Ano95c]. need [Mil95]. needed [Coc95f]. needs [Row96]. negotiating [McC96]. Neon [Haa94]. NetRenderMan [Cra94]. Nets [Phi97f]. Network [Cra94, Lie93]. network-rendering [Cra94]. never [Doy94]. newly [Pot96c]. News [Ano95-49]. NewTek [Ano95-44]. Next [SR90]. NICE [Ano98-58]. nine [Phi97a]. no [Smi96b]. not-so-subtle [Mah97]. nothing [Woh97]. novel [Nov96]. NT [Cad93, Por95a]. number [Por94]. numerous [Gar95]. NURBS [Ano97-41]. NYC [Coc95e]. NYC-based [Coc95e].

O [Rob92b]. object [Pot96c]. object-oriented [Pot96c]. objections [Pot96a]. Objects [LeW93]. obstacles [Sim95a]. oceans [Phi96a]. Odyssey [Ano97-50]. Off [Coc95f]. offer [Gar95]. offering [Coc96a]. offerings [Pot96c]. offers [Mac95]. often [Sim95a]. Oil [Ano98-28]. old [Ano95h]. old-fashioned [Kin97]. Olympic [Ano96-77]. Olympics [Ano95-61]. once [Rob96c]. one [Rob94c]. only [Doy94]. Onto [Coc94c]. opener [Ano96-71]. OpenGL [Ano96d]. Opens [Nov95]. Opera [Ano98s]. operating [Mae96]. opportunities [Phi95a]. Opportunity [Mae96]. Optical [Sho93a]. optical-effects [Ano95-42]. Optimal [Pot96a]. optimism [Emm94]. optimization [Pot96a]. optimize [Ano95-57]. Optimized [Emm92b]. Option [Row96, Ano90a]. Options [LB91b]. Orbit [Ano98-56]. Order [Mah93d]. ordinary [Sch97e]. organic [Ano98-34]. oriented [Pot96c]. other [Rob95h]. Our [Por91b]. Outer [Fre91]. Outlook [Ano98-30]. output [Yar95, LB91b]. Outputting [Pyr96]. outstripping [For97]. overcame [Fan96]. overview [Ano92a]. own [Ano97q].

pace [For97]. Pacific [Ano95-52]. Pack [CC98]. Package [McM91, Ano90g]. packages [Cra94, Ano92b]. PACS [Fre90c]. page [Rob94c, Joh90]. page-layout [Rob94c]. pager [Ano97s]. pages [Ano95-63].

Paint [Rob94c, Ano95-42]. Painter [Ano96-56]. Painting [Rob91b, Rob91d]. Paints [Dun97]. panels [Ano95-33]. Paper [Ano91j]. Parade [Rob94g]. Paradise [Fre90c]. Parametric [Ano95-45]. Parasolid [Pot97d]. Paris [Har96d]. Park [Mac90d, Ano95-52]. Part [Mil95, Ano90h]. Particles [Rob93e]. particular [Row96]. parts [Pot96d, Tho95b]. passive [Phi95e]. past [Hir96, Mah91c]. patience [Coc97d]. Pattern [Ano97-53]. Pattern-recognition [Ano97-53]. Pay [Ano98-73]. PC [Ros90, Mac91b]. PC-based [Ano90g]. PC/ [Ano90l]. PCB [Pfi90b]. Pcholkin [Ano95v]. PCs [Sch95b]. PDM [Ano96i]. Peach [Ano96k]. Peek [CGW92a]. [Ano90l]. Apollo [Ano90j]. CAM [Ano95-47, Ano95f]. E [Ano95-58]. Engineer [Ano95p]. Eye [Ano95-38]. FastSOLID [Ano97-37]. graphics [Ano95-34]. Hate [Cra91]. JR [Ano95-45]. Kinetix [Ano96-60]. Photo-Realism [Ano90f]. server [Tuc95a]. Wavefront [Ano98-45]. Pennello [Ano95h]. Pentium [Sch95b]. People [Bel93]. perfect [Ano95o, Cav92]. perfectly [Row96]. perform [Har96d]. performance [Rob96a]. performed [Rob96c]. Performers [Rob97d]. personal [Ano97s]. personality [Ano96t]. phenomena [Phi95f]. Philadelphia [Phi97c]. Photo [DeY95, Rob90a]. Photo-Paint [Ano95-42]. photo-quality [Emm96a]. Photo-Retouching [Rob90a]. photogrammetric [Kin97]. photographer [Ano95v]. photographic [Nov95]. photographic-quality [Nov95]. photographs [Kin97].

Photography [Ive92b]. Photorealism [Emm94]. photorealistic [Gar95]. Photoshop [Ano92j]. Physical [Rob93d]. picking [Mil95]. picture [Ano95o, Cav92]. picture-perfect [Ano95o, Cav92]. Pictures [Pot95f, Rob92c]. Pictures/ Kinetix [Ano96-60]. pie [Sch97e]. Piecing [Ano97-53]. Pioneers [Rob97d]. Piranesi [Ano98-50]. piration [Rob98a]. Pistons [Ano95-60]. pitfalls [Ano95d]. pivotal [Mac94]. Pixar [Cra94, Ano96-60]. PixelPutty [Ano95-46]. place [Phi96c]. Plain [Ano91j]. Planets [Fre91]. Planning [Ano95-61, Yar95]. Plant [Ano98h]. plastic [Tho95b, Sho90a]. play [Doy94, Bel93]. players [Por94]. playground [Ano95-53]. Playing [SH90]. plays [Ano96-81]. Plotters [Yar95]. plug [Ano95-41, Rob95d]. plug-in [Ano95-41]. Plug-Ins [For97, Rob95d]. Plus [Ano95-42]. point [Smi96a]. poised [Pot97b]. popular [Yar95]. popularity [Yar95]. populated [Rob95h]. Portable [McM92]. Portfolio [Ano95v, Ano91w]. Portifolio [Ano92x]. Poser [Ano97-38, Ano95-41]. Positran [Ano96-55]. Positron [Ano95-39]. post [Emm95a]. post-production [Emm95a]. post-sales [McC96]. poster [Ano96-28]. Potatoes [Ano92l]. potential [Hel97]. Pouring [Sho90a]. Power [Kru93, Por95a]. PowerAnimator [Ano96u]. powerful [Sch97e, Rob93e]. powers [Ano96j]. Practical [Woh90]. Practices [Pot95d]. pragmatic [Rob94a]. Predicting [Cad93]. Prehistoric [Ano98q]. Premature [Fre90c]. Premieres [Ano90f]. Preparing [Emm95c]. prerendered [Coc95f]. presentation [Ano92b]. Presentations [Rob90d]. presidential [Ano96-64]. Press [Spr94]. pretty [Cra94]. Preview [Doy95, Ano92-38]. previous [Pot96c]. previously [For97]. price [Pot95f]. price-performance [Rob96a]. prices [Sch95b]. Prime [Ano91s, Rob93f]. Prime-Time [Rob93f]. Principle [Ano92-45]. Print [Rob91a]. printers [Gar95]. Printing [Ano91j, Cav92]. priorities [Phi96d, Ano91-27]. Prix [Ano98-43]. Pro [Ano95-42, Ano95p]. Pro/E [Ano95-58]. Pro/ Engineer [Ano95p]. Pro/JR [Ano95-45]. problems [Coc97d, Rob94a]. Procedural [Phi97e]. process [Tho95b]. processing [Rob94c, Fre90b]. produce [Ano96-32]. produced [Ano96-49]. Producers [Phi96f]. product [Ano95-37, Ano91y]. production [Emm95a]. production-worthy [Kur97]. products [Pot95a, Ano93-37]. Professional [LB91c]. professionals [Har96a]. program [Ano95-42]. Programs [Coc94b, Mac95]. progress [Ano96-67]. progressed [Smi96a]. project [Rom97]. projection [Ano95-33]. Projector [Sch97b]. projectors [Ano96-38]. proliferate [Ano96j]. Promise [McM92]. promises [Doy95]. promising [Sch95b]. promote [Spr94]. prompt [Pot96b]. prompting [Phi96g]. Pronto [Ano98-41]. Proofs [CP92]. proposed [Ano95-62]. Proprietary [Ano98b]. prospective [Ano96i]. prosper [Ano96f]. Protecting [Ano92-30]. Protection [Doy91b]. Prototypes [Mey97, Woh90]. Prototyping [Mah95d]. provide [Har96a]. provided [Phi96e]. provides [Kur97]. province [For97]. Proving [Rob93f]. Psygnosis [Ano97o]. PTC [Ano95p]. publicly [Pot96e].

Publishing [Ano96-55, McM90a]. purchase [Ano96i]. purchasers [Ano92-30]. purchasing [Yar95]. purposes [Pot95a]. Push [Pot94]. Pushed [Pot96d]. Pushes [Ano98j]. pushing [Ano97p]. put [Pot97a]. puts [Ano96-76]. putting [Rob95c].

Q [Ano97-35]. Quadra [Ano91g]. Quake [Ano96l]. quality [Oli94]. quantifying [Coh96]. Quest [Ano93-62]. question [Mae96]. quickly [Sha95]. QuickTime [Nov95, Nov95]. quite [Coc96a].

racing [Ano96m]. radiosity [Ash96]. Raffaello [Ano96l]. rages [Pot97d]. RAID [Har96b]. Rain [Pfi92c]. Raising [Sor90c]. range [Mac95]. Rapid [Ano96-45, Mah95d]. rapid-prototyping [Pot97c]. rapture [Mae95c]. rather [Rob94d]. ray [Ash96, Ano96-50]. Razor [Ano95-42]. re [Ano96-75]. re-create [Ano96-75]. re-creates [Ano96-75]. Reaches [Nov98b]. Read [Rob97f]. Reading [Coh96, Mah91b]. Ready [Rob94j]. Real [Ano91d, Coc95f]. Real-Time [Ano91d, Coc95f]. Realism [For93c, Ano90f]. realistic [Ano95-27]. realistically [Phi96c]. Realitie [Ano95-41]. reality [Rob95a, Dav91]. really [Ano95c]. realtime [Ano96-74]. reasons [Pot96a]. ReBoot [Ano95-55]. Rebuilding [Mah91c]. recognition [Ano97-53]. recognize [Ano97p]. recognized [Pot95a]. reconstruction [Ano97t]. Recorders [Mac90c]. recording [Doy94]. Recreating [Mah97]. Recycle [Ano95h]. Redefines [Ano91-28]. Redesigning [Ano96-77]. Rediscovering [Fre91]. reduce [Phi97a]. Rel [Ano96-59]. related [Pyr96, Pot96e]. Relationship [Cra91]. Relative [Mah91a]. relatively [Por94]. release [Pot97b]. releases [Pot96c]. reliable [Mah97]. Relic [Ano97o]. relies [Ano96-74]. relying [Nov95]. Removable [Ano97-32]. Removing [Pot93b]. Renault [Ano96-49]. rendered [Rob95h]. rendering [Ano92-37, Ano91f]. renderings [Ano95-62]. RenderMan [Ano96-60, For93c]. RenderWare [Ano96d]. RenderZone [Ano95-45]. Reordering [Ano91-27]. Repertoire [Ano90n]. replace [Pot96b]. Report [Pot96c]. represents [Phi95a]. Reproducing [Ano97-55]. Republican [Ano95-60]. requires [Emm96c]. Res [Hel97]. Research [Mah92e, Ano95-58]. researchers [Phi95d]. resources [Vin97]. respond [Sha95]. result [Sch97e]. results [Ano96-32]. retina [Ano96-66]. Retouching [Rob90a]. rev [Ano95-60]. revealed [McC96]. Revenge [Ano98u]. Reverse [Ano96-78]. Reversing [Coh98b]. review [Pot95a, Ano92-37]. Reviews [Ano95-42, Ano93-47]. revise [Pot96c]. Revolution [Ano95-44, Rob91b]. Rhino [Ano97-41]. rich [Phi96d]. Richard [Ano96-36]. Richer [Sch98b]. Riddler [Ano98u]. Ride [Rob94j]. ride-simulation [Phi96f]. Rides [Mac90d]. Riding [Phi96f]. right [Emm96c]. Riven [For98]. Road [Has92]. roar [Ano97n]. Rob [Ano95s]. robot [Ano95j]. robust [Coc97d]. Rod [Mah95c]. Rodin [Rob92d]. role [Ano96-49, Por90a]. Roll [Rob92d]. Rolls [Ano90j]. ROM [Spr94, Ano95-63]. ROMs [Doy94]. roof [Ano96-79]. room [Ano96-65]. Rooms [Ano96l]. Round [Cla96]. Roundup [And90]. Rousing [Fri92]. roving [Ano95-51]. Roy [Ano95u]. Royal [Ano96-79]. rules [Phi95e, Por91a]. Run [Pfi94a]. runs [Hel97]. Russian [Ano95v].

saddle [Ano96-78]. safely [Pot97a]. saga [Rob96b]. sales [McC96]. salesperson [McC96]. same [Cla96]. sampling [Ano96z]. San [Ano95-60]. Sav [Rob98h]. scan [Ano96-40]. scanner [Mil95]. Scanners [Mil95]. scanning [Phi96b, Hol93]. scene [Mac94]. Scenery [Rob90c]. scenes [Spr94, Nov95]. sci [Rob95b]. sci-fi [Rob95b]. Science [Phi95f, Ive92a]. science-fiction [Rob96b]. scientific [App93]. scientists [Phi95f]. Scores [Ano91h]. scorpion [Ano98-33]. scramble [Sch97e]. Screen [Sor90b, Sor90b]. Screens [Ano98p]. Sculpting [Rob90c]. sculpts [Ano96-28]. Search [Emm91b]. Second [Coo93, Ano97-42]. Secret [Ano98u]. secrets [McC96]. See [Emm86]. Seeing [Mah91d]. seems [Rob94c]. Segments [Kee86]. Select [Pot94]. Selecting [Sch97b]. sell [Ano95-63]. semiconductor [Ano97-53]. Sense [Mac93a]. sensory [Rob96b]. sequel [Ano96l]. Series [Ano97-39]. Serious [Ano91o, Coo92]. seriously [Ano95-53]. serve [Har96d]. service [Vin97]. Set [Rob95h]. Set-Up [Rob95h]. Sets [Ano90b]. several [Pyr96]. SGI [Ano96-63, Ano91-28]. shaded [Ano95r]. Shades [Rob92e]. share [Rom97]. shared [Rom97]. Sharing [Ano98i]. Sharks [Ano98q]. Shine [Spo98]. shines [Ano96-49]. Shining [Haa94]. shipped [Ano96n]. Shop [Ano94g, Ano94h]. Shopping [Mil95, Vas90d]. short [Ano98c]. should [Mac95]. Show [Has92]. showcase [Ano96-49, Ano95y]. showing [Rob96d]. showpiece [Pot96e]. shy [Ano96w]. siblings [Pot97c]. Side [Ano97-43].

SIGGRAPH [Ano95t, Ano92-38]. Significant [Phi96g]. significantly [Pot96c]. Silicon [Ano90l]. Silver [Sor90b]. simple [Mah97]. Simplifies [Ano98h]. Simpson [Ano95-59]. simulated [Ano96-77]. simulates [Ano96-69]. Simulating [Ano95i]. simulation [Ano95q, Ano95-52]. simulations [Phi95d]. sites [Ano96k]. size [Emm96c]. Sketch [Ano92-37]. skillful [Phi97a]. skills [Coc95f]. slice [Sch97e]. slide [Ano95p]. Small [Fre90b, Por94]. Smaller [Pot97c]. Smart [Rob90g, RV90]. smile [Mah97]. Smithsonian [Ano98-69]. Snapshots [Ano95-63]. sneak [Ano95t]. snowboards [Ano95-58]. Softimage [Ano96-56, Mae96]. software [Rob94d, Rob90g]. software-which [Rob94c]. Soldiers [Ano98-51]. Solid [Pot91a, Ano97-40]. SolidDesigner [Ano97-37]. SolidView [Ano97-41]. SolidWorks [Ano96-58]. Solo [Ano95-46]. Solution [Pot91a]. some [Doy94]. somehow [Woh97]. something [FA95, Emm92c]. Sony [Ano91-29]. sophisticated [Doy94]. Sorting [Pot96g, Mac93c]. sorts [Ano96-64]. Sound [Ive92a]. soundly [Pot97b]. Space [Min94]. sparks [Ano96-30]. spatial [Ano92-30]. special [Rob94f, Ano95-48]. special-effects [Rob94f]. specialized [Ano95g]. species [Rob95b]. specifically [Pot95a]. Specular [Ano95-38]. speech [Mah95c]. Speed [Ano96-51]. Spin [Sch95c]. Spinal [Ano98-70]. Spoken [Hod98c]. spokescharacter [Ano98-68]. Sports [Row98c]. Spotlight [Ano96-62, Mac91a]. spread [For97]. Squeezes [Coo90b].

Squeezing [Coc94c]. stabilization [Vac96]. stage [Ano97-61, Por90b]. stand [Emm96a, Ive91]. standard [Ano95a]. Standards [Ano90b]. Stanford [Ano97-54]. Stardent [Ano90m]. stars [Rob96c, Mah91d]. Starship [Ano97z]. starting [Kur97]. State [Pfi93b]. Statements [LB90b]. Station [Min94]. step [Vin97, Ano95a]. Stereo [Rob90f]. Stereoscopic [Sch97d]. still [Por94]. Stimulating [Coc98d]. sting [Ano98-33]. stood [Ano97p]. stopping [Phi96a]. storage [Har96a, Lyn95]. Story [War98b, Rob95f]. Straddles [Ano90l]. Strata [Ano97-39, Ano95-39]. strategies [Pot96e, Lyn95]. Streamline [Emm96d]. Streamlining [Pfi90b]. streets [Cla96]. Strim [Ano95-41]. string [Ano96-49]. striving [Phi96a, For93c]. Stroke [Dun97]. Strokes [Rob97c]. Stroll [Ano91d]. strong [Row96]. structure [Ano95-63]. student [Ano95-62]. Students [Nov93b]. Studio [Ano96j, Pop94, Ano95l]. StudioPro [Ano97-39, Ano96-55]. Studios [Rob94e]. Stuff [Mah93e]. Stunning [Ano96-33]. style [Emm94, Sho91]. Stylist [Row98a]. sub [Emm96a]. Subconscious [Ano98n]. sublimation [Gar95]. subtle [Mah97]. subtleties [Emm96c]. Success [Hol94]. such [Yar95]. suited [Coc95f]. Summer [Ano95y, Ano96-67]. Sun [Ano96-63]. sunlight [Ano95i]. Super [Sto90, PF90]. Supercomputer [Ano90m]. Supercomputers [Ano90b]. Supplement [Ano95-47]. support [Rob94d]. surface [Smi96a]. Surfaces [Smi96a, Wol93]. Surgeon [Ano97-63]. surgery [Ano96-69]. surprisingly [Por94]. Survey [PF90]. Survived [Rob97h]. sustain [Sha95]. swear [Rob96c]. sweeping [Ano97j]. Swirling [Ano91-31]. Sync [Ano95-42]. synced [Rob95h]. SyQuest [Ano96-77]. system [Rob95b]. systems [Mae95c, Ano95-39].

T2 [Rob96b]. Tablets [Emm96c]. Tactics [Rob96d]. tailored [Pot95a]. take [Emm94]. taken [Pot97b]. Takes [Ive91]. taking [Cra94, Has92]. Tale [Pfi92a]. Talk [Ano93-49]. talking [Pot96e]. tall [Ano98-51]. Tangled [Rob90e]. Targeting [Sho90b]. Targets [Ano91-30]. task [Mah97]. teach [Rob94d]. teaches [App93]. Team [Ano90j]. Tech [Ano95-51, Ano92-39]. Technical [Sim95a]. Techniques [Rob95c]. technological [Coc95a]. technologically [Phi96f]. technologies [Doy94]. Technology [Rob90b, Lan92b]. technology-rich [Phi96d]. TechWatch [Ano98-62]. teeny [Ano97-54]. teeny-tiny [Ano97-54]. Tektronix [Ano90n]. Telemedicine [Ano96-67]. Television [Spr94]. Tells [War98b]. temple [Ano97r]. tens [Pot96d]. term [Coc95e]. Terminal [Win90]. Terminator [Sor91d]. terms [Emm96a]. Terrain [Lan93]. test [Coh96]. texture [Rob94c]. textured [Rob95b]. textures [Phi95f]. TFX [Row97e]. Thankfully [Mac95]. Thanks [Vac96]. theater [Ano95-56]. theaters [Coc95e]. Theatre [Ano95t, Rob91e]. theatrical [Ano97-61]. their [Doy94]. them [Yar95]. Theme [Mac90d]. theme-park [Rob96b]. Therapy [Phi97f]. There [Pyr96]. thermal [Gar95]. these [Pot95a]. they’re [Yar95]. think [Ano97a]. Thinking [Sch97d]. third [Rob96b]. those [Rob94c, Rob92f]. though [Rob96d]. thousands [Pot96d]. threaded [Ano97-53]. threat [Ano96h]. three [Phi97a]. Thrill [Ano98-70]. throat [Mah95c]. Thumbelina [Pfi94b]. Tidal [Ano98-69]. Time [Ano91d, Dav92]. time-consuming [Nov96]. Times [Pot92a]. tiniest [Ano97-55]. tiny [Ano97-54]. tips [Ano95q]. Tires [Bri93]. Titanic [Ano97z]. title [Ano97n]. titling [Emm95a]. Toaster [Ano91-34]. today [Doy94]. together [Ano97-53]. tomorrow [Phi97e]. tongue [Mah95c]. too [Nov96]. Took [Coo93]. Tool [LS90]. Toolbox [Coc94a]. Toolmaker [Ano98q]. Tools [McM91]. Toons [Rob94h]. top [Phi96d]. Tornados [Fan96]. Touch [Gra94, LB91c]. tough [Pot97b]. tour [Ano95-54]. touted [Mae96]. touting [Smi96b]. toy [Ano96-75, Ano95z, Rob95f]. toyland [Mah95a, Phi95b]. tracing [Ash96]. Track [Rob98g]. Tracking [Mah97]. trade [Ano97-52]. traditionally [Pot96a]. tragic [Ano98-68]. Training [Ano97-65]. Transcending [Emm94]. transfer [Gar95]. transitions [Coc96a]. Translation [Ano95-46]. Trapping [Sho93b]. traps [Ano95d]. Travels [Ano98-71, Mah93f]. Treasures [Tes94]. Treatment [Sho90b]. tree [Ano95m]. trend [Pot97a]. Trial [Vas90c]. trio [Ano92b]. trip [Ano96l]. TriSpectives [Ano95-38]. TriStar [Rob96c]. Triumph [Rob95f]. trivial [Mah97]. Troopers [Rob97g]. truck [Phi97a]. True [Ano91-33]. trueSpace [Ano95-40]. trueSpace3 [Ano97-40]. truly [Ano97p]. truth [Woh97]. try [Hir96]. trying [Pot97a]. Tuning [Emm96c]. Turns [Ano98q]. TV [Coc98a, Coc94b]. Twister [Fan96]. two [Ash96]. Type [Emm95a]. types [Rob94a]. typography [Emm95a].

Ultimate [Pic91]. Ultra [Ano95-27]. Ultra-realistic [Ano95-27].

Unbelievably [Rob94f]. Under-USD [Ano96-46]. undergone [Hir96]. underground [Ano96-72]. underlying [Pot96c]. understand [Phi95f]. understanding [Mil95]. uninterrupted [Har96a]. unique [Ano97-52]. university [Phi96c]. university-based [Phi96c]. Unix [Ano90k]. Unlocking [Phi95f]. Unraveling [For98]. UnReal [Ano96-60]. Untamed [Rob95g]. Unto [Rob92a]. Unveils [Ano90e]. Update [Sch97d]. USD [Pot96g]. use [Pot95a]. used [Spr94]. User [Ano95-63, Ano95-60]. user-friendly [Phi96g]. users [Cra94]. uses [Ano95-51]. using [Tuc95a]. utilize [Gar95]. utilized [Pot95d].

V [Ano98-48, Har96a]. V1.2 [Ano95-46]. V1.5 [Ano96-57]. V1.75 [Ano96-55]. V2.0 [Ano95-46]. V2.01 [Ano96-55]. V2.7 [Ano96-59]. V3.0 [Ano96-50, Ano95-38]. V3.51 [Ano97-44]. V4 [Ano95-44]. V4.0 [Ano96-58]. V6.0 [Ano95-41]. V7 [Ano96-51]. valuable [Ano96-49]. value [Pot95f]. valve [Ano95-57]. Vanilla [Tuc96]. variety [Emm94]. various [Hel97]. Vatican [Ano96l]. Vegas [Ano96-74, Ano95-59]. vehicle [Ano95-52]. Vehicles [Row98b]. Vellum [Ano97-46]. vendors [Pot95a]. Venture [Ano91t]. Versailles [Ano97-29]. Version [Ano95-39]. versions [Phi96a]. via [Rom97]. Video [McM90b, Mac91b]. video-capture [Ano96-43]. Video/ graphics [Ano95-34]. videos [Ano96-49]. view [Vac96]. Viewpoint [Ano98-73]. virtual [Ano92-44, PE92]. virtual-model [Hel97]. virtual-reality [Rob95a]. Virtuality [Ano95-45]. Virtualized [Ano97-54]. Visible [LMVP91].

Vision [Ano95-39]. Visiting [Nov93c]. visual [Rob96c, Ano96-61]. visual-effects [Rob96c]. Visualization [Por94, FP91]. Visualizing [Phi97g, Sta90]. Viz [Ano97-36, Rob91c]. Vladimir [Ano95v]. Volis [Ano95-46]. Volume [LMVP91]. volumes [Phi97e]. Vote [Ano96-64]. VR [Bri93, Phi95c]. VRML [Vac96]. vs [Ano96d].

walking [Ano96p, Mah94f]. walkthrough [Ano95-60]. want [Hel97]. Warner [Rob95a]. warp [Coc96a]. warriors [Tuc95b]. Wars [Ano98f]. was [Rob95b]. Watch [Ano95-51, Ano92-39]. Water [Sor91e, Sor91e]. Wave [Phi96f, SR90]. Wavefront [Ano96-51]. Waves [Sor91c]. wax [Gar95]. wax-transfer [Gar95]. way [Ano95-55]. ways [Doy94]. wealth [Coc96b]. Web [Rob90e, Ano96j]. Web-enabled [Pot97e]. weight [Emm96c]. well [Mah95c, Mac90a]. were [Pot96e]. Whether [Gar95]. which [Doy94]. While [Pot95a]. Whirling [Ano91-31]. Who [Nov93b]. who’ll [Vin97]. Wicked [Ano90e]. wide [Rob94a]. Wide-format [Ano97-33]. widely [Pot95f]. Wild [Rob94j, Mac90d]. Wildebeests [Pfi94a]. will [Hir96]. win [Ano95-56]. Window [Mae96]. Windows [Ano92b, Cad93]. Windows-based [Ano92b, Ano97-34]. Winds [Por91c]. Winner [Mah94b]. winners [Ano95w]. winning [Rob94f]. wins [Ano96-73]. within [Rob95b]. without [Rob94c]. Won [Pot96b]. wonder [Smi96b]. wonderland [Ano96-76]. Woolly [Mac90d]. word [Rob94c]. work [Rob94f, Emm94]. Workgroup [Ano96i]. Working [Ano95-44]. works [Ano95x]. Workshop [Ano95h]. Workstation [Ano90j, And90]. workstation-level [Sch95b]. workstation-quality [Sch95b]. workstation/server [Tuc95a]. Workstations [Fic86]. World [Hol93]. worlds [Kin95]. worthy [Kur97]. wound [Ano96-67]. wound-healing [Ano96-67]. Wright [Ano97-64].

X [Win90]. Xaos [Ano95h].

XD88 [Ano90n].

Year [Pfi93a]. years [Tuc95a]. you’d [Rob96c]. you’ll [Coh96]. you’re [Row96]. yourself [Vin97, Sen90b]. you’ve [Pot96g].

Z2.7 [Ano95-45]. ZD [Row96]. ’Zillas [Rob98f]. Zork [Ano97k].

References

Alofs:1994:CCL

[Alo94] J. Alofs. Coping with copyright law. Computer Graphics World, 17(10):40–??, October 1994. CO-DEN CGWODH. ISSN 0271-4159.

Anderson:1990:WR

[And90] Ken Anderson. Workstation Roundup. Computer Graphics World, 13(11):69–??, November 1990. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1990:MAO

[Ano90a] Anonymous. A Mac Animation Option. Computer Graphics World, 13(12):99–??, December 1990. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1990:ASS

[Ano90b] Anonymous. Alliant Sets Standards with Supercomputers. Computer Graphics World, 13(2):23– ??, February 1990. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1990:AAH

[Ano90c] Anonymous. Amiga 3000, AmigaVision Heat Up Multimedia Market. Computer Graphics World, 13(6):29–??, June 1990. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1990:NM

[Ano90d] Anonymous. And Now, MacVisualization. Computer Graphics World, 13(8):141–??, August 1990. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1990:AUW

[Ano90e] Anonymous. Apple Unveils its “Wicked Fast” Macintosh. Computer Graphics World, 13(4):23– ??, April 1990. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1990:APA

[Ano90f] Anonymous. Autodesk Premieres 3D Animation/Photo-Realism. Computer Graphics World, 13(8): 31–??, August 1990. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1990:CLP

[Ano90g] Anonymous. Cadam Launches its PC-based 3D Package. Computer Graphics World, 13(11):19– ??, November 1990. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1990:CID

[Ano90h] Anonymous. CDI’s Industrial Design Goes from Art to Part. Computer Graphics World, 13(1):19– ??, January 1990. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1990:DED

[Ano90i] Anonymous. Digital Enhances its DECstation Family. Computer Graphics World, 13(5):25–??, May 1990. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1990:HAT

[Ano90j] Anonymous. HP/Apollo Team Rolls Out its First Workstation Line. Computer Graphics World, 13(7):25–??, July 1990. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1990:INU

[Ano90k] Anonymous. IBM’s new Unix line. Computer Graphics World, 13(3): 25–??, March 1990. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1990:SGS

[Ano90l] Anonymous. Silicon Graphics Straddles PC/ Workstation Market. Computer Graphics World, 13 (12):21–??, December 1990. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1990:SED

[Ano90m] Anonymous. Stardent Enters the Desktop Supercomputer Market. Computer Graphics World, 13(10):25–??, October 1990. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1990:TAI

[Ano90n] Anonymous. Tektronix Adds Imaging to its XD88 Repertoire.

Computer Graphics World, 13(9): 21–??, September 1990. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:G

[Ano91a] Anonymous. 3D from the Ground Up. Computer Graphics World, 14(5):87–??, May 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:MDa

[Ano91b] Anonymous. 76 MIPS on a Desktop. Computer Graphics World, 14(4):19–??, April 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:HMM

[Ano91c] Anonymous. A Healthy MCAD Market. Computer Graphics World, 14(11):23–??, November 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:RS

[Ano91d] Anonymous. A Real-Time Stroll.

Computer Graphics World, 14(2): 111–??, February 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:Ab

[Ano91e] Anonymous.  AEC.  Com-

puter Graphics World, 14(10):67– ??, October 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:ER

[Ano91f] Anonymous. An Emphasis on Rendering. Computer Graphics World, 14(11):101–??, November 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:AQJ

[Ano91g] Anonymous. Apple’s Quadra:

The Juiciest Mac Yet. Computer Graphics World, 14(11):17– ??, November 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:AS

[Ano91h] Anonymous. Autocad 11 Scores. Computer Graphics World, 14(4): 105–??, April 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:C

[Ano91i] Anonymous.  CAD.  Com-

puter Graphics World, 14(9):69– ??, September 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:CPP

[Ano91j] Anonymous. Color Printing on Plain Paper. Computer Graphics World, 14(7):25–??, July 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:DL

[Ano91k] Anonymous. Designer Letters. Computer Graphics World, 14 (11):73–??, November 1991. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1991:DND

[Ano91l] Anonymous. Designing in a New Dimension. Computer Graphics World, 14(3):23–??, March 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:DVC

[Ano91m] Anonymous. Desktop Video Contender. Computer Graphics World, 14(4):105–??, April 1991. CODEN CGWODH. ISSN 0271-

Anonymous:1991:FE

[Ano91n] Anonymous. Front End. Computer Graphics World, 14(10):6– ??, October 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:IGS

[Ano91o] Anonymous. IBM Gets Serious with Graphics. Computer Graphics World, 14(5):19–??, May 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:Ic

[Ano91p] Anonymous. Imaging. Computer Graphics World, 14(9):75– ??, September 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:Ib

[Ano91q] Anonymous. Industry. Computer Graphics World, 14(9):27– ??, September 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:Ia

[Ano91r] Anonymous.  Input.  Com-

puter Graphics World, 14(9):11– ??, September 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:IIL

[Ano91s] Anonymous. Is it Live? Or is it Prime Time Live? Computer Graphics World, 14(11):6– ??, November 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:MPV

[Ano91t] Anonymous. Mac Painting Venture. Computer Graphics World, 14(6):107–??, June 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:MDb

[Ano91u] Anonymous. MGM on the Desktop? Computer Graphics World, 14(7):119–??, July 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:MM

[Ano91v] Anonymous. Movable Mannequins. Computer Graphics World, 14(1):21–??, January 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:Pa

[Ano91w] Anonymous. Portfolio. Computer Graphics World, 14(9):64– ??, September 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:Pb

[Ano91x] Anonymous. Portfolio. Computer Graphics World, 14(10):64– ??, October 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1991:PSa

[Ano91y] Anonymous. Product Spotlight. Computer Graphics World, 14(9): 21–??, September 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:PSb

[Ano91z] Anonymous. Product Spotlight. Computer Graphics World, 14 (10):17–??, October 1991. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1991:RP

[Ano91-27] Anonymous. Reordering Priorities. Computer Graphics World, 14(11):124–??, November 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:SRB

[Ano91-28] Anonymous. SGI Redefines the Blues. Computer Graphics World, 14(8):25–??, August 1991. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1991:SLU

[Ano91-29] Anonymous. Sony Launches Unix Laptop. Computer Graphics World, 14(2):21–??, February 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:STD

[Ano91-30] Anonymous. Stardent Targets the Desktop. Computer Graphics World, 14(6):21–??, June 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:SWL

[Ano91-31] Anonymous. Swirling, Whirling Life Forms. Computer Graphics World, 14(11):66–??, November 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:Aa

[Ano91-32] Anonymous. The Amiga 3000.

Computer Graphics World, 14(1): 93–??, January 1991. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1991:PTC

[Ano91-33] Anonymous. The PC’s True Colors. Computer Graphics World, 14 (11):85–??, November 1991. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1991:VT

[Ano91-34] Anonymous. The Video Toaster.

Computer Graphics World, 14(3): 89–??, March 1991. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:GMO

[Ano92a] Anonymous. A graphics market overview. Computer Graphics World, 15(3):23–??, March 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:TWP

[Ano92b] Anonymous. A trio of Windowsbased presentation packages. Computer Graphics World, 15(5): 85–??, May 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Ag

[Ano92c] Anonymous.  AEC.  Com-

puter Graphics World, 15(11):65– ??, November 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Aa

[Ano92d] Anonymous. Applications. Computer Graphics World, 15(7):87– ??, July 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Af

[Ano92e] Anonymous. Applications. Computer Graphics World, 15(8):67– ??, August 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Ab

[Ano92f] Anonymous. Applications. Computer Graphics World, 15(10):71– ??, October 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Ac

[Ano92g] Anonymous. Applications. Computer Graphics World, 15(11):71– ??, November 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Ae

[Ano92h] Anonymous. Applications. Computer Graphics World, 15(12):63– ??, December 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:CPE

[Ano92i] Anonymous. Creativity Par Excellence. Computer Graphics World, 15(3):55–??, March 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:FP

[Ano92j] Anonymous. Fun with Photoshop.

Computer Graphics World, 15(5): 74–??, May 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:I

[Ano92k] Anonymous. Industry. Computer Graphics World, 15(7):27–??, July 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:MC

[Ano92l] Anonymous. Meat-and-Potatoes CAD. Computer Graphics World, 15(3):81–??, March 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:MD

[Ano92m] Anonymous. Multimedia Development. Computer Graphics World, 15(3):69–??, March 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:Od

[Ano92n] Anonymous. Output. Computer Graphics World, 15(7):144– ??, July 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Oe

[Ano92o] Anonymous. Output. Computer Graphics World, 15(8):100– ??, August 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Of

[Ano92p] Anonymous. Output. Computer Graphics World, 15(9):108– ??, September 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Oa

[Ano92q] Anonymous. Output. Computer Graphics World, 15(10):108–??, October 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Ob

[Ano92r] Anonymous. Output. Computer Graphics World, 15(11):116–??, November 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Oc

[Ano92s] Anonymous. Output. Computer Graphics World, 15(12):96– ??, December 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Pc

[Ano92t] Anonymous. Portfolio. Computer Graphics World, 15(7):60–??, July 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:Pd

[Ano92u] Anonymous. Portfolio. Computer Graphics World, 15(8):64–??, August 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:Pa

[Ano92v] Anonymous. Portfolio. Computer Graphics World, 15(10):64– ??, October 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Pb

[Ano92w] Anonymous. Portfolio. Computer Graphics World, 15(11):60– ??, November 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Pe

[Ano92x] Anonymous. Portifolio. Computer Graphics World, 15(9):54– ??, September 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:PGd

[Ano92y] Anonymous. Product Guide.

Computer Graphics World, 15(8): 75–??, August 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:PGe

[Ano92z] Anonymous. Product Guide.

Computer Graphics World, 15(9): 71–??, September 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:PGa

[Ano92-27] Anonymous. Product Guide.

Computer Graphics World, 15 (10):79–??, October 1992. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1992:PGb

[Ano92-28] Anonymous. Product Guide.

Computer Graphics World, 15 (11):77–??, November 1992. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1992:PGc

[Ano92-29] Anonymous. Product Guide. Computer Graphics World, 15 (12):69–??, December 1992. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1992:PPD

[Ano92-30] Anonymous. Protecting purchasers of digital spatial data. Computer Graphics World, 15(5): 104–??, May 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Rd

[Ano92-31] Anonymous. Review. Computer Graphics World, 15(7):97–??, July 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:Re

[Ano92-32] Anonymous. Review. Computer Graphics World, 15(8):79–??, August 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:Rf

[Ano92-33] Anonymous. Review. Computer Graphics World, 15(9):77– ??, September 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Ra

[Ano92-34] Anonymous. Review. Computer Graphics World, 15(10):83– ??, October 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Rb

[Ano92-35] Anonymous. Review. Computer Graphics World, 15(11):81– ??, November 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:Rc

[Ano92-36] Anonymous. Review. Computer Graphics World, 15(12):75– ??, December 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:RMR

[Ano92-37] Anonymous. Review: Modelling and rendering with Alias Sketch. Computer Graphics World, 15(6): 71–??, June 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:SP

[Ano92-38] Anonymous. SIGGRAPH ’92 Preview. Computer Graphics World, 15(7):102–??, July 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:TWd

[Ano92-39] Anonymous. Tech Watch. Computer Graphics World, 15(8):21– ??, August 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:TWe

[Ano92-40] Anonymous. Tech Watch. Computer Graphics World, 15(9):22– ??, September 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:TWa

[Ano92-41] Anonymous. Tech Watch. Computer Graphics World, 15(10):18– ??, October 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:TWb

[Ano92-42] Anonymous. Tech Watch. Computer Graphics World, 15(11):22– ??, November 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:TWc

[Ano92-43] Anonymous. Tech Watch. Computer Graphics World, 15(12):14– ??, December 1992. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1992:MVR

[Ano92-44] Anonymous. The market for virtual reality. Computer Graphics World, 15(5):27–??, May 1992. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1992:MP

[Ano92-45] Anonymous. The “Most Principle”. Computer Graphics World, 15(3):112–??, March 1992. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1993:Aa

[Ano93a] Anonymous. Applications. Computer Graphics World, 16(1):77– ??, January 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ae

[Ano93b] Anonymous. Applications. Computer Graphics World, 16(2):59– ??, February 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Af

[Ano93c] Anonymous. Applications. Computer Graphics World, 16(3):57– ??, March 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ag

[Ano93d] Anonymous. Applications. Computer Graphics World, 16(4):73– ??, April 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ah

[Ano93e] Anonymous. Applications. Computer Graphics World, 16(6):71– ??, June 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ai

[Ano93f] Anonymous. Applications. Computer Graphics World, 16(7):89– ??, July 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Aj

[Ano93g] Anonymous. Applications. Computer Graphics World, 16(8):65– ??, August 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ak

[Ano93h] Anonymous. Applications. Computer Graphics World, 16(9):65– ??, September 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ab

[Ano93i] Anonymous. Applications. Computer Graphics World, 16(10):93– ??, October 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ac

[Ano93j] Anonymous. Applications. Computer Graphics World, 16(11):79– ??, November 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ad

[Ano93k] Anonymous. Applications. Computer Graphics World, 16(12):65– ??, December 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Oa

[Ano93l] Anonymous. Output. Computer Graphics World, 16(1):100– ??, January 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ob

[Ano93m] Anonymous. Output. Computer Graphics World, 16(2):88– ??, February 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Oc

[Ano93n] Anonymous. Output. Computer Graphics World, 16(3):94– ??, March 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Od

[Ano93o] Anonymous. Output. Computer Graphics World, 16(5):72–??, May 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:Oe

[Ano93p] Anonymous. Output. Computer Graphics World, 16(9):96– ??, September 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Pa

[Ano93q] Anonymous. Portfolio. Computer Graphics World, 16(1):72–??, January 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:Pe

[Ano93r] Anonymous. Portfolio. Computer Graphics World, 16(3):44– ??, March 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Pf

[Ano93s] Anonymous. Portfolio. Computer Graphics World, 16(5):34–??, May 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:Pg

[Ano93t] Anonymous. Portfolio. Computer Graphics World, 16(6):62– ??, June 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ph

[Ano93u] Anonymous. Portfolio. Computer Graphics World, 16(8):62–??, August 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:Pi

[Ano93v] Anonymous. Portfolio. Computer Graphics World, 16(9):60– ??, September 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Pb

[Ano93w] Anonymous. Portfolio. Computer Graphics World, 16(10):88– ??, October 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Pc

[Ano93x] Anonymous. Portfolio. Computer Graphics World, 16(11):74– ??, November 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Pd

[Ano93y] Anonymous. Portfolio. Computer Graphics World, 16(12):62– ??, December 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:PGa

[Ano93z] Anonymous. Product Guide. Computer Graphics World, 16(1): 81–??, January 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:PGd

[Ano93-27] Anonymous. Product Guide. Computer Graphics World, 16(2): 65–??, February 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:PGe

[Ano93-28] Anonymous. Product Guide. Computer Graphics World, 16(3): 65–??, March 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:PGf

[Ano93-29] Anonymous. Product Guide. Computer Graphics World, 16(4): 65–??, April 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:PGg

[Ano93-30] Anonymous. Product Guide. Computer Graphics World, 16(5): 59–??, May 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:PGh

[Ano93-31] Anonymous. Product Guide. Computer Graphics World, 16(6): 77–??, June 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:PGi

[Ano93-32] Anonymous. Product Guide. Computer Graphics World, 16(7): 95–??, July 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:PGj

[Ano93-33] Anonymous. Product Guide.

Computer Graphics World, 16(8): 73–??, August 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:PGk

[Ano93-34] Anonymous. Product Guide.

Computer Graphics World, 16(9): 71–??, September 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:PGb

[Ano93-35] Anonymous. Product Guide.

Computer Graphics World, 16 (10):99–??, October 1993. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1993:PGc

[Ano93-36] Anonymous. Product Guide.

Computer Graphics World, 16 (11):85–??, November 1993. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1993:Pj

[Ano93-37] Anonymous. Products. Computer Graphics World, 16(11):93– ??, November 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Rc

[Ano93-38] Anonymous. Review. Computer Graphics World, 16(1):83–??, January 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:Re

[Ano93-39] Anonymous. Review. Computer Graphics World, 16(2):73– ??, February 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Rf

[Ano93-40] Anonymous. Review. Computer Graphics World, 16(3):71– ??, March 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Rg

[Ano93-41] Anonymous. Review. Computer Graphics World, 16(4):81– ??, April 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Rh

[Ano93-42] Anonymous. Review. Computer Graphics World, 16(5):61–??, May 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:Ri

[Ano93-43] Anonymous. Review. Computer Graphics World, 16(6):83– ??, June 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Rj

[Ano93-44] Anonymous. Review. Computer Graphics World, 16(7):101– ??, July 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Rk

[Ano93-45] Anonymous. Review. Computer Graphics World, 16(9):75– ??, September 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Rd

[Ano93-46] Anonymous. Review. Computer Graphics World, 16(10):105–??, October 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Ra

[Ano93-47] Anonymous. Reviews. Computer Graphics World, 16(11):88– ??, November 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:Rb

[Ano93-48] Anonymous. Reviews. Computer Graphics World, 16(12):70– ??, December 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:TTa

[Ano93-49] Anonymous. Tech Talk. Computer Graphics World, 16(7):33–??, July 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:TTb

[Ano93-50] Anonymous. Tech Talk. Computer Graphics World, 16(8):17–??, August 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:TTc

[Ano93-51] Anonymous. Tech Talk. Computer Graphics World, 16(9):13– ??, September 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:TWa

[Ano93-52] Anonymous. Tech Watch. Computer Graphics World, 16(1):14– ??, January 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:TWd

[Ano93-53] Anonymous. Tech Watch. Computer Graphics World, 16(2):14– ??, February 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:TWe

[Ano93-54] Anonymous. Tech Watch. Computer Graphics World, 16(3):18– ??, March 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:TWf

[Ano93-55] Anonymous. Tech Watch. Computer Graphics World, 16(4):16– ??, April 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:TWg

[Ano93-56] Anonymous. Tech Watch. Computer Graphics World, 16(5):10– ??, May 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:TWh

[Ano93-57] Anonymous. Tech Watch. Computer Graphics World, 16(6):21– ??, June 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:TWi

[Ano93-58] Anonymous. Tech Watch. Computer Graphics World, 16(7):16– ??, July 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:TWb

[Ano93-59] Anonymous. Tech Watch. Computer Graphics World, 16(10):12– ??, October 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:TWc

[Ano93-60] Anonymous. Tech Watch. Computer Graphics World, 16(11):16– ??, November 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:T

[Ano93-61] Anonymous. Technology. Computer Graphics World, 16(2):55– ??, February 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1993:QT

[Ano93-62] Anonymous. The Quest for “True” 3D. Computer Graphics World, 16(9):35–??, September 1993. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1993:V

[Ano93-63] Anonymous.  Video.  Com-

puter Graphics World, 16(3):47– ??, March 1993. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Aa

[Ano94a] Anonymous. Applications. Computer Graphics World, 17(1):45– ??, January 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Ab

[Ano94b] Anonymous. Applications. Computer Graphics World, 17(2):57– ??, February 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Ad

[Ano94c] Anonymous. Applications. Computer Graphics World, 17(3):65– ??, March 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Af

[Ano94d] Anonymous. Applications. Computer Graphics World, 17(4):49– ??, April 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Ah

[Ano94e] Anonymous. Applications. Computer Graphics World, 17(5):43– ??, May 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Aj

[Ano94f] Anonymous. Applications. Computer Graphics World, 17(6):55– ??, June 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:MSa

[Ano94g] Anonymous. Model shop. Computer Graphics World, 17(1):51– ??, January 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:MSb

[Ano94h] Anonymous. Model Shop. Computer Graphics World, 17(2):51– ??, February 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:MSc

[Ano94i] Anonymous. Model Shop. Computer Graphics World, 17(4):53– ??, April 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:MSd

[Ano94j] Anonymous. Model Shop. Computer Graphics World, 17(5):49– ??, May 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:MSe

[Ano94k] Anonymous. Model Shop. Computer Graphics World, 17(6):61– ??, June 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Pa

[Ano94l] Anonymous. Portfolio. Computer Graphics World, 17(3):58– ??, March 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Pb

[Ano94m] Anonymous. Portfolio. Computer Graphics World, 17(4):44– ??, April 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Pc

[Ano94n] Anonymous. Portfolio. Computer Graphics World, 17(6):48– ??, June 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:PGa

[Ano94o] Anonymous. Product Guide.

Computer Graphics World, 17(1): 57–??, January 1994. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1994:PGb

[Ano94p] Anonymous. Product Guide.

Computer Graphics World, 17(2): 61–??, February 1994. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1994:PGc

[Ano94q] Anonymous. Product Guide.

Computer Graphics World, 17(3): 71–??, March 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:PGd

[Ano94r] Anonymous. Product Guide.

Computer Graphics World, 17(4): 59–??, April 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:PGe

[Ano94s] Anonymous. Product Guide. Computer Graphics World, 17(5): 53–??, May 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:PGf

[Ano94t] Anonymous. Product Guide. Computer Graphics World, 17(6): 67–??, June 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Ra

[Ano94u] Anonymous. Reviews. Computer Graphics World, 17(1):60–??, January 1994. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1994:Rb

[Ano94v] Anonymous. Reviews. Computer Graphics World, 17(2):66– ??, February 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Rc

[Ano94w] Anonymous. Reviews. Computer Graphics World, 17(3):75– ??, March 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Rd

[Ano94x] Anonymous. Reviews. Computer Graphics World, 17(4):63– ??, April 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:Re

[Ano94y] Anonymous. Reviews. Computer Graphics World, 17(5):58–??, May 1994. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1994:Rf

[Ano94z] Anonymous. Reviews. Computer Graphics World, 17(6):71– ??, June 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:TWa

[Ano94-27] Anonymous. Tech Watch. Computer Graphics World, 17(1):12– ??, January 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:TWb

[Ano94-28] Anonymous. Tech Watch. Computer Graphics World, 17(2):14– ??, February 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1994:TWc

[Ano94-29] Anonymous. Tech Watch. Computer Graphics World, 17(6):13– ??, June 1994. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:CCCa

[Ano95a] Anonymous. CAD/CAM Connection — Broad support for the STEP data-exchange standard. Computer Graphics World, 18(7):19–??, July 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:CCCb

[Ano95b] Anonymous. CAD/CAM Connection — Emphasizing the value of innovative design engineers. Computer Graphics World, 18(12):19– ??, December 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:CCCc

[Ano95c] Anonymous. CAD/CAM Connection — Is “fancy” rendering really necessary? Computer Graphics World, 18(10):19–??, October 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:CCCd

[Ano95d] Anonymous. CAD/CAM Connection — Some traps and pitfalls to watch out for. Computer Graphics World, 18(8):25–??, August 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:CCCe

[Ano95e] Anonymous. CAD/CAM Connection — The challenges of 3D digitizing. Computer Graphics World, 18(11):21–??, November 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:CCCf

[Ano95f] Anonymous. CAD/CAM Connection — The language of CAD. Computer Graphics World, 18(6): 21–??, June 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:CCCg

[Ano95g] Anonymous. CAD/CAM Connection — The need for specialized CAID software. Computer Graphics World, 18(9):19–??, September 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:DWR

[Ano95h] Anonymous. Digital Workshop — Recycle an old image using Xaos Tools’ Pennello. Computer Graphics World, 18(4):79– ??, April 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:DWS

[Ano95i] Anonymous. Digital Workshop — Simulating sunlight. Computer Graphics World, 18(6):70– ??, June 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:MSCa

[Ano95j] Anonymous. Model shop — creating a robot in Hash Inc’s Animation Master. Computer Graphics World, 18(4):71–??, April 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:MSCb

[Ano95k] Anonymous. Model Shop — Creating cool effects with KPT Gradient Designer. Computer Graphics World, 18(12):68–??, December 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:MSD

[Ano95l] Anonymous. Model Shop — Designing a watch with Alias Studio. Computer Graphics World, 18(5): 65–??, May 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:MSMa

[Ano95m] Anonymous. Model Shop — Modeling a believable tree. Computer Graphics World, 18(8):83–??, August 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:MSMb

[Ano95n] Anonymous. Model Shop — Modeling a dinosaur with metaballs. Computer Graphics World, 18(7): 79–??, July 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:MSMc

[Ano95o] Anonymous. Model Shop — Modeling a picture-perfect fly. Computer Graphics World, 18(11):67– ??, November 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:MSMd

[Ano95p] Anonymous. Model Shop — Modeling a slide with PTC’s Pro/ Engineer. Computer Graphics World, 18(6):65–??, June 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:MSS

[Ano95q] Anonymous. Model Shop — Some tips for designing a 3D database for simulation applications. Computer Graphics World, 18(10):67– ??, October 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:MSB

[Ano95r] Anonymous. Model Shop — The benefits of shaded graphics modeling. Computer Graphics World, 18(9):51–??, September 1995. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1995:PIR

[Ano95s] Anonymous. Portfolio — 3D illustrator Rob Magiera. Computer Graphics World, 18(5):54–??, May 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:PSPa

[Ano95t] Anonymous. Portfolio — A sneak peek at this year’s SIGGRAPH Electronic Theatre Competition. Computer Graphics World, 18(8): 72–??, August 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:PRD

[Ano95u] Anonymous. Portfolio — Roy Dream’s Modern Masters of 3D. Computer Graphics World, 18 (12):56–??, December 1995. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1995:PRP

[Ano95v] Anonymous. Portfolio — Russian photographer Vladimir Pcholkin. Computer Graphics World, 18(4): 52–??, April 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:PSWa

[Ano95w] Anonymous. Portfolio — Some winners of Intergraph’s Golden Mouse Awards Competition. Computer Graphics World, 18(6): 52–??, June 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:PSWb

[Ano95x] Anonymous. Portfolio — Some works from Disney animator Lee Lanier. Computer Graphics World, 18(10):56–??, October 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:PSM

[Ano95y] Anonymous. Portfolio — Summer movie showcase. Computer Graphics World, 18(7):70–??, July 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:PTS

[Ano95z] Anonymous. Portfolio — Toy Story’s believable hair. Computer Graphics World, 18(11):58– ??, November 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:PUC

[Ano95-27] Anonymous. Portfolio — Ultrarealistic CG humans. Computer Graphics World, 18(9):42– ??, September 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:PGS

[Ano95-28] Anonymous. Product Guide — 16-bit sound cards. Computer Graphics World, 18(5):74–??, May 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:PGC

[Ano95-29] Anonymous. Product Guide — 21-inch color monitors. Computer Graphics World, 18(7):92–??, July 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:PGM

[Ano95-30] Anonymous. Product Guide — 3D model libraries. Computer Graphics World, 18(12):80–??, December 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:PGD

[Ano95-31] Anonymous. Product Guide — Digital cameras. Computer Graphics World, 18(6):78–??, June 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:PGH

[Ano95-32] Anonymous. Product Guide — Head-mounted displays. Computer Graphics World, 18(8):101– ??, August 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:PGL

[Ano95-33] Anonymous. Product Guide — LCD projection panels. Computer Graphics World, 18(11):70– ??, November 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:PGV

[Ano95-34] Anonymous. Product Guide — Video/graphics accelerators. Computer Graphics World, 18(9): 67–??, September 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:PAP

[Ano95-35] Anonymous. Products — Autofact ’95 Preview. Computer Graphics World, 18(11):78–??, November 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:PSPb

[Ano95-36] Anonymous. Products — SIG-GRAPH ’95 Preview. Computer Graphics World, 18(8):105– ??, August 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:PSN

[Ano95-37] Anonymous. Products — Special NAB product preview. Computer Graphics World, 18(4):92– ??, April 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:RET

[Ano95-38] Anonymous. Reviews — 3D/Eye’s TriSpectives. Linker Systems’ Animation Stand. Specular’s Infini-D V3.0. Computer Graphics World, 18(11):73–??, November 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:RBS

[Ano95-39] Anonymous. Reviews — Bentley Systems’ MicroStation modeler. Positron Publishing’s MeshPaint 3D. Strata’s Vision 3d version 4.0. Computer Graphics World, 18(6): 73–??, June 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:RCT

[Ano95-40] Anonymous. Reviews — Caligan’s trueSpace 2.0. HSC’s Live Picture 2.0,. Autodesk’s Animator Studio. Computer Graphics World, 18(7): 85–??, July 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:RFD

[Ano95-41] Anonymous. Reviews — Fractal Design’s Poser. Matra Datavision’s Strim V6.0. Dynamic Realities’ Impact plug-in for Light-Wave. Computer Graphics World, 18(8):89–??, August 1995. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1995:RIC

[Ano95-42] Anonymous. Reviews — In:Sync Corp’s Razor Pro desktop video editor Digimation’s LenZFX 1 3, on optical-effects generator for 3D Studio. Corel’s Photo-Paint 5 Plus image editing and paint program. Computer Graphics World, 18(4): 87–??, April 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:RMS

[Ano95-43] Anonymous. Reviews — Microcadam’s solid models. AnimaTek’s World Builder. the MGA Millennium from Matrox. Computer Graphics World, 18(10):73– ??, October 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:RNL

[Ano95-44] Anonymous. Reviews — NewTek’s LightWave 3D V4 0. Knowledge Revolution’s Working Model V3.0. Computer Graphics World, 18 (12):75–??, December 1995. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1995:RPT

[Ano95-45] Anonymous. Reviews — Parametric Technology’s Pro/JR. Virtuality’s Elysium. VR development system. form. Z2.7 Render-Zone from auto.des.sys. Computer Graphics World, 18(9):57– ??, September 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:RVG

[Ano95-46] Anonymous. Reviews — the Volis Group’s PixelPutty Solo V1.2. Data Translation’s Media 100 V2.0. BeInfinite’s InfiniteFX. Computer Graphics World, 18(5): 69–??, May 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:SCCa

[Ano95-47] Anonymous. Special CAD/CAM supplement: 3D digitizing for engineering. Computer Graphics World, 18(3):S1–??, March 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:SCCb

[Ano95-48] Anonymous. Special CAD/CAM Supplement: Designing With Low-Cost 3D CAD Software. Computer Graphics World, 18(5): 1–??, May 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:SNPb

[Ano95-49] Anonymous. Spotlight: News and Products. Computer Graphics World, 18(9):8–??, September 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:SNPa

[Ano95-50] Anonymous. Spotlight: News and Products. Computer Graphics World, 18(11):8–??, November 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:TWC

[Ano95-51] Anonymous. Tech watch — A computer connection to the mind. A roving virtual reality van. VR glove uses conductive sense technology. Computer Graphics World, 18(5):17–??, May 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:TWI

[Ano95-52] Anonymous. Tech Watch — Intel’s Indeo Video Interactive. Pacific Bell’s enhanced Media Park. ADAMS/ Car for advanced vehicle simulation. Computer Graphics World, 18(11):17–??, November 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:TWV

[Ano95-53] Anonymous. Tech Watch — Virtual playground for seriously ill children. Computer Graphics World, 18(7):13–??, July 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:UFVa

[Ano95-54] Anonymous. User Focus — A virtual tour of France’s Lascaux Cave. CAD gives handcrafted look to ceramic dinnerware. Computer Graphics World, 18(10):61– ??, October 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:UFA

[Ano95-55] Anonymous. User Focus — Architects use familiar tools in a new way Canadian television’s Re-Boot cartoon. Computer Graphics World, 18(11):61–??, November 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:UFCb

[Ano95-56] Anonymous. User Focus — CAD and CFD contribute to America’s Cup win. Live theater uses computer generated scenery. Computer Graphics World, 18(7):75– ??, July 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:UFCa

[Ano95-57] Anonymous. User Focus — Computer analyses help engineers optimize heart-valve design. Computer Graphics World, 18(9):47– ??, September 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:UFD

[Ano95-58] Anonymous. User Focus — Designing snowboards with Auto-CAD and Pro/E Visualization aids infertility research. Computer Graphics World, 18(5):59–??, May 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:UFH

[Ano95-59] Anonymous. User Focus — Homer and Bart Simpson go 3D. Animations highlight downtown Las Vegas. Computer Graphics World, 18 (12):63–??, December 1995. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1995:UFI

[Ano95-60] Anonymous. User Focus — Interactive walkthrough helps San Diego clinch ’96. Republican National Convention bid. Dynamic animations rev up Detroit Pistons fans. Computer Graphics World, 18(4):58–??, April 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:UFP

[Ano95-61] Anonymous. User Focus — Planning for the ’96 Olympics with 3D CAD. CAD aids in car-collision analyses. Computer Graphics World, 18(6):58–??, June 1995. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1995:UFVb

[Ano95-62] Anonymous. User Focus — VR effects in Lawnmower Man II. Photorealistic renderings make the grade for proposed student center. Computer Graphics World, 18(8): 77–??, August 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1995:USA

[Ano95-63] Anonymous. User snapshots — animations help sell cereal. creating cartoon characters for new CD-ROM game. 3D illustrations brighten pages of Money Magazine for Kids. analyzing the structure of the Golden Gate Bridge.

Computer Graphics World, 18(4): 60–??, April 1995. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:CCCf

[Ano96a] Anonymous. CAD/CAM connection — defining new CAD standards. Computer Graphics World, 19(9):21–??, September 1996. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1996:CCCg

[Ano96b] Anonymous. CAD/CAM connection — focusing technology on business benefits. Computer Graphics World, 19(10):23– ??, October 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:CCCb

[Ano96c] Anonymous. CAD/CAM connection — more CAD traps and pitfalls. Computer Graphics World, 19(3):19–??, March 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:CCCd

[Ano96d] Anonymous. CAD/CAM connection — RenderWare vs. OpenGL. Computer Graphics World, 19(7): 21–??, July 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:CCCc

[Ano96e] Anonymous. CAD/CAM connection — the real value of 3D rendering to architects. Computer Graphics World, 19(6):19– ??, June 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:CCCe

[Ano96f] Anonymous. CAD/CAM connection — Unix workstations will prosper despite Windows NT’s success. Computer Graphics World, 19(8):21–??, August 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:CCCh

[Ano96g] Anonymous. CAD/CAM connection — what does VRML have to offer design engineers? Computer Graphics World, 19(12):22– ??, December 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:CCCa

[Ano96h] Anonymous. CAD/CAM connection — Windows: a threat to Unix? Computer Graphics World, 19(2):25–??, February 1996. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1996:CCS

[Ano96i] Anonymous. CAD/CAM special report: Workgroup vs. enterprisewide PDM — should prospective buyers of product data management systems purchase and install enterprise-wide systems or local (workgroup) systems? what are the benefits and limitations of each? Computer Graphics World, 19(9):S1–??, September 1996. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1996:DSCa

[Ano96j] Anonymous. Digital studio — 3D chip powers arcade game. low-end 3D chips proliferate. free 3D models on the Web. Computer Graphics World, 19(4):17– ??, April 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:DSCb

[Ano96k] Anonymous. Digital studio — “CG” movies and their Web sites. James and the Giant Peach. Amazing Studio’s Heart of Darkness. Computer Graphics World, 19(8):17–??, August 1996. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1996:DSQ

[Ano96l] Anonymous. Digital studio — quake the 3D sequel to Doom. experiencing the Vatican’s “Raffaello Rooms”. A virtual trip to Mars. Computer Graphics World, 19(10):19–??, October 1996. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1996:DSU

[Ano96m] Anonymous. Digital studio — the ultimate racing game. motioncaptured characters. creating digital life. Computer Graphics World, 19(12):19–??, December 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:HPC

[Ano96n] Anonymous. Hottest products of 1995 — CGW’s 2nd Annual Editors’ Choice Awards honor the most innovative products introduced and shipped in 1995. Computer Graphics World, 19(1):40– ??, January 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:MSA

[Ano96o] Anonymous. Model shop — animating displacement maps. Computer Graphics World, 19(11):85– ??, November 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:MSCb

[Ano96p] Anonymous. Model shop — create a walking alien with Character Studio. Computer Graphics World, 19(8):87–??, August 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:MSCa

[Ano96q] Anonymous. Model shop — creating a convincing composite. Computer Graphics World, 19(7):67– ??, July 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:MSD

[Ano96r] Anonymous. Model shop — dealing with details. Computer Graphics World, 19(5):73–??, May 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:MSMd

[Ano96s] Anonymous. Model shop — modeling an airplane with Hash’s Animation Master. Computer Graphics World, 19(10):65–??, October 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:MSMb

[Ano96t] Anonymous. Model shop — modeling faces that exude personality when animated. Computer Graphics World, 19(2):58–??, February 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:MSMa

[Ano96u] Anonymous. Model shop — modeling the human face with Alias’ PowerAnimator. Computer Graphics World, 19(1):66–??, January 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:MSSb

[Ano96v] Anonymous. Model shop — some modeling tips for game developers. Computer Graphics World, 19(4): 59–??, April 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:MSSa

[Ano96w] Anonymous. Model shop — some modeling tips for the Boolean-shy. Computer Graphics World, 19(3): 69–??, March 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:MSSc

[Ano96x] Anonymous. Model shop — Sony Pictures Imageworks creates CG butterflies for feature film The Craft. Computer Graphics World, 19(6):79–??, June 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:MSMc

[Ano96y] Anonymous. Model shop — the modular approach to modeling. Computer Graphics World, 19(9): 75–??, September 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PSW

[Ano96z] Anonymous. Portfolio — A sampling of works from this year’s SIGGRAPH Electronic Theatre. Computer Graphics World, 19(8): 74–??, August 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:PAA

[Ano96-27] Anonymous. Portfolio — Amblin animators give Balto a handdrawn look. Computer Graphics World, 19(3):62–??, March 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PAS

[Ano96-28] Anonymous. Portfolio — Animator “sculpts” Iron Man poster. Computer Graphics World, 19(1): 56–??, January 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PCM

[Ano96-29] Anonymous. Portfolio — CG magic in Disney’s Hunchback. Computer Graphics World, 19(7): 60–??, July 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:PGCb

[Ano96-30] Anonymous. Portfolio — Graphisoft contest sparks architecture students’ creativity. Computer Graphics World, 19(9):68–??, September 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:PIG

[Ano96-31] Anonymous. Portfolio — Intergraph’s Golden Mouse Award winners. Computer Graphics World, 19(6):66–??, June 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PLT

[Ano96-32] Anonymous. Portfolio — lowerend tools produce high-end results. Computer Graphics World, 19(5):62–??, May 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PSR

[Ano96-33] Anonymous. Portfolio — stunning room designs. Computer Graphics World, 19(12):58–??, December 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PAI

[Ano96-34] Anonymous. Portfolio — the art of industrial design. Computer Graphics World, 19(10):54– ??, October 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:PIH

[Ano96-35] Anonymous. Portfolio — the imagery of Henk Dawson. Computer Graphics World, 19(11):74– ??, November 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:PWG

[Ano96-36] Anonymous. Portfolio — the work of graphic artist Richard Bucci. Computer Graphics World, 19(4): 48–??, April 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:PGCc

[Ano96-37] Anonymous. Product guide — CD recorders and software. Computer Graphics World, 19(12):71– ??, December 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:PGCa

[Ano96-38] Anonymous. Product guide — CRT projectors. Computer Graphics World, 19(6):90–??, June 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PGD

[Ano96-39] Anonymous. Product guide — direct-imaging plotters. Computer Graphics World, 19(3):76– ??, March 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:PGH

[Ano96-40] Anonymous. Product guide — high-end scan converters. Computer Graphics World, 19(2):66– ??, February 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:PGLb

[Ano96-41] Anonymous. Product guide — LED plotters. Computer Graphics World, 19(10):74–??, October 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PGLa

[Ano96-42] Anonymous. Product guide — low-cost 3D accelerators. Computer Graphics World, 19(9):86– ??, September 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:PGPa

[Ano96-43] Anonymous. Product guide — PC-based video-capture cards. Computer Graphics World, 19(1): 78–??, January 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PGPb

[Ano96-44] Anonymous. Product guide — presentation monitors. Computer Graphics World, 19(5):82–??, May 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PGR

[Ano96-45] Anonymous. Product guide — rapid prototyping systems. Computer Graphics World, 19(11): 100–??, November 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PGU

[Ano96-46] Anonymous. Product guide — Under-USD 10,000 desktop film recorders. Computer Graphics World, 19(7):76–??, July 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PNP

[Ano96-47] Anonymous. Products — NAB product preview. Computer Graphics World, 19(4):71–??, April 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:PSP

[Ano96-48] Anonymous. Products — SIG-GRAPH ’96 product preview. Computer Graphics World, 19(8): 103–??, August 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:RQR

[Ano96-49] Anonymous. Renault’s quest for realism — Renault’s design and visualization expertise shines in Les Citadines, the latest in a string of videos the French automaker has produced to not only showcase its graphics expertise but also to play a valuable role in its design process. Computer Graphics World, 19(3):25–??, March 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:RAA

[Ano96-50] Anonymous. Reviews — Adobe’s After Effects V3.0. Ray Dream’s Studio Avid’s beyond reality. Computer Graphics World, 19(3): 71–??, March 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:RAW

[Ano96-51] Anonymous. Reviews — Alias: Wavefront’s PowerAnimator V7. Speed Razor Mach 111 from in:sync. Computer Graphics World, 19(5):77–??, May 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:RCC

[Ano96-52] Anonymous. Reviews — Corel CAD. HiJaak Pro 4.0. Computer Graphics World, 19(12):67– ??, December 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:RIS

[Ano96-53] Anonymous. Reviews — Intergraph’s Solid Edge. Autodesk’s Mechanical Desktop. Macromedia’s Director 5.0. Computer Graphics World, 19(9):81–??, September 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:RKS

[Ano96-54] Anonymous. Reviews — Kinetix’s 3D Studio Max. Bryce 2.0 from MetaTools. Computer Graphics World, 19(7):71–??, July 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:RMM

[Ano96-55] Anonymous. Reviews — MetaReyes’ Metaballs V2.01. Strata’s Studio-Pro Blitz V1.75. Positran Publishing’s Dynamic Motion Module. Computer Graphics World, 19 (1):73–??, January 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:RMS

[Ano96-56] Anonymous. Reviews — Microsoft’s Softimage 3D for NT. Fractal’s Painter 4. MetaTools’ Kai’s Power Tools V3.0. Computer Graphics World, 19(4):65– ??, April 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:RNL

[Ano96-57] Anonymous. Reviews — NewTek’s LightWave 5.0. Positron Publishing’s MeshPaint 3D V1.5. Computer Graphics World, 19(10):69– ??, October 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:RSS

[Ano96-58] Anonymous. Reviews — Solid-Works 95 solid modeler. Macromedia’s Extreme 3D. Hash Inc’s Animation Master V4.0. Computer Graphics World, 19(6):85– ??, June 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:RSP

[Ano96-59] Anonymous. Reviews — Solid-Works 96. PTC’s Pro/E rel 17.0. ElectricImage V2.7. Computer Graphics World, 19(11):91–??, November 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:RUP

[Ano96-60] Anonymous. Reviews — Un-Real pictures/Kinetix’s Character Studio. Pixar’s RenderMan artists tools. Computer Graphics World, 19(8):93–??, August 1996. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1996:RVS

[Ano96-61] Anonymous. Reviews — Visual Software’s Visual Reality 2.0. Computer Graphics World, 19(2): 62–??, February 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:SNP

[Ano96-62] Anonymous. Spotlight: News and products. Computer Graphics World, 19(4):8–??, April 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:SSM

[Ano96-63] Anonymous. Sun, SGI, and Macromedia to extend Java. Computer Graphics World, 19(1):8, January 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:TWS

[Ano96-64] Anonymous. Tech watch — 3D scenes from visual databases. NASA eases way through huge databases. Vote America sorts out presidential campaign issues and candidates. Computer Graphics World, 19(3):15–??, March 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:TWVb

[Ano96-65] Anonymous. Tech watch — A virtual emergency room. Computer Graphics World, 19(9):17– ??, September 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:TWVc

[Ano96-66] Anonymous. Tech watch — A virtual retina display. healing with holograms. group VR environment. Computer Graphics World, 19(11):19–??, November 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:TWM

[Ano96-67] Anonymous. Tech watch — mapping wound-healing progress. reality meets virtual reality. telemedicine debuts at Summer Olympics. Computer Graphics World, 19(7):17–??, July 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:TWN

[Ano96-68] Anonymous. Tech watch — new system for 3D motion capture. time multiplexing enables “perfect” 3D display. Computer Graphics World, 19(1):19–??, January 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:TWVa

[Ano96-69] Anonymous. Tech watch — VR simulates surgery. forum addresses multimedia options. digital archive of the earth. Computer Graphics World, 19(5):17–??, May 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:UFMa

[Ano96-70] Anonymous. User focus — 3D models for medical education. Computer Graphics World, 19(2): 55–??, February 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:UFB

[Ano96-71] Anonymous. User focus — A bold opener for MTV features on animated fetus. Computer Graphics World, 19(11):79–??, November 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:UFNb

[Ano96-72] Anonymous. User focus — A new ride film goes underground. Computer Graphics World, 19(12):61– ??, December 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:UFA

[Ano96-73] Anonymous. User focus — AEC firm wins design bid with help of interactive simulation. CD-ROM re-creates one of China’s lost treasures. Computer Graphics World, 19(6):73–??, June 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:UFGa

[Ano96-74] Anonymous. User focus — going back in time in Las Vegas for Casino. AEC company relies on realtime virtual reality. Computer Graphics World, 19(1):61–??, January 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:UFMb

[Ano96-75] Anonymous. User focus — Magnet Pictures re-creates action figure for toy commercial. animators re-create Arizona fire. Computer Graphics World, 19(3):65– ??, March 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:UFNa

[Ano96-76] Anonymous. User focus — new interface puts filmmakers inside scientific wonderland. Computer Graphics World, 19(8):81–??, August 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:UFRb

[Ano96-77] Anonymous. User focus — redesigning SyQuest with CAD. exploring Olympic locale with simulated flythrough. Computer Graphics World, 19(9):71– ??, September 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:UFRa

[Ano96-78] Anonymous. User focus — reverse engineering elevates science of saddle making. Computer Graphics World, 19(7):63–??, July 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:UFS

[Ano96-79] Anonymous. User focus — simulating a facelift for the Royal Albert Hall 3D simulation key evidence in leaky-roof lawsuit. Computer Graphics World, 19(5):69– ??, May 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1996:UFC

[Ano96-80] Anonymous. User focus — the challenge of creating a new kind of VR experience. Computer Graphics World, 19(10):59–??, October 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:UFGb

[Ano96-81] Anonymous. User focus — the “Great American Virtual Home”. CAD plays role in design of Dark-Star military aircraft. Computer Graphics World, 19(4):53– ??, April 1996. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:CCCb

[Ano97a] Anonymous. CAD/CAM connection — architects need to think like filmmakers. Computer Graphics World, 20(4):19– ??, April 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:CCCi

[Ano97b] Anonymous. CAD/CAM connection — common-sense rules for CAD. Computer Graphics World, 20(11):21–??, November 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:CCCe

[Ano97c] Anonymous. CAD/CAM connection — displaying large assemblies. Computer Graphics World, 20(7):19–??, July 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:CCCf

[Ano97d] Anonymous. CAD/CAM connection — displaying large assemblies. Computer Graphics World, 20(8):23–??, August 1997. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1997:CCCg

[Ano97e] Anonymous. CAD/CAM connection — feature-based virtual prototyping. Computer Graphics World, 20(9):23–??, September 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:CCCh

[Ano97f] Anonymous. CAD/CAM connection — getting the most from models. Computer Graphics World, 20(10):23–??, October 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:CCCc

[Ano97g] Anonymous. CAD/CAM connection — solid modeling for industrial design. Computer Graphics World, 20(5):19–??, May 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:CCCj

[Ano97h] Anonymous. CAD/CAM connection — the changing landscape of mechanical CAD. Computer Graphics World, 20(12):21– ??, December 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:CCCd

[Ano97i] Anonymous. CAD/CAM connection — the future of 3D graphics boards. Computer Graphics World, 20(6):15–??, June 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:CCCa

[Ano97j] Anonymous. CAD/CAM connection — the Web promises to bring sweeping changes. Computer Graphics World, 20(2):15– ??, February 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:DSIb

[Ano97k] Anonymous. Digital studio — an intergalactic adventure. it’s back to Zork. Computer Graphics World, 20(10):17–??, October 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:DSCb

[Ano97l] Anonymous. Digital studio — constructing Lego Island. Dilbert does 3D. Computer Graphics World, 20 (12):17–??, December 1997. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1997:DSCa

[Ano97m] Anonymous. Digital studio — creating an LBE experience. A unique real-time character. Computer Graphics World, 20(6):11– ??, June 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:DSIa

[Ano97n] Anonymous. Digital studio — the invisible effects of “roar”. A crosscultural edutainment title. Computer Graphics World, 20(8):17– ??, August 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:DSM

[Ano97o] Anonymous. Digital studio — the magic behind The Relic; a new game from Psygnosis. Computer Graphics World, 20(4):13– ??, April 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:HPY

[Ano97p] Anonymous. The hottest products of the year — it’s time once again to recognize the most impressive products of the year. the winners of this year’s CGW innovation awards are a select group of 25 new products that stood out from the crowd. these are products that are truly pushing the leading edge of computer. Computer Graphics World, 20(1):46–??, January 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:MSCb

[Ano97q] Anonymous. Model shop — create your own 3D digitizer. Computer Graphics World, 20(4):75– ??, April 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:MSCa

[Ano97r] Anonymous. Model shop — creating a lost temple with 3D Studio Max. Computer Graphics World, 20(2):61–??, February 1997. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1997:MSD

[Ano97s] Anonymous. Model shop — designing a personal pager. Computer Graphics World, 20(3):65– ??, March 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:PDC

[Ano97t] Anonymous. Portfolio — A digital cell reconstruction. Computer Graphics World, 20(9):74– ??, September 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:PMM

[Ano97u] Anonymous. Portfolio — A modeler with a mission. Computer Graphics World, 20(5):58–??, May 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:PWS

[Ano97v] Anonymous. Portfolio — A winning student animation. Computer Graphics World, 20(7):58–??, July 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:PA

[Ano97w] Anonymous. Portfolio — Anastasia. Computer Graphics World, 20 (12):60–??, December 1997. CO-DEN CGWODH. ISSN 0271-4159. Anonymous:1997:PCLb

[Ano97x] Anonymous. Portfolio — creating landscapes. Computer Graphics World, 20(6):74–??, June 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:PHS

[Ano97y] Anonymous. Portfolio — hot summer movies. Computer Graphics World, 20(8):88–??, August 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:PST

[Ano97z] Anonymous. Portfolio — Starship Titanic. Computer Graphics World, 20(10):74–??, October 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:PAA

[Ano97-27] Anonymous. Portfolio — the architect as artist. Computer Graphics World, 20(1):56–??, January 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:PCLa

[Ano97-28] Anonymous. Portfolio — The City of Lost Children. Computer Graphics World, 20(4):68– ??, April 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:PIV

[Ano97-29] Anonymous. Portfolio — the imagery of Versailles. Computer Graphics World, 20(2):52– ??, February 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:PVH

[Ano97-30] Anonymous. Portfolio — visible humans. Computer Graphics World, 20(3):56–??, March 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:PZG

[Ano97-31] Anonymous. Portfolio — Zork Grand Inquisitor. Computer Graphics World, 20(11):78–??, November 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:PGR

[Ano97-32] Anonymous. Product guide — removable disk drives. Computer Graphics World, 20(2):71– ??, February 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:PGWa

[Ano97-33] Anonymous. Product guide — wide-format ink jet plotters. Computer Graphics World, 20(3):77– ??, March 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:PGWb

[Ano97-34] Anonymous. Product guide — Windows-based solid modelers. Computer Graphics World, 20(5): 66–??, May 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:QWE

[Ano97-35] Anonymous. Q and A: Webenabled engineering. Computer Graphics World, 20(11):71–??, November 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:RSVb

[Ano97-36] Anonymous. Reviews — 3D Studio Viz. Ray Dream 5.0. Computer Graphics World, 20(10):81– ??, October 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:RCF

[Ano97-37] Anonymous. Reviews — Cadkey 97/FastSOLID. SolidDesigner 5.1. Cinema 4D. Computer Graphics World, 20(11):85–??, November 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:RFP

[Ano97-38] Anonymous. Reviews — Fractal’s Poser 2.0; DeBabelizer Pro. Computer Graphics World, 20(4): 83–??, April 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:RIA

[Ano97-39] Anonymous. Reviews — I-DEAS Artisan Series. Strata StudioPro 2.0. Computer Graphics World, 20(6):81–??, June 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:RIS

[Ano97-40] Anonymous. Reviews — Intergraph’s Solid Edge 3.0. Caligari’s trueSpace3. Computer Graphics World, 20(7):65–??, July 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:RRO

[Ano97-41] Anonymous. Reviews — Rhino 3D offers NURBS; MicroScribe digitizer SolidView 2.0. Computer Graphics World, 20(3):71– ??, March 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:RSN

[Ano97-42] Anonymous. Reviews — Second Nature’s Hypermatter. PowerAnimator 8.0. Computer Graphics World, 20(8):99–??, August 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:RSE

[Ano97-43] Anonymous. Reviews — Side Effects’ Houdini. Computer Graphics World, 20(2):67–??, February 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:RSVa

[Ano97-44] Anonymous. Reviews — Softimage 3D V3.51. Speed Razor 3.5. Computer Graphics World, 20(1): 61–??, January 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:RSA

[Ano97-45] Anonymous. Reviews — Solid-Works 97; Allplan FT; Extreme 3D 2.0. Computer Graphics World, 20(5):61–??, May 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:RVV

[Ano97-46] Anonymous. Reviews — Vellum 3D version 3.0. AutoCAD rel. 14. Computer Graphics World, 20(9): 81–??, September 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:RWB

[Ano97-47] Anonymous. Reviews — World Builder 2.0. Infini-D 4.1. Computer Graphics World, 20(12):69– ??, December 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:SNPa

[Ano97-48] Anonymous. Spotlight: News and products. Computer Graphics World, 20(8):8–??, August 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:SNPb

[Ano97-49] Anonymous. Spotlight: News and products. Computer Graphics World, 20(12):8–??, December 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:TWC

[Ano97-50] Anonymous. Tech watch — A Canadian odyssey. CG-enhanced brain research. A new semiconductor design. high-end rendering option. Computer Graphics World, 20(11):17–??, November 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:TWF

[Ano97-51] Anonymous. Tech watch — A floating display. real-time motion capture. designing a bridge. A virtual lab. Computer Graphics World, 20(7):13–??, July 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:TWU

[Ano97-52] Anonymous. Tech watch — A unique collaborative environment for scientists. A virtual trade show. Computer Graphics World, 20(1):13–??, January 1997. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1997:TWP

[Ano97-53] Anonymous. Tech watch — pattern-recognition system. piecing together history. 3D semiconductor simulation. multi-threaded architecture. Computer Graphics World, 20(9):15–??, September 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:TWS

[Ano97-54] Anonymous. Tech watch — Stanford researchers use lasers to create color in a cube; developing virtualized reality at Carnegie Mellon; chemists create teenytiny monitors; high-speed video on the Internet. Computer Graphics World, 20(3):19–??, March 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:TWW

[Ano97-55] Anonymous. Tech watch — the world’s tiniest display; a new VR installation; motion capture for medicine; reproducing 3D surfaces. Computer Graphics World, 20(5):15–??, May 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:UFA

[Ano97-56] Anonymous. User focus — animating biological defenses. Computer Graphics World, 20(9):77– ??, September 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:UFB

[Ano97-57] Anonymous. User focus — Babylon 5’s last hurrah. Computer Graphics World, 20(12):65– ??, December 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:UFC

[Ano97-58] Anonymous. User focus — Chrysler embraces VR. Computer Graphics World, 20(7):61–??, July 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:UFF

[Ano97-59] Anonymous. User focus — FEA plays a role in medical imaging. Computer Graphics World, 20(1): 59–??, January 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:UFJ

[Ano97-60] Anonymous. User focus — Japanese music video features CG characters. Computer Graphics World, 20(2):55–??, February 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:UFSa

[Ano97-61] Anonymous. User focus — scientists employ a unique visualization display. the stage itself comes alive in this theatrical production. Computer Graphics World, 20(3): 59–??, March 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:UFSb

[Ano97-62] Anonymous. User focus — simulating emergency medicine. Computer Graphics World, 20(8):95– ??, August 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:UFSc

[Ano97-63] Anonymous. User focus — surgeon gets 3D insight. Computer Graphics World, 20(10):77– ??, October 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1997:UFH

[Ano97-64] Anonymous. User focus — the house that Wright (could have) built. Computer Graphics World, 20(11):81–??, November 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:UFT

[Ano97-65] Anonymous. User focus — training simulation on a PC. Computer Graphics World, 20(4):71– ??, April 1997. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:CIA

[Ano98a] Anonymous. 1997 CGW innovation awards. Computer Graphics World, 21(1):36–??, January 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:APC

[Ano98b] Anonymous. Antz: Proprietary character-animation software helped launch this critter invasion. Computer Graphics World, 21(7):61–??, July 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:BSF

[Ano98c] Anonymous. Bingo!: A short film puts new animation software to the test. Computer Graphics World, 21(7):33–??, July 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:BLT

[Ano98d] Anonymous. A bug’s life: The team behind Toy Story collaborates on a “buggy” animation. Computer Graphics World, 21(7): 56–??, July 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:CCCa

[Ano98e] Anonymous. CAD/CAM — connection — digital manufacturing. Computer Graphics World, 21(3): 23–??, March 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:CCCc

[Ano98f] Anonymous. CAD/CAM — connection. kernel wars. Computer Graphics World, 21(6):19– ??, June 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:CCCb

[Ano98g] Anonymous. CAD/CAM — connection: Making CAD/CAM, CAE easier. Computer Graphics World, 21(4):21–??, April 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:CSP

[Ano98h] Anonymous. Collaborative 3D simplifies plant design. Computer Graphics World, 21(9):44– ??, September 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:CSC

[Ano98i] Anonymous. Connection — sharing CAD models on the Web. Computer Graphics World, 21(5): 21–??, May 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:DPE

[Ano98j] Anonymous. DARPA pushes for an engineering internet. Computer Graphics World, 21(9):42– ??, September 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:DER

[Ano98k] Anonymous. Departments — events and resources. Computer Graphics World, 21(10):68– ??, October 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:DIS

[Ano98l] Anonymous. Departments — insight — space bananas. Computer Graphics World, 21(10):76– ??, October 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:DMO

[Ano98m] Anonymous. Departments — market outlook — A hungry animation market. Computer Graphics World, 21(10):13–??, October 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:DPU

[Ano98n] Anonymous. Departments — portfolio — unlocking the subconscious. Computer Graphics World, 21(10):58–??, October 1998. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1998:DSN

[Ano98o] Anonymous. Departments — spotlight — news and products. Computer Graphics World, 21 (10):8–??, October 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:DTW

[Ano98p] Anonymous. Departments — tech watch — dissection on the Web. an interface for large screens. Computer Graphics World, 21(10):16– ??, October 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:DUF

[Ano98q] Anonymous. Departments — user focus — adding harmony to 3D. creating prehistoric sharks. CAD toolmaker turns user. how drafting works. Computer Graphics World, 21(10):60–??, October 1998. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1998:DVB

[Ano98r] Anonymous. Departments — viewpoint — benchmark myths. Computer Graphics World, 21 (10):21–??, October 1998. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1998:DSG

[Ano98s] Anonymous. Digital studio — 3D goes to the opera. cool looks for Barbie. Computer Graphics World, 21(6):15–??, June 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:DSD

[Ano98t] Anonymous. Digital studio — developing Kingsley. Computer Graphics World, 21(2):15– ??, February 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:DST

[Ano98u] Anonymous. Digital studio — top secret ride film. the riddler’s revenge. Computer Graphics World, 21(4):15–??, April 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:ERa

[Ano98v] Anonymous. Events and resources. Computer Graphics World, 21(7):132–??, July 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:ERb

[Ano98w] Anonymous. Events and resources. Computer Graphics World, 21(8):80–??, August 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:ERc

[Ano98x] Anonymous. Events and resources. Computer Graphics World, 21(9):76–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:GSW

[Ano98y] Anonymous. Godzilla: See what it took to bring this giant lizard and his babies to life. Computer Graphics World, 21(7):49–??, July 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:I

[Ano98z] Anonymous. Insight. Computer Graphics World, 21(7):160– ??, July 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:IMA

[Ano98-27] Anonymous. Insight — magnetic attraction. Computer Graphics World, 21(8):92–??, August 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:IMO

[Ano98-28] Anonymous. Insight — mixing oil and water. Computer Graphics World, 21(9):84–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:IAD

[Ano98-29] Anonymous. International airport design. Computer Graphics World, 21(9):38–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:MO

[Ano98-30] Anonymous. Market outlook. Computer Graphics World, 21(7): 17–??, July 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:MOM

[Ano98-31] Anonymous. Market outlook — moving to MCAD’s midrange. Computer Graphics World, 21(9): 17–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:MOP

[Ano98-32] Anonymous. Market outlook — PC graphics grows up. Computer Graphics World, 21(8):17–??, August 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:MSA

[Ano98-33] Anonymous. Model shop — animating a scorpion’s sting. Computer Graphics World, 21(2):71– ??, February 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:MSC

[Ano98-34] Anonymous. Model shop — creating organic models. Computer Graphics World, 21(1):67–??, January 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:P

[Ano98-35] Anonymous. Portfolio. Computer Graphics World, 21(7):112– ??, July 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:PLT

[Ano98-36] Anonymous. Portfolio — A light touch. Computer Graphics World, 21(6):70–??, June 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:PAL

[Ano98-37] Anonymous. Portfolio — art lessons. Computer Graphics World, 21(9):60–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:PAE

[Ano98-38] Anonymous. Portfolio — artist ex machina. Computer Graphics World, 21(1):60–??, January 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:PDS

[Ano98-39] Anonymous. Portfolio — digital set design. Computer Graphics World, 21(3):86–??, March 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:PFE

[Ano98-40] Anonymous. Portfolio — Flubber’s emotional range. Computer Graphics World, 21(2):62– ??, February 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:PMM

[Ano98-41] Anonymous. Portfolio — making movies, pronto. Computer Graphics World, 21(4):78–??, April 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:PNB

[Ano98-42] Anonymous. Portfolio — natural beauty. Computer Graphics World, 21(5):68–??, May 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:PPA

[Ano98-43] Anonymous. Portfolio — Prix Ars Electronica winners. Computer Graphics World, 21(8):64–??, August 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:RE

[Ano98-44] Anonymous. Resources and events. Computer Graphics World, 21(5):86–??, May 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:RSM

[Ano98-45] Anonymous. Reviews — 3D Studio Max 2.0. Alias/Wavefront Composer 4.5. Computer Graphics World, 21(1):73–??, January 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:RAD

[Ano98-46] Anonymous. Reviews — ArchiDesign. DataCAD 8.0. Computer Graphics World, 21(6):83–??, June 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:RBS

[Ano98-47] Anonymous. Reviews — Bryce 3D. SolidWorks 98. Computer Graphics World, 21(4):85–??, April 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:RCV

[Ano98-48] Anonymous. Reviews — Catalyst V 1.1. IntelliCAD98. Computer Graphics World, 21(3):97– ??, March 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:RLM

[Ano98-49] Anonymous. Reviews — Lightscape 3.1, Michanicol Desktop 2.0. Computer Graphics World, 21(2):77– ??, February 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:RPD

[Ano98-50] Anonymous. Reviews — Piranesi. Design Wave 2.0. Computer Graphics World, 21(5):77– ??, May 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:SSC

[Ano98-51] Anonymous. Small soldiers: The characters are short, but animating them was a tall order. Computer Graphics World, 21(7):38– ??, July 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:S

[Ano98-52] Anonymous. Spotlight. Computer Graphics World, 21(7):10–??, July 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:SNPa

[Ano98-53] Anonymous. Spotlight — news and products. Computer Graphics World, 21(8):10–??, August 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:SNPb

[Ano98-54] Anonymous. Spotlight — news and products. Computer Graphics World, 21(9):10–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:SRD

[Ano98-55] Anonymous. A STEP in the right direction. Computer Graphics World, 21(9):48–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:TCO

[Ano98-56] Anonymous. Taking collaboration into orbit. Computer Graphics World, 21(9):34–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:TW

[Ano98-57] Anonymous. Tech watch. Computer Graphics World, 21(7):22– ??, July 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:TWMb

[Ano98-58] Anonymous. Tech watch — A movie with heart. the NICE way to learn. Computer Graphics World, 21(5):13–??, May 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:TWN

[Ano98-59] Anonymous. Tech watch — A new way to view volumes. realistic facial animation. Computer Graphics World, 21(9):18–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:TWMa

[Ano98-60] Anonymous. Tech watch — modeling life’s complexities. VR in medicine 3D camera technology. blue lasers for optical storage. Computer Graphics World, 21(1): 17–??, January 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:TWMc

[Ano98-61] Anonymous. Tech watch — music and graphics in harmony. averting aviation disasters. visualization drives Mars exploration. Computer Graphics World, 21(8): 20–??, August 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:TBC

[Ano98-62] Anonymous. TechWatch — from bones to clones. Computer Graphics World, 21(3):17–??, March 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:TWT

[Ano98-63] Anonymous. TechWatch — walking through VR. Computer Graphics World, 21(3):17–??, March 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:UF

[Ano98-64] Anonymous. User focus. Computer Graphics World, 21(7):117– ??, July 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:UFG

[Ano98-65] Anonymous. User focus — A geological adventure. Computer Graphics World, 21(1):63–??, January 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:UFN

[Ano98-66] Anonymous. User focus — A new bike design. Computer Graphics World, 21(2):67–??, February 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:UFA

[Ano98-67] Anonymous. User focus — analyzing jet fighters. virtual world cup coverage. 3D printers in toyland. cyber athlete meets the press. Computer Graphics World, 21(9): 64–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:UFSa

[Ano98-68] Anonymous. User focus — simulating a tragic crash. A spokescharacter with attitude. Computer Graphics World, 21(6): 75–??, June 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:UFSb

[Ano98-69] Anonymous. User focus — simulating eye surgery. Deep Impact’s tidal wave building. Smithsonian exhibits with CAD. Computer Graphics World, 21(8):69–??, August 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:UFTa

[Ano98-70] Anonymous. User focus — the thrill of the chase. spinal map. Computer Graphics World, 21(3): 89–??, March 1998. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1998:UFTb

[Ano98-71] Anonymous. User focus — time travels. Computer Graphics World, 21(4):81–??, April 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:UFV

[Ano98-72] Anonymous. User focus — virtual dancers get moving. Computer Graphics World, 21(5):73–??, May 1998. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1998:VMC

[Ano98-73] Anonymous. Viewpoint — making CAD pay. Computer Graphics World, 21(8):27–??, August 1998. CODEN CGWODH. ISSN 0271-4159.

Applications:1993:IET

[App93] Applications. Interactive exhibit teaches scientific concepts; Designing and modeling the car of the future. Computer Graphics World, 16(5):49–??, May 1993. CODEN CGWODH. ISSN 0271-4159.

Ashdown:1996:LAH

[Ash96] Ian Ashdown. Lighting for architects — how do you model light within a computer graphics program? the answer encompasses two basic approaches: ray tracing and radiosity. Computer Graphics World, 19(8):38–??, August 1996. CODEN CGWODH. ISSN 0271-4159.

Beal:1991:SD

[Bea91] Ellen Beal. Smart Documents.

Computer Graphics World, 14(5): 53–??, May 1991. CODEN CG-WODH. ISSN 0271-4159.

Belleville:1993:GPP

[Bel93] Laureen Belleville. The Games People Play. Computer Graphics World, 16(12):22–??, December 1993. CODEN CGWODH. ISSN 0271-4159.

Blackburn:1998:M

[BM98] Dave Blackburn and Karen Moltenbrey. Moonstruck. Computer Graphics World, 21(9):50– ??, September 1998. CODEN CG-WODH. ISSN 0271-4159.

Boyd:1998:CE

[Boy98a] Lane Boyd. Collaborative efforts. Computer Graphics World, 21(9): 33–??, September 1998. CODEN CGWODH. ISSN 0271-4159.

Boyd:1998:DFS

[Boy98b] Lane Boyd. Digital factories. Computer Graphics World, 21(5): 45–??, May 1998. CODEN CG-WODH. ISSN 0271-4159.

Brill:1992:FII

[Bri92] Louis M. Brill. Facing Interface Issues. Computer Graphics World, 15(4):48–??, April 1992. CODEN CGWODH. ISSN 0271-4159.

Brill:1993:KTV

[Bri93] Louis M. Brill. Kicking the Tires of VR Software. Computer Graphics World, 16(6):40–??, June 1993. CODEN CGWODH. ISSN 0271-4159.

Bruvel:1996:MF

[Bru96] Gil Bruvel. Making faces. Computer Graphics World, 19(1):66– 69, January 1996. CODEN CG-WODH. ISSN 0271-4159.

Barnes-Svarney:1992:GG

[BS92] Patricia Barnes-Svarney. Galactic Graphics. Computer Graphics World, 15(12):36–??, December 1992. CODEN CGWODH. ISSN 0271-4159.

Bar-Zeev:1994:LBV

[BZJ94] Avi Bar-Zeev and Robert Jacobson. Looking Back From the Virtual Future. Computer Graphics World, 17(5):30–??, May 1994. CODEN CGWODH. ISSN 0271-4159.

CadCam:1993:PIW

[Cad93] Cad/Cam. Predicting the impact of Windows NT on Unix. Computer Graphics World, 16(4):67– ??, April 1993. CODEN CG-WODH. ISSN 0271-4159.

Cavuoto:1992:PP

[Cav92] Jim Cavuoto. Picture-Perfect Printing. Computer Graphics World, 15(1):44–??, January 1992. CODEN CGWODH. ISSN 0271-4159.

Copeland:1998:LP

[CC98] Tom Copeland and Donna Coco.

Leaders of the pack. Computer Graphics World, 21(9):22– ??, September 1998. CODEN CG-WODH. ISSN 0271-4159.

CGWStaff:1992:PF

[CGW92a] CGW Staff. A Peek Into the Future. Computer Graphics World, 15(12):42–??, December 1992. CODEN CGWODH. ISSN 0271-

CGWStaff:1992:IGP

[CGW92b] CGW Staff. Images of Graphics Past. Computer Graphics World, 15(12):52–??, December 1992. CODEN CGWODH. ISSN 0271-4159.

Clarkson:1996:ART

[Cla96] Mark Clarkson. 3D accelerators, round two — A new generation of 3D graphics accelerator boards for Windows NT are coming to market just a little over a year after the first generation hit the streets. and for just about the same price as last year’s models, you can get a lot more functionality. Computer Graphics World, 19(11):67– ??, November 1996. CODEN CG-WODH. ISSN 0271-4159.

Coco:1994:IGD

[Coc94a] D. Coco. Inside the Game Developer’s Toolbox. Computer Graphics World, 17(10):18–??, October 1994. CODEN CGWODH. ISSN 0271-4159.

Coco:1994:DPI

[Coc94b] Donna Coco. Developing Programs for Interactive TV. Computer Graphics World, 17(6):33– ??, June 1994. CODEN CG-WODH. ISSN 0271-4159.

Coco:1994:SVD

[Coc94c] Donna Coco. Squeezing Video Onto the Desktop. Computer Graphics World, 17(3):45–??, March 1994. CODEN CGWODH. ISSN 0271-4159.

Coco:1995:PRT

[Coc95a] D. Coco. 3D on the PC: Recent technological advances in both hardware and software are making cost-effective 3D accelerators for the PC a reality. Computer Graphics World, 18(2):22–??, February 1995. CODEN CGWODH. ISSN 0271-4159.

Coco:1995:CP

[Coc95b] Donna Coco. 3D conquers the PC.

Computer Graphics World, 18(2): ??, February 1995. CODEN CG-WODH. ISSN 0271-4159.

Coco:1995:BLH

[Coc95c] Donna Coco. Breathing life Into 3D Humans. Computer Graphics World, 18(8):40–??, August 1995. CODEN CGWODH. ISSN 0271-4159.

Coco:1995:CSE

[Coc95d] Donna Coco. Compositing and Special Effects. Computer Graphics World, 18(5):46–??, May 1995. CODEN CGWODH. ISSN 0271-4159.

Coco:1995:IGB

[Coc95e] Donna Coco. Interactivity Goes Big Screen: NYC-based Interfilm brings new meaning to the term interactive as its all-digital, big-screen films come to theaters nationwide. Computer Graphics World, 18(3):46–??, March 1995. CODEN CGWODH. ISSN 0271-4159.

Coco:1995:RGT

[Coc95f] Donna Coco. Real-Time 3D Games Take Off — The skills needed to create a real-time 3D game are a bit different from those best-suited to 2D or to prerendered 3D development. Computer Graphics World, 18(12):22– ??, December 1995. CODEN CG-WODH. ISSN 0271-4159.

Coco:1995:WWE

[Coc95g] Donna Coco. Water, Water Everywhere. Computer Graphics World, 18(9):38–??, September 1995. CO-DEN CGWODH. ISSN 0271-4159.

Coco:1996:DSE

[Coc96a] Donna Coco. Desktop special effects — today’s desktop specialeffects market is quite diverse, boasting a range of packages offering not only enhanced morphing functionality but also the ability to warp and distort images and video as well as create digitalvideo effects and transitions. Computer Graphics World, 19(7):44– ??, July 1996. CODEN CG-WODH. ISSN 0271-4159.

Coco:1996:NTG

[Coc96b] Donna Coco. New tools for 3D gamers — in the last year, vendors of 3D software took a particular interest in game developers, bringing to market a wealth of new products with modeling and animation features geared specifically at their needs. Computer Graphics World, 19(12):29–??, December 1996. CODEN CGWODH. ISSN 0271-4159.

Coco:1997:ASR

[Coc97a] Donna Coco. Animation special report: Choosing the “right” animation system. Computer Graphics World, 20(8):62–??, August 1997. CODEN CGWODH. ISSN 0271-4159.

Coco:1997:CHG

[Coc97b] Donna Coco. Creating humans for games. Computer Graphics World, 20(10):26–??, October 1997. CODEN CGWODH. ISSN 0271-4159.

Coco:1997:CIC

[Coc97c] Donna Coco. Creating intelligent creatures. Computer Graphics World, 20(7):22–??, July 1997. CODEN CGWODH. ISSN 0271-4159.

Coco:1997:DEV

[Coc97d] Donna Coco. Developers embrace VRML 2.0 — with the arrival of VRML 2.0, Web developers are creating 3D environments that are more exciting and robust than ever. but that doesn’t mean there aren’t problems. the technology’s still new, and working with it requires patience, understanding, and a healthy degree of creativity to navigate around obstacles. Computer Graphics World, 20(4): 32–??, April 1997. CODEN CG-WODH. ISSN 0271-4159.

Coco:1997:IDY

[Coc97e] Donna Coco. Interacting with dinosaurs. Computer Graphics World, 20(6):42–??, June 1997. CODEN CGWODH. ISSN 0271-4159.

Coco:1997:MCA

[Coc97f] Donna Coco. Motion capture advances. Computer Graphics World, 20(11):37–??, November 1997. CODEN CGWODH. ISSN 0271-4159.

Coco:1997:TVD

[Coc97g] Donna Coco. Tools for VRML 2.0 development. Computer Graphics World, 20(12):37–??, December 1997. CODEN CGWODH. ISSN 0271-4159.

Coco:1998:STB

[Coc98a] Donna Coco. 3D stars on TV. Computer Graphics World, 21(4): 40–??, April 1998. CODEN CG-WODH. ISSN 0271-4159.

Coco:1998:CGY

[Coc98b] Donna Coco. CG that gives you the creeps. Computer Graphics World, 21(3):52–??, March 1998. CODEN CGWODH. ISSN 0271-4159.

Coco:1998:SAE

[Coc98c] Donna Coco. Space: An artistic exploration of the final frontier. Computer Graphics World, 21(8): 33–??, August 1998. CODEN CG-WODH. ISSN 0271-4159.

Coco:1998:SS

[Coc98d] Donna Coco. Stimulating simulations. Computer Graphics World, 21(7):92–??, July 1998. CODEN CGWODH. ISSN 0271-4159.

Coco:1998:WDD

[Coc98e] Donna Coco. A Web developer’s dream. Computer Graphics World, 21(8):54–??, August 1998. CODEN CGWODH. ISSN 0271-4159.

Cohn:1996:RCB

[Coh96] Davois S. Cohn. Reading computer benchmarks — although flawed, computer benchmarks provide one of the few ways of quantifying system performance. but you’ve got to be able to analyze the test results correctly, or you’ll be fooled into making assumptions about performance that will be entirely misleading. Computer Graphics World, 19(12):49– ??, December 1996. CODEN CG-WODH. ISSN 0271-4159.

Cohn:1998:DP

[Coh98a] David Cohn. Digital parts. Computer Graphics World, 21(7):103– ??, July 1998. CODEN CG-WODH. ISSN 0271-4159.

Cohn:1998:RPA

[Coh98b] David Cohn. Reversing the process. Computer Graphics World, 21(10):42–??, October 1998. CO-DEN CGWODH. ISSN 0271-4159.

Cook:1990:DA

[Coo90a] Rick Cook. Designing Assemblies.

Computer Graphics World, 13 (11):42–??, November 1990. CO-DEN CGWODH. ISSN 0271-4159.

Cook:1990:ICS

[Coo90b] Rick Cook. Image Compression Squeezes to the Fore. Computer Graphics World, 13(9):70– ??, September 1990. CODEN CG-WODH. ISSN 0271-4159.

Cook:1992:SE

[Coo92] Rick Cook. Serious Entertainment. Computer Graphics World, 15(5):40–??, May 1992. CODEN CGWODH. ISSN 0271-4159.

Cook:1993:AII

[Coo93] Rick Cook. The Amiga: Is it Time You Took a Second Look? Computer Graphics World, 16(8):30– ??, August 1993. CODEN CG-WODH. ISSN 0271-4159.

Cavuoto:1992:DPD

[CP92] James Cavuoto and David Pope. Digital Proofs on Demand. Computer Graphics World, 15(6):59– ??, June 1992. CODEN CG-WODH. ISSN 0271-4159.

Cramblitt:1991:LHR

[Cra91] Bob Cramblitt. A Love/Hate Relationship. Computer Graphics World, 14(1):55–??, January 1991. CODEN CGWODH. ISSN 0271-4159.

Cramblitt:1994:RNE

[Cra94] B. Cramblitt. Rendering on a Network: With the exception of animators who are taking advantage of the multiplatform capabilities of Pixar’s NetRenderMan, users of commercial animation packages have been pretty much excluded from the network-rendering revolution. Computer Graphics World, 17(8):28–??, August 1994. CO-DEN CGWODH. ISSN 0271-4159.

Davis:1991:RC

[Dav91] Dwight B. Davis. Reality Check. Computer Graphics World, 14(6): 49–??, June 1991. CODEN CG-WODH. ISSN 0271-4159.

Davis:1992:BC

[Dav92] Dwight Davis. Big-Time Changes.

Computer Graphics World, 15(7): 42–??, July 1992. CODEN CG-WODH. ISSN 0271-4159.

DeYoung:1995:PF

[DeY95] H. Garrett DeYoung. Photo finishes. Computer Graphics World, 18(1):??, January 1995. CODEN CGWODH. ISSN 0271-4159.

Doyle:1991:CC

[Doy91a] Audrey V. Doyle. Cruise Control. Computer Graphics World, 14(10):29–??, October 1991. CO-DEN CGWODH. ISSN 0271-4159.

Doyle:1991:EP

[Doy91b] Audrey V. Doyle. Environmental Protection. Computer Graphics World, 14(7):78–??, July 1991. CODEN CGWODH. ISSN 0271-4159.

Doyle:1994:PMT

[Doy94] A. V. Doyle. Playing with the Music: With today’s interactive music CD-ROMs, which combine sophisticated computer graphics and multimedia technologies, people can not only play the music created by some of their favorite recording artists, they can play with the music in ways never before available. Computer Graphics World, 17(8):38–??, August 1994. CODEN CGWODH. ISSN 0271-4159.

Doyle:1995:SPY

[Doy95] Audrey Doyle. SIGGRAPH ’95 Preview — This year’s conference and exhibition, to be held next month in Los Angeles. promises to challenge, fascinate, and enlighten. Computer Graphics World, 18(7):65–??, July 1995. CODEN CGWODH. ISSN 0271-4159.

Dunniway:1997:PBS

[Dun97] Troy Dunniway. 3D paints a broader stroke. Computer Graphics World, 20(1):18–??, January 1997. CODEN CGWODH. ISSN 0271-4159.

Emigh:1991:MED

[Emi91] Jacqueline Emigh. Managing Engineering Data. Computer Graphics World, 14(1):67–??, January 1991. CODEN CGWODH. ISSN 0271-4159.

Emmett:1986:DMB

[Emm86] Arielle Emmett. Do managers believe what they see? Computer Graphics World, 9(9):50–52, 54, 57, September 1986. CODEN CG-WODH. ISSN 0271-4159.

Emmett:1991:GS

[Emm91a] Arielle Emmett. Guardians of Style. Computer Graphics World, 14(6):30–??, June 1991. CODEN CGWODH. ISSN 0271-4159.

Emmett:1991:SMH

[Emm91b] Arielle Emmett. In Search of the Miracle Hologram. Computer Graphics World, 14(2):44– ??, February 1991. CODEN CG-WODH. ISSN 0271-4159.

Emmett:1991:AL

[Emm91c] Arielle Emmett. The ABCs of LCDs. Computer Graphics World, 14(9):52–??, September 1991. CO-DEN CGWODH. ISSN 0271-4159.

Emmett:1992:ND

[Emm92a] Arielle Emmett. A New Dynamism. Computer Graphics World, 15(6):30–??, June 1992. CODEN CGWODH. ISSN 0271-4159.

Emmett:1992:OD

[Emm92b] Arielle Emmett. Optimized Designs. Computer Graphics World, 15(2):30–??, February 1992. CO-DEN CGWODH. ISSN 0271-4159.

Emmett:1992:SE

[Emm92c] Arielle Emmett. Something for Everybody. Computer Graphics World, 15(7):29–??, July 1992. CODEN CGWODH. ISSN 0271-4159.

Emmett:1994:TPC

[Emm94] A. Emmett. Transcending Photorealism: Computer artists are expressing new optimism for the future, as they take a variety of approaches in an effort to achieve a more flowing, expressive style in their work. Computer Graphics World, 17(9):30–??, September 1994. CODEN CGWODH. ISSN 0271-4159.

Emmett:1995:TTA

[Emm95a] Arielle Emmett. 3D Type — With today’s animated 3D titling packages, which offer an array of typography, compositing, and post-production features, conjuring up something more than just the jaded flying font is fairly easy. Computer Graphics World, 18 (11):35–??, November 1995. CO-DEN CGWODH. ISSN 0271-4159.

Emmett:1995:CLP

[Emm95b] Arielle Emmett. Color Laser Printing Technology. Computer Graphics World, 18(6):46–50, June 1995. CODEN CGWODH. ISSN 0271-4159.

Emmett:1995:PCA

[Emm95c] Arielle Emmett. Preparing for a Career in Animation. Computer Graphics World, 18(8):55–??, August 1995. CODEN CGWODH. ISSN 0271-4159.

Emmett:1996:DPD

[Emm96a] Arielle Emmett. Dye-sublimation printers — dye-sub devices stand alone in terms of their ability to deliver true photo-quality output. Computer Graphics World, 19(10):47–??, October 1996. CO-DEN CGWODH. ISSN 0271-4159.

Emmett:1996:SFR

[Emm96b] Arielle Emmett. Scanners and film recorders at the high end — thanks to the explosion in the use of CG technology in films, high-end film scanners and film recorders. Computer Graphics World, 19(9):53–??, September 1996. CODEN CGWODH. ISSN 0271-4159.

Emmett:1996:TTC

[Emm96c] Arielle Emmett. Tuning in to tablets — choosing which 2D digitizer is right for your application requires careful analysis of such subtleties as ergonomics, weight, and size. Computer Graphics World, 19(2):44–??, February 1996. CODEN CGWODH. ISSN 0271-4159.

Emmett:1996:VSL

[Emm96d] Arielle Emmett. Vendors streamline large-format scanners — the increasing use of large-format scanners is driving improvements in technology as well as software. Computer Graphics World, 19(1): 46–??, January 1996. CODEN CGWODH. ISSN 0271-4159.

Emmett:1998:VJV

[Emm98] Arielle Emmett. Virtual journeys.

Computer Graphics World, 21(5): 34–??, May 1998. CODEN CG-WODH. ISSN 0271-4159.

Forcade:1995:FFE

[FA95] Tim Forcade and Mark Anderson.

Feeling frustrated? explode something. Computer Graphics World, 18(2):??, February 1995. CODEN CGWODH. ISSN 0271-4159.

Fangmeier:1996:DDT

[Fan96] Stefen Fangmeier. Designing digital tornados — how the artists and designers at Industrial Light and Magic overcame the many challenges they encountered when creating Twister’s tornados. Computer Graphics World, 19(8):58– ??, August 1996. CODEN CG-WODH. ISSN 0271-4159.

Fichera:1986:LWF

[Fic86] Richard Fichera. Low-end workstations force market barriers down. Computer Graphics World, 9(9):34–35, 37, 42, September 1986. CODEN CGWODH. ISSN 0271-4159.

Friedhoff:1990:EB

[FK90] Richard Mark Friedhoff and Thomas Kiely. The Eye of the Beholder. Computer Graphics World, 13(8):46–??, August 1990. CODEN CGWODH. ISSN 0271-4159.

Flynn:1992:NAC

[Fly92] Bill Flynn. A New Age in Color Hard-Copy. Computer Graphics World, 15(8):45–??, August 1992. CODEN CGWODH. ISSN 0271-4159.

Forcade:1992:EV

[For92] Tim Forcade. Engineering Visualization. Computer Graphics World, 15(11):37–??, November 1992. CODEN CGWODH. ISSN 0271-4159.

Forcade:1993:EHEa

[For93a] Tim Forcade. Evaluating 3D on the High End. Computer Graphics World, 16(10):44–??, October 1993. CODEN CGWODH. ISSN 0271-4159.

Forcade:1993:EHEb

[For93b] Tim Forcade. Evaluating 3D on the High End. Computer Graphics World, 16(11):57–??, November 1993. CODEN CGWODH. ISSN 0271-4159.

Forcade:1993:SRR

[For93c] Tim Forcade. Striving for Realism with RenderMan. Computer Graphics World, 16(4):24– ??, April 1993. CODEN CG-WODH. ISSN 0271-4159.

Forcade:1997:PPO

[For97] Tim Forcade. 3D plug-ins proliferate — once the province of 3D Studio, plug-in fever has spread to numerous other 3D modeling and animation packages. in fact, the frantic pace at which 3D plug-ins are appearing is outstripping and driving the previously centralized 3D software development process. Computer Graphics World, 20(2): 32–??, February 1997. CODEN CGWODH. ISSN 0271-4159.

Forcade:1998:URC

[For98] Tim Forcade. Unraveling Riven.

Computer Graphics World, 21(2): 49–??, February 1998. CODEN CGWODH. ISSN 0271-4159.

Freiherr:1991:VNB

[FP91] Greg Freiherr and Gary Pfitzer.

Visualization’s New Breed. Computer Graphics World, 14(1):44– ??, January 1991. CODEN CG-WODH. ISSN 0271-4159.

Freiherr:1990:CHF

[Fre90a] Greg Freiherr. CAD on the High Frontier. Computer Graphics World, 13(11):54–??, November 1990. CODEN CGWODH. ISSN 0271-4159.

Freiherr:1990:GSI

[Fre90b] Greg Freiherr. Getting Small with Image Processing. Computer Graphics World, 13(9):44– ??, September 1990. CODEN CG-WODH. ISSN 0271-4159.

Freiherr:1990:PPP

[Fre90c] Greg Freiherr. PACS: A Premature Paradise. Computer Graphics World, 13(4):52–??, April 1990. CODEN CGWODH. ISSN 0271-4159.

Freiherr:1991:ROP

[Fre91] Greg Freiherr. Rediscovering the Outer Planets. Computer Graphics World, 14(7):36–??, July 1991. CODEN CGWODH. ISSN 0271-4159.

Friedhoff:1991:M

[Fri91] Richard Mark Friedhoff. Microvisualization. Computer Graphics World, 14(5):38–??, May 1991. CODEN CGWODH. ISSN 0271-4159.

Fritz:1992:RP

[Fri92] Mark Fritz. Rousing Presentations. Computer Graphics World, 15(8):31–??, August 1992. CO-DEN CGWODH. ISSN 0271-4159.

GarrettDeYoung:1995:PFW

[Gar95] H. Garrett DeYoung. Photo Finishes: Whether they utilize dyesublimation, inkjet, or thermal wax-transfer technology, today’s low-cost photorealistic printers offer numerous design and marketing advantages to hundreds of companies. Computer Graphics World, 18(1):46–??, January 1995. CODEN CGWODH. ISSN 0271-4159.

Glatzer:1998:CDC

[Gla98] Hal Glatzer. CD or DVD? Computer Graphics World, 21(2):57– ??, February 1998. CODEN CG-WODH. ISSN 0271-4159.

Graves:1993:HA

[Gra93a] Gaye L. Graves. Hell-Bent on Adventure. Computer Graphics World, 16(7):71–??, July 1993. CODEN CGWODH. ISSN 0271-4159.

Graves:1993:MM

[Gra93b] Gaye L. Graves. The Magic of Metaballs. Computer Graphics World, 16(5):26–??, May 1993. CODEN CGWODH. ISSN 0271-4159.

Graves:1994:DTD

[Gra94] Gaye L. Graves. Don’t Touch That Dial! Computer Graphics World, 17(3):17–??, March 1994. CODEN CGWODH. ISSN 0271-4159.

Grabowski:1996:WVL

[Gra96] Ralph Grabowski. Will VRML 2.0 let me design a bolt faster? Computer Graphics World, 19(12):22, 24, December 1996.

Gross:1991:MMM

[Gro91] Daniel Gross. Merging Man and Machine. Computer Graphics World, 14(5):47–??, May 1991. CODEN CGWODH. ISSN 0271-4159.

Haase:1994:NSE

[Haa94] Bruce Haase. The Neon: A Shining Example of Design Efficiency. Computer Graphics World, 17(2): 41–??, February 1994. CODEN CGWODH. ISSN 0271-4159.

Hansen:1990:BCB

[Han90] Rebecca Hansen. Breaking the Color Barrier. Computer Graphics World, 13(7):38–??, July 1990.

CODEN CGWODH. ISSN 0271-4159.

Harbert:1996:AVD

[Har96a] Tam Harbert. A/V drive vendors expand their offerings — today’s product offerings provide video and animation professionals with such benefits as increased storage space and uninterrupted data flow. Computer Graphics World, 19(4): 43–??, April 1996. CODEN CG-WODH. ISSN 0271-4159.

Harbert:1996:SRS

[Har96b] Tam Harbert. Sorting out RAID storage — the need for faster data transfer and more storage space is driving many graphics professionals to use RAID systems. Computer Graphics World, 19(9):61– ??, September 1996. CODEN CG-WODH. ISSN 0271-4159.

Hars:1996:CEM

[Har96c] Adele Hars. CAD in Europe — many Americans feel Europeans are 10 years behind in their use of CAD technology, while numerous Europeans believe the Americans will . . .. Computer Graphics World, 19(5):21–??, May 1996. CODEN CGWODH. ISSN 0271-4159.

Hars:1996:MMC

[Har96d] Adele Hars. Masters of motion capture — one of the leaders in performance animation technology is Medialab, an animation production house based in Paris. its real time characters perform on television and serve as corporate mascots. Computer Graphics World, 19(10):26–??, October 1996. CODEN CGWODH. ISSN 0271-4159.

Hastings:1992:TSR

[Has92] Bryan Hastings. Taking the Show on the Road. Computer Graphics World, 15(8):38–??, August 1992. CODEN CGWODH. ISSN 0271-4159.

Helsel:1997:HVD

[Hel97] Sandra Helsel. High-res VR displays — today, the range of displays for virtual reality runs the gamut from head-mounted displays to virtual-model displays. each of these different types of displays has their own unique advantages, and for those who want to understand the potential of VR, a good place to begin is to understand the various VR display options. Computer Graphics World, 20(5):49–??, May 1997. CODEN CGWODH. ISSN 0271-4159.

Hirsch:1994:CD

[Hir94] Evan Marc Hirsch. 3D CAD on the desktop. Computer Graphics World, 17(2):22–??, February 1994. CODEN CGWODH. ISSN 0271-4159.

Hirsch:1996:CLA

[Hir96] Evan Marc Hirsch. A changing landscape — the animation marketplace has undergone some dramatic changes over the past few years. here, we try to make sense of these changes and help determine the impact they will have on the animation industry. Computer Graphics World, 19(4):21– ??, April 1996. CODEN CG-WODH. ISSN 0271-4159.

Hirsch:1998:NTO

[Hir98] Evan Hirsch. New technology in the old country. Computer Graphics World, 21(3):70–??, March 1998. CODEN CGWODH. ISSN 0271-4159.

Hodges:1998:IJF

[Hod98a] Mark Hodges. It just feels right. Computer Graphics World, 21 (10):48–??, October 1998. CO-DEN CGWODH. ISSN 0271-4159.

Hodges:1998:VRT

[Hod98b] Mark Hodges. Virtual reality in training. Computer Graphics World, 21(8):45–??, August 1998. CODEN CGWODH. ISSN 0271-4159.

Hodges:1998:VSH

[Hod98c] Mark Hodges. Visualization spoken here. Computer Graphics World, 21(4):55–??, April 1998. CODEN CGWODH. ISSN 0271-4159.

Holch:1993:SCW

[Hol93] Kristy Holch. Scanning in a Colorful World. Computer Graphics World, 16(10):59–??, October 1993. CODEN CGWODH. ISSN 0271-4159.

Holsinger:1994:FSD

[Hol94] Erik Holsinger. Finding success with desktop video effects. Computer Graphics World, 17(1):22– ??, January 1994. CODEN CG-WODH. ISSN 0271-4159.

Hollander:1998:TUV

[Hol98] Richard Hollander. 3D tracking:

A user’s view. Computer Graphics World, 21(5):60–??, May 1998. CODEN CGWODH. ISSN 0271-4159.

Iversen:1991:ATS

[Ive91] Wesley Iversen. Animation Takes the Stand. Computer Graphics World, 14(11):48–??, November 1991. CODEN CGWODH. ISSN 0271-4159.

Iversen:1992:SS

[Ive92a] Wes Iversen. The Sound of Science. Computer Graphics World, 15(1):54–??, January 1992. CO-DEN CGWODH. ISSN 0271-4159.

Iversen:1992:DP

[Ive92b] Wesley R. Iversen. Digital Photography. Computer Graphics World, 15(2):67–??, February 1992. CO-DEN CGWODH. ISSN 0271-4159.

Iversen:1992:PF

[Ive92c] Wesley R. Iversen. Photo Finish.

Computer Graphics World, 15(3): 75–??, March 1992. CODEN CG-WODH. ISSN 0271-4159.

Johnston:1990:CP

[Joh90] Peter Johnston. From Computer to Page. Computer Graphics World, 13(1):28–??, January 1990. CODEN CGWODH. ISSN 0271-4159.

Keep:1986:SNA

[Kee86] Peter Keep. Segments: a natural approach to graphics. Computer Graphics World, 9(9):60–62, 64, September 1986. CODEN CG-WODH. ISSN 0271-4159.

King:1993:H

[Kin93] Douglas King. Heads Up. Computer Graphics World, 16(11):41– ??, November 1993. CODEN CG-WODH. ISSN 0271-4159.

King:1995:MVW

[Kin95] Douglas King. Modeling virtual worlds. Computer Graphics World, 18(2):67–68, February 1995. CODEN CGWODH. ISSN 0271-4159.

King:1997:PMY

[Kin97] Doug King. Picture-perfect models — are you still creating 3D models the old-fashioned way? now there are a handful of commercial software packages that employ photogrammetric algorithms that enable you to create 3D models from photographs. true, they’re not the perfect solution for every application, but for. . .. Computer Graphics World, 20(1):39–??, January 1997. CO-DEN CGWODH. ISSN 0271-4159.

Keery:1991:BBG

[KPV91] Nina Keery, Gary Pfitzer, and Audrey Vasilopoulos. Broadening the Boundaries of Graphic Art and Illustration. Computer Graphics World, 14(2):65–??, February 1991. CODEN CGWODH. ISSN 0271-4159.

Krum:1993:MPP

[Kru93] Ted Krum. More Power on the PC. Computer Graphics World, 16(9):54–??, September 1993. CO-DEN CGWODH. ISSN 0271-4159.

Kurland:1997:MMC

[Kur97] Raymond Kurland. Mid-range mechanical CAD: A comparison — high-end CAD/CAM programs are starting to feel the heat from low-cost, production-worthy, midrange CAD packages. but just how sophisticated are those programs, and how do they compare in their modeling capabilities? this benchmark comparison provides some clues. Computer Graphics World, 20(1):29–??, January 1997. CO-DEN CGWODH. ISSN 0271-4159.

Lang:1990:DI

[Lan90] Laura Lang. Designer Interiors. Computer Graphics World, 13 (12):34–??, December 1990. CO-DEN CGWODH. ISSN 0271-4159.

Lang:1992:GCL

[Lan92a] Laura Lang. GIS Comes to Life. Computer Graphics World, 15(10):27–??, October 1992. CO-DEN CGWODH. ISSN 0271-4159.

Lang:1992:TF

[Lan92b] Laura Lang. Technology Fusion.

Computer Graphics World, 15(3): 63–??, March 1992. CODEN CG-WODH. ISSN 0271-4159.

Lang:1993:TM

[Lan93] Laura Lang. Terrain Modeling.

Computer Graphics World, 16(9): 22–??, September 1993. CODEN CGWODH. ISSN 0271-4159.

LeWinter:1990:AC

[LB90a] Renee LeWinter and Cynthia Baron. Artistic Challenge. Computer Graphics World, 13(2):48– ??, February 1990. CODEN CG-WODH. ISSN 0271-4159.

LeWinter:1990:BS

[LB90b] Renee LeWinter and Cynthia Baron. Bold Statements. Computer Graphics World, 13(8):74– ??, August 1990. CODEN CG-WODH. ISSN 0271-4159.

LeWinter:1991:AA

[LB91a] Renee LeWinter and Cynthia Baron. Artistic Animation. Computer Graphics World, 14(7):58– ??, July 1991. CODEN CG-WODH. ISSN 0271-4159.

LeWinter:1991:OOA

[LB91b] Renee LeWinter and Cynthia Baron. Output Options for Artists. Computer Graphics World, 14(1):34–??, January 1991. CODEN CGWODH. ISSN 0271-4159.

LeWinter:1991:PT

[LB91c] Renee LeWinter and Cynthia Baron. The Professional’s Touch. Computer Graphics World, 14(9): 30–??, September 1991. CODEN CGWODH. ISSN 0271-4159.

LeWinter:1993:OYD

[LeW93] Renee LeWinter. The Objects on Your Desire. Computer Graphics World, 16(8):54–??, August 1993. CODEN CGWODH. ISSN 0271-4159.

Liebman:1993:VN

[Lie93] Sheldon Liebman. Video on a Network. Computer Graphics World, 16(4):46–??, April 1993. CODEN CGWODH. ISSN 0271-4159.

Leland:1992:TWD

[LJ92] Jon Leland and Cliff Johnson.

The Tangled World of Desktop Video. Computer Graphics World, 15(10):46–??, October 1992. CO-DEN CGWODH. ISSN 0271-4159.

LoPiccolo:1991:VV

[LMVP91] Philip LoPiccolo, Diana Phillips Mahoney, Audrey Vasilopoulos, and Stephen Porter. The Visible Volume. Computer Graphics World, 14(4):44–??, April 1991. CODEN CGWODH. ISSN 0271-4159.

Long:1998:CB

[Lon98] Ben Long. Character building.

Computer Graphics World, 21(7): 49–??, July 1998. CODEN CG-WODH. ISSN 0271-4159.

Lang:1990:NTG

[LS90] Laura Lang and Vicki Speed.

A New Tool for GIS. Computer Graphics World, 13(10):40– ??, October 1990. CODEN CG-WODH. ISSN 0271-4159.

Lang:1992:EC

[LS92] Laura Lang and Vicki Speed. Environmental Consciousness. Computer Graphics World, 15(8):57– ??, August 1992. CODEN CG-WODH. ISSN 0271-4159.

Lyn:1995:SS

[Lyn95] Craig Lyn. Storage Strategies.

Computer Graphics World, 18(4): 35–??, April 1995. CODEN CG-WODH. ISSN 0271-4159.

MacNicol:1990:AAW

[Mac90a] Gregory MacNicol. 2D Animation: Alive and Well. Computer Graphics World, 13(3):40– ??, March 1990. CODEN CG-WODH. ISSN 0271-4159.

MacNicol:1990:FRA

[Mac90b] Gregory MacNicol. Film Recorders at the High End. Computer Graphics World, 13(8):102–??, August 1990. CODEN CG-WODH. ISSN 0271-4159.

MacNicol:1990:FRM

[Mac90c] Gregory MacNicol. Film Recorders: The Marketplace Myths. Computer Graphics World, 13(7):62– ??, July 1990. CODEN CG-WODH. ISSN 0271-4159.

MacNicol:1990:WWT

[Mac90d] Gregory MacNicol. Wild and Woolly Theme Park Rides. Computer Graphics World, 13(9):34– ??, September 1990. CODEN CG-WODH. ISSN 0271-4159.

MacNicol:1991:SA

[Mac91a] Gregory MacNicol. Spotlight on Animation. Computer Graphics World, 14(6):65–??, June 1991. CODEN CGWODH. ISSN 0271-4159.

MacNicol:1991:VMP

[Mac91b] Gregory MacNicol. Video Meets PC. Computer Graphics World, 14(2):55–??, February 1991. CO-DEN CGWODH. ISSN 0271-4159.

MacNicol:1992:VEH

[Mac92] Gregory MacNicol. Video Editing Hits the Desktop. Computer Graphics World, 15(4):32– ??, April 1992. CODEN CG-WODH. ISSN 0271-4159.

MacKrell:1993:MSR

[Mac93a] John MacKrell. Making Sense of a Revolution. Computer Graphics World, 16(11):26–??, November 1993. CODEN CGWODH. ISSN 0271-4159.

MacNicol:1993:AIE

[Mac93b] Gregory MacNicol. 3D Animation: Inexpensive and Effective. Computer Graphics World, 16(4): 37–??, April 1993. CODEN CG-WODH. ISSN 0271-4159.

MacNicol:1993:STL

[Mac93c] Gregory MacNicol. Sorting Through Large-Screen Monitors. Computer Graphics World, 16(8): 42–??, August 1993. CODEN CG-WODH. ISSN 0271-4159.

MacNicol:1994:ALL

[Mac94] G. MacNicol. The Art of Lighting:

The lighting of a scene is pivotal to the power of an image. Computer Graphics World, 17(12):41– ??, December 1994. CODEN CG-WODH. ISSN 0271-4159.

Macstri:1995:AMA

[Mac95] George Macstri. 3D Animation on the Mac… Animators creating complex effects should be critical about the feature sets of the software they choose. Thankfully, the Mac offers a wide range of 3D animation programs. Computer Graphics World, 18(7):44–??, July 1995. CODEN CGWODH. ISSN 0271-4159.

Maestri:1995:AM

[Mae95a] George Maestri. 3D animation on the mac. Computer Graphics World, 18(7):??, July 1995. CO-DEN CGWODH. ISSN 0271-4159.

Maestri:1995:CM

[Mae95b] George Maestri. Capturing motion. Computer Graphics World, 18(12):??, December 1995. CO-DEN CGWODH. ISSN 0271-4159.

Maestri:1995:CMM

[Mae95c] George Maestri. Capturing Motion — Motion-rapture systems are quickly entering the front lines of digital animation production, capturing motion for characters in everything from video games to feature films. But how do you go about implementing the technology? Computer Graphics World, 18(12):47–??, December 1995. CODEN CGWODH. ISSN 0271-4159.

Maestri:1996:WOG

[Mae96] George Maestri. A window of opportunity? given the dramatic promises being touted for the Windows NT operating system, the question for many animation houses is whether now is the right time to invest in NT machines. to help answer that question, we take a look at the first 3D animation programs that attempt to take advantage of NT: NewTek’s LightWave 3D, Microsoft’s Softimage 3D, and Autodesk’s 3D Studio MAX. Computer Graphics World, 19(4):32–??, April 1996.

CODEN CGWODH. ISSN 0271-4159.

Maestri:1997:AAU

[Mae97] George Maestri. Affordable animation. Computer Graphics World, 20(11):54–??, November 1997. CODEN CGWODH. ISSN 0271-4159.

Maguire:1991:IM

[Mag91] John G. Maguire. Interactive Magazines. Computer Graphics World, 14(8):33–??, August 1991. CODEN CGWODH. ISSN 0271-4159.

Maguire:1992:MYS

[Mag92] John G. Maguire. Movies on Your Screen. Computer Graphics World, 15(2):51–??, February 1992. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1990:AM

[Mah90] Diana Phillips Mahoney. The Art of Medicine. Computer Graphics World, 13(12):47–??, December 1990. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1991:IAR

[Mah91a] Diana Mahoney. It’s All Relative.

Computer Graphics World, 14(8): 42–??, August 1991. CODEN CG-WODH. ISSN 0271-4159.

Mahoney:1991:RM

[Mah91b] Diana Phillips Mahoney. Reading Minds. Computer Graphics World, 14(9):85–??, September 1991. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1991:RP

[Mah91c] Diana Phillips Mahoney. Rebuilding the Past. Computer Graphics World, 14(5):30–??, May 1991. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1991:SS

[Mah91d] Diana Phillips Mahoney. Seeing Stars. Computer Graphics World, 14(5):79–??, May 1991. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1992:CC

[Mah92a] Diana Phillips Mahoney. Campus Computing. Computer Graphics World, 15(11):53–??, November 1992. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1992:DHC

[Mah92b] Diana Phillips Mahoney. Designing for Health Care. Computer Graphics World, 15(2):39– ??, February 1992. CODEN CG-WODH. ISSN 0271-4159.

Mahoney:1992:LGB

[Mah92c] Diana Phillips Mahoney. Let the Games Begin. Computer Graphics World, 15(5):54–??, May 1992. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1992:P

[Mah92d] Diana Phillips Mahoney. Picture This. Computer Graphics World, 15(9):43–??, September 1992. CO-DEN CGWODH. ISSN 0271-4159.

Mahoney:1992:MRL

[Mah92e] Dianna Phillips Mahoney. Moviemaking in the Research Lab. Computer Graphics World, 15(12):19– ??, December 1992. CODEN CG-WODH. ISSN 0271-4159.

Mahoney:1993:AL

[Mah93a] Diana Phillips Mahoney. Artificial Life. Computer Graphics World, 16(2):37–??, February 1993. CO-DEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1993:DD

[Mah93b] Diana Phillips Mahoney. Drug Discovery. Computer Graphics World, 16(4):54–??, April 1993. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1993:MMU

[Mah93c] Diana Phillips Mahoney. Movies That Move Us. Computer Graphics World, 16(12):40–??, December 1993. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1993:SOC

[Mah93d] Diana Phillips Mahoney. Seeing Order in Chaos. Computer Graphics World, 16(7):51–??, July 1993. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1993:SL

[Mah93e] Diana Phillips Mahoney. The Stuff of Life. Computer Graphics World, 16(6):54–??, June 1993. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1993:TT

[Mah93f] Diana Phillips Mahoney. Time Travels. Computer Graphics World, 16(10):38–??, October 1993. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1994:VS

[Mah94a] D. P. Mahoney. Virtual Science. Computer Graphics World, 17(7): 20–??, July 1994. CODEN CG-WODH. ISSN 0271-4159.

Mahoney:1994:WI

[Mah94b] Diana Phillips Mahoney. And the winner is. Computer Graphics World, 17(1):40–??, January 1994. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1994:AAL

[Mah94c] Diana Phillips Mahoney. Avant-Garde Architects Look to CAD. Computer Graphics World, 17(3): 36–??, March 1994. CODEN CG-WODH. ISSN 0271-4159.

Mahoney:1994:C

[Mah94d] Diana Phillips Mahoney. Cityscapes. Computer Graphics World, 17(4): 36–??, April 1994. CODEN CG-WODH. ISSN 0271-4159.

Mahoney:1994:DND

[Mah94e] Diana Phillips Mahoney. Driving a New Design Mode. Computer Graphics World, 17(2):46– ??, February 1994. CODEN CG-WODH. ISSN 0271-4159.

Mahoney:1994:WTA

[Mah94f] Diana Phillips Mahoney. Walking Through Architectural Designs. Computer Graphics World, 17(6):22–??, June 1994. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1995:CT

[Mah95a] Diana Phillips Mahoney. CAD in toyland. Computer Graphics World, 18(4):??, April 1995. CO-DEN CGWODH. ISSN 0271-4159.

Mahoney:1995:DV

[Mah95b] Diana Phillips Mahoney. Driving VR. Computer Graphics World, 18(5):??, May 1995. CODEN CG-WODH. ISSN 0271-4159.

Mahoney:1995:FAA

[Mah95c] Diana Phillips Mahoney. Facial Animation: According to freelance animator Rod Altschul, speech involves not just lip movement, but also myriad facial muscles as well as those of the throat and tongue. Computer Graphics World, 18(1):60–62, January 1995. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1995:RPM

[Mah95d] Diana Phillips Mahoney. Rapid Prototyping in Medicine. Computer Graphics World, 18(2):42– ??, February 1995. CODEN CG-WODH. ISSN 0271-4159.

Mahoney:1996:VRP

[Mah96] Diana Phillips Mahoney. Virtual reality on the PC. Computer Graphics World, 19(5):??, May 1996. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1997:FTR

[Mah97] Diana Mahoney. Face tracking — recreating the subtle and notso-subtle expressions of the human face is no trivial task. even a simple smile involves the interplay of myriad muscles. yet given the booming market for animated characters, the push is on to create reliable face-tracking technology. Computer Graphics World, 20(4):23–??, April 1997. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1998:MCG

[Mah98a] Diana Philips Mahoney. Merging CAD and GIS. Computer Graphics World, 21(3):45–??, March 1998. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1998:B

[Mah98b] Diana Phillips Mahoney. Brainstorm. Computer Graphics World, 21(7):79–??, July 1998. CODEN CGWODH. ISSN 0271-4159.

Mahoney:1998:CST

[Mah98c] Diana Phillips Mahoney. Commercial success. Computer Graphics World, 21(4):65–??, April 1998. CODEN CGWODH. ISSN 0271-4159.

McCloskey:1996:LMD

[McC96] Bill McCloskey. Let’s make a deal — the secrets to negotiating a great deal on high-end animation software and hardware — and maintaining good post-sales and technical support — as revealed by a former software salesperson. Computer Graphics World, 19(2): 38–??, February 1996. CODEN CGWODH. ISSN 0271-4159.

McMillan:1990:DCP

[McM90a] Tom McMillan. Desktop Color Publishing. Computer Graphics World, 13(1):36–??, January 1990. CODEN CGWODH. ISSN 0271-

McMillan:1990:IVB

[McM90b] Tom McMillan. Interactive Video Bolsters Health Care Education. Computer Graphics World, 13(5): 44–??, May 1990. CODEN CG-WODH. ISSN 0271-4159.

McMillan:1990:ME

[McM90c] Tom McMillan. Multimedia and Education. Computer Graphics World, 13(10):68–??, October 1990. CODEN CGWODH. ISSN 0271-4159.

McMillan:1991:NTP

[McM91] Tom McMillan. New Tools for Package Design. Computer Graphics World, 14(8):62–??, August 1991. CODEN CGWODH. ISSN 0271-4159.

McMillan:1992:PPC

[McM92] Tom McMillan. The Promise of Portable Color. Computer Graphics World, 15(9):30–??, September 1992. CODEN CGWODH. ISSN 0271-4159.

Meyer:1997:VPS

[Mey97] Ann Meyer. Virtual prototypes.

Computer Graphics World, 20(9): 65–??, September 1997. CODEN CGWODH. ISSN 0271-4159.

Milburn:1993:FH

[Mil93] Ross Milburn. Flying High. Computer Graphics World, 16(2):44– ??, February 1993. CODEN CG-WODH. ISSN 0271-4159.

Milburn:1995:SSP

[Mil95] Ken Milburn. Shopping for scanners — part of the art of picking a scanner is understanding what you need it for and how that need matches up with the capabilities. Computer Graphics World, 18(7): 57–??, July 1995. CODEN CG-WODH. ISSN 0271-4159.

MintzTesta:1994:BSS

[Min94] Bridget Mintz Testa. Building a Space Station. Computer Graphics World, 17(5):24–??, May 1994. CODEN CGWODH. ISSN 0271-4159.

Morrison:1993:RA

[Mor93] Mike Morrison. Ready… and Action! Computer Graphics World, 16(11):49–??, November 1993. CODEN CGWODH. ISSN 0271-4159.

Novitski:1993:CC

[Nov93a] B. Novitski. Constructive Communication. Computer Graphics World, 16(6):35–??, June 1993. CODEN CGWODH. ISSN 0271-4159.

Novitski:1993:SWM

[Nov93b] B. Novitski. Students Who Mean Business. Computer Graphics World, 16(7):77–??, July 1993. CODEN CGWODH. ISSN 0271-4159.

Novitski:1993:VLC

[Nov93c] B. J. Novitski. Visiting Lost Cities. Computer Graphics World, 16(1):48–??, January 1993. CO-DEN CGWODH. ISSN 0271-4159.

Novitski:1993:SM

[Nov93d] Bobby-Jo Novitski. Sketch Modeling. Computer Graphics World, 16(5):45–??, May 1993. CODEN CGWODH. ISSN 0271-4159.

Novitski:1995:QVO

[Nov95] B. J. Novitski. QuickTime VR Opens Doors for Architects? With Apple’s QuickTime VR, architects — even those relying on relatively modest equipment — can “navigate” photographic-quality scenes. Computer Graphics World, 18 (11):50–??, November 1995. CO-DEN CGWODH. ISSN 0271-4159.

Novitski:1996:AAA

[Nov96] B. J. Novitski. Adventures in architecture — architects who find photorealism too expensive and time-consuming to achieve are experimenting with novel techniques for. Computer Graphics World, 19(6):23–??, June 1996. CODEN CGWODH. ISSN 0271-4159.

Novitski:1998:AAI

[Nov98a] B. J. Novitski. An architectural awakening. Computer Graphics World, 21(6):22–??, June 1998. CODEN CGWODH. ISSN 0271-4159.

Novitski:1998:DDR

[Nov98b] B. J. Novitski. Digital design reaches new heights. Computer Graphics World, 21(9):54– ??, September 1998. CODEN CG-WODH. ISSN 0271-4159.

Oliver:1994:EVE

[Oli94] D. Oliver. Enriched Video Editing: As hardware advances have enriched the quality of captured video, the tools used to edit video have improved as well. Computer Graphics World, 17(12):20– ??, December 1994. CODEN CG-WODH. ISSN 0271-4159.

Porter:1992:VR

[PE92] Stephen Porter and Arielle Emmett. Virtual Reality. Computer Graphics World, 15(3):42– ??, March 1992. CODEN CG-WODH. ISSN 0271-4159.

Porter:1990:SS

[PF90] Stephen Porter and Richard Fichera. Super Survey. Computer Graphics World, 13(8):90–??, August 1990. CODEN CGWODH. ISSN 0271-4159.

Pfitzer:1990:DME

[Pfi90a] Gary Pfitzer. Designing Made Easy. Computer Graphics World, 13(10):102–??, October 1990. CO-DEN CGWODH. ISSN 0271-4159.

Pfitzer:1990:SPD

[Pfi90b] Gary Pfitzer. Streamlining PCB Design. Computer Graphics World, 13(5):87–??, May 1990. CODEN CGWODH. ISSN 0271-4159.

Pfitzer:1991:MM

[Pfi91] Gary Pfitzer. Music and Motion. Computer Graphics World, 14(7): 68–??, July 1991. CODEN CG-WODH. ISSN 0271-4159.

Pfitzer:1992:ATM

[Pfi92a] Gary Pfitzer. Ancient Tale, Modern Impact. Computer Graphics World, 15(11):46–??, November 1992. CODEN CGWODH. ISSN 0271-4159.

Pfitzer:1992:CC

[Pfi92b] Gary Pfitzer. Computer Carpentry. Computer Graphics World, 15 (1):81–??, January 1992. CODEN CGWODH. ISSN 0271-4159.

Pfitzer:1992:IRF

[Pfi92c] Gary Pfitzer. Inside the Rain Forest. Computer Graphics World, 15(7):54–??, July 1992. CODEN CGWODH. ISSN 0271-4159.

Pfitzer:1992:LM

[Pfi92d] Gary Pfitzer. The Lawnmower Man. Computer Graphics World, 15(4):44–??, April 1992. CODEN CGWODH. ISSN 0271-4159.

Pfitzer:1993:YF

[Pfi93a] Gary Pfitzer. A Year in Film.

Computer Graphics World, 16 (12):52–??, December 1993. CO-DEN CGWODH. ISSN 0271-4159.

Pfitzer:1993:SA

[Pfi93b] Gary Pfitzer. State of the Arts.

Computer Graphics World, 16(3): 24–??, March 1993. CODEN CG-WODH. ISSN 0271-4159.

Pfitzer:1994:WR

[Pfi94a] G. Pfitzer. Wildebeests on the Run. Computer Graphics World, 17(7):52–??, July 1994. CODEN CGWODH. ISSN 0271-4159.

Pfitzer:1994:T

[Pfi94b] Gary Pfitzer. Thumbelina. Computer Graphics World, 17(4):31– ??, April 1994. CODEN CG-WODH. ISSN 0271-4159.

PhillipsMahoney:1995:AFI

[Phi95a] Diana Phillips Mahoney. Architecture for the Fun of It — The burgeoning digital entertainment market represents some exciting new opportunities for a growing number of architects. Computer Graphics World, 18(12):36– ??, December 1995. CODEN CG-WODH. ISSN 0271-4159.

PhillipsMahoney:1995:CT

[Phi95b] Diana Phillips Mahoney. CAD in Toyland. Computer Graphics World, 18(4):42–??, April 1995. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1995:DV

[Phi95c] Diana Phillips Mahoney. Driving VR. Computer Graphics World, 18(5):22–??, May 1995. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1995:MRT

[Phi95d] Diana Phillips Mahoney. Medical researchers take simulations to heart. Computer Graphics World, 18(9):47–49, September 1995. CO-DEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1995:MVR

[Phi95e] Diana Phillips Mahoney. Modeling for Virtual Reality — When developing virtual environments, modelers must follow a set of rules that are more complex than those adhered to by creators of passive animations. Computer Graphics World, 18(10):45–??, October 1995. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1995:UMS

[Phi95f] Diana Phillips Mahoney. Unlocking the mysteries of science — A variety of tools, including colors, textures, and animation, combine to help scientists better understand complex natural phenomena. Computer Graphics World, 18(7):22–??, July 1995. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1996:ANI

[Phi96a] Diana Phillips Mahoney. Animating nature — it’s not easy to mimic mother nature. yet that’s not stopping computer artists and animators from striving to create ever more believable digital versions of oceans, mountains, forests, hurricanes, and avalanches. Computer Graphics World, 19(12):40–??, December 1996. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1996:ASM

[Phi96b] Diana Phillips Mahoney. The art and science of medical visualization — in conjunction with dramatic advances in scanning technologies, computer-aided visualization tools are. Computer Graphics World, 19(7):25–??, July 1996. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1996:FAD

[Phi96c] Diana Phillips Mahoney. The fabric of animation — due to a variety of factors, realistically animating CG clothing is extraordinarily difficult. as such, most of the significant work being done in this area is taking place in universitybased research facilities. Computer Graphics World, 19(8):29– ??, August 1996. CODEN CG-WODH. ISSN 0271-4159.

PhillipsMahoney:1996:FPT

[Phi96d] Diana Phillips Mahoney. Feature prototyping — today, comfort and ease of use are top priorities when it comes to designing technology-rich products, such as avionics control. Computer Graphics World, 19(10):38–??, October 1996. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1996:LFC

[Phi96e] Diana Phillips Mahoney. L5: First city in space — this new Imax 3D film is a futuristic adventure that combines 3D computer graphics and live-action footage with actual space imagery-provided by NASA. Computer Graphics World, 19 (11):39–??, November 1996. CO-DEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1996:RWM

[Phi96f] Diana Phillips Mahoney. Riding the wave of motion-based movies — producers of ride-simulation films are increasingly looking to computer-generated imagery to create scenes that films are increasingly looking to computergenerated imagery to create scenes that are either technologically impossible or impractical to achieve. Computer Graphics World, 19(2): 30–??, February 1996. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1996:VRP

[Phi96g] Diana Phillips Mahoney. Virtual reality on the PC — significant enhancements in PC hardware are prompting many vendors of virtual-reality software to develop more user-friendly interfaces to enhance the appeal of their PCbased offerings. Computer Graphics World, 19(5):53–??, May 1996. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1997:AFA

[Phi97a] Diana Phillips Mahoney. All fired up — on average, it takes from three to five years to design and produce a new fire truck. but the engineers and designers at Freightliner managed to reduce that process to just nine months through some skillful use of its digital tools. Computer Graphics World, 20(5): 25–??, May 1997. CODEN CG-WODH. ISSN 0271-4159. PhillipsMahoney:1997:MBC

[Phi97b] Diana Phillips Mahoney. Moving beyond CAD. Computer Graphics World, 20(6):20–??, June 1997. CODEN CGWODH. ISSN 0271-4159. PhillipsMahoney:1997:PPA

[Phi97c] Diana Phillips Mahoney. Philadelphia 2000. Computer Graphics World, 20(6):30–??, June 1997. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1997:PTI

[Phi97d] Diana Phillips Mahoney. The power of touch. Computer Graphics World, 20(8):41–??, August 1997. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1997:PAP

[Phi97e] Diana Phillips Mahoney. Procedural animation — procedural animation techniques, which use algorithms or mathematical expressions to drive the movement of objects or volumes over time, already are delivering some significant benefits to animators. but the technology continues to advance, and tomorrow’s animation programs will make even greater use of these techniques. Computer Graphics World, 20(5):39–??, May 1997. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1997:VTN

[Phi97f] Diana Phillips Mahoney. Virtual therapy nets real results. Computer Graphics World, 20(12):52– ??, December 1997. CODEN CG-WODH. ISSN 0271-4159.

PhillipsMahoney:1997:VVU

[Phi97g] Diana Phillips Mahoney. Visualizing volumes. Computer Graphics World, 20(7):42–??, July 1997. CODEN CGWODH. ISSN 0271-4159.

PhillipsMahoney:1998:LCS

[Phi98] Diana Phillips Mahoney. Launching a construction simulation. Computer Graphics World, 21(8): 60–??, August 1998. CODEN CG-WODH. ISSN 0271-4159.

Pickover:1991:US

[Pic91] Clifford A. Pickover. The Ultimate Survey. Computer Graphics World, 14(11):57–??, November 1991. CODEN CGWODH. ISSN 0271-4159.

Pope:1994:EOB

[Pop94] D. Pope. Exploring the Options Beyond 3D Studio. Computer Graphics World, 17(11):44– ??, November 1994. CODEN CG-WODH. ISSN 0271-4159.

Porter:1990:FFC

[Por90a] Stephen Porter. FEA Faces a Changing Role. Computer Graphics World, 13(4):38–??, April 1990. CODEN CGWODH. ISSN 0271-4159.

Porter:1990:MS

[Por90b] Stephen Porter. Made for the Stage. Computer Graphics World, 13(8):60–??, August 1990. CO-DEN CGWODH. ISSN 0271-4159.

Porter:1991:DR

[Por91a] Stephen Porter. Designing by the Rules. Computer Graphics World, 14(11):28–??, November 1991. CODEN CGWODH. ISSN 0271-4159.

Porter:1991:EOE

[Por91b] Stephen Porter. Educating Our Engineers. Computer Graphics World, 14(10):46–??, October 1991. CODEN CGWODH. ISSN 0271-4159.

Porter:1991:WC

[Por91c] Stephen Porter. Winds of Change.

Computer Graphics World, 14(2): 34–??, February 1991. CODEN CGWODH. ISSN 0271-4159.

Porter:1992:IJL

[Por92] Stephen Porter. Interview: Jaron Lanier. Computer Graphics World, 15(4):61–??, April 1992. CODEN CGWODH. ISSN 0271-4159.

Porter:1994:EVN

[Por94] S. Porter. Engineering Visualization: The number of players in the visualization market is surprisingly large for a market that’s still relatively small. Computer Graphics World, 17(11):23–??, November 1994. CODEN CGWODH. ISSN 0271-4159.

Porter:1995:NHP

[Por95a] Stephen Porter. NT Hardware Power. Computer Graphics World, 18(8):66–??, August 1995. CO-DEN CGWODH. ISSN 0271-4159.

Porter:1995:WA

[Por95b] Stephen Porter. What’s Up With ACIS? Computer Graphics World, 18(9):30–??, September 1995. CO-DEN CGWODH. ISSN 0271-4159.

Potter:1991:SS

[Pot91a] Caren Potter. A Solid Solution? Computer Graphics World, 14(8): 51–??, August 1991. CODEN CG-WODH. ISSN 0271-4159.

Potter:1991:HF

[Pot91b] Caren Potter. The Human Factor. Computer Graphics World, 14(3): 61–??, March 1991. CODEN CG-WODH. ISSN 0271-4159.

Potter:1992:MAM

[Pot92a] Caren D. Potter. Mechanism Analysis Moves with the Times.

Computer Graphics World, 15(5): 30–??, May 1992. CODEN CG-WODH. ISSN 0271-4159.

Potter:1992:MME

[Pot92b] Caren D. Potter. Modeling Made Easy. Computer Graphics World, 15(11):28–??, November 1992. CODEN CGWODH. ISSN 0271-4159.

Potter:1993:FC

[Pot93a] Caren D. Potter. FEA to the Core.

Computer Graphics World, 16(5): 21–??, May 1993. CODEN CG-WODH. ISSN 0271-4159.

Potter:1993:RBE

[Pot93b] Caren D. Potter. Removing Barriers to Electro-Mechanical Design. Computer Graphics World, 16(2): 32–??, February 1993. CODEN CGWODH. ISSN 0271-4159.

Potter:1994:SFP

[Pot94] Caren D. Potter. A Select Few Push the Edge of Design. Computer Graphics World, 17(4):20– ??, April 1994. CODEN CG-WODH. ISSN 0271-4159.

Potter:1995:AEW

[Pot95a] Caren D. Potter. Animation for Engineers: While the use of animation by engineers is not new, a handful of vendors have recognized a growing demand for animation as a design review or engineering communications tool and have created products tailored specifically for these purposes. Computer Graphics World, 18(1):54–59, January 1995. CO-DEN CGWODH. ISSN 0271-4159.

Potter:1995:ACD

[Pot95b] Caren D. Potter. Architectural CAD on the Desktop. Computer Graphics World, 18(6):25– ??, June 1995. CODEN CG-WODH. ISSN 0271-4159.

Potter:1995:DTC

[Pot95c] Caren D. Potter. Digital tools for collaborative engineering. Computer Graphics World, 18(8):??, August 1995. CODEN CG-WODH. ISSN 0271-4159.

Potter:1995:SCCa

[Pot95d] Caren D. Potter. Special CAD/ CAM Supplement — Best Practices in Solid Modeling — What is it about solid modeling that makes it so difficult to do well, and how might it be better utilized? Computer Graphics World, 18(11):S1– ??, November 1995. CODEN CG-WODH. ISSN 0271-4159.

Potter:1995:SCCb

[Pot95e] Caren D. Potter. Special CAD/ CAM Supplement: Digital Tools for Collaborative Engineering. Computer Graphics World, 18(8): S1–??, August 1995. CODEN CG-WODH. ISSN 0271-4159.

Potter:1995:APP

[Pot95f] Caren D. Potter. The Allure of Pretty Pictures: As solid modeling becomes more widely used in engineering circles and rendering technology drops in price, the mechanical engineering community begins to see the value in using “pretty pictures” as a communications medium. Computer Graphics World, 18(3):26–??, March 1995. CODEN CGWODH. ISSN 0271-4159.

Potter:1996:AOD

[Pot96a] Caren D. Potter. Achieving the optimal design — because the design optimization capabilities in most finite-element analysis programs have traditionally been difficult to use for a number of reasons, designers and engineers haven’t really been making use of the tools. but as vendors enhance their design optimization offerings, previous objections to the use of this technology are beginning to disappear. Computer Graphics World, 19(5):21–??, May 1996. CODEN CGWODH. ISSN 0271-4159.

Potter:1996:CCSb

[Pot96b] Caren D. Potter. CAD/CAM special report: CAD contracts won and lost — why do companies buy the CAD systems they buy, and what factors might prompt users to replace an . . .. Computer Graphics World, 19(6):??, June 1996. CODEN CGWODH. ISSN 0271-4159.

Potter:1996:CCSa

[Pot96c] Caren D. Potter. CAD/CAM special report: New CAD architectures — as established CAD software vendors revise their product offerings, they’re devising underlying architectures that differ significantly from those of previous releases. one common ingredient in these newly architected CAD systems is the incorporation of object-oriented technology. Computer Graphics World, 19(3):S1– ??, March 1996. CODEN CG-WODH. ISSN 0271-4159.

Potter:1996:DLA

[Pot96d] Caren D. Potter. Designing large assemblies — pushed by the sophisticated demands of the aerospace and automotive industries, CAD/CAM technology is finally becoming adept at handling complex assemblies containing thousands and even tens of thousands of parts. Computer Graphics World, 19(11):S1– ??, November 1996. CODEN CG-WODH. ISSN 0271-4159.

Potter:1996:EIF

[Pot96e] Caren D. Potter. Engineering on the Internet — A few years ago, few CAD vendors were talking publicly about the Internet and its graphical showpiece, the World Wide Web. now, all of them have formulated Internet strategies, and a handful have already announced Internet-related products. Computer Graphics World, 19(8):51–??, August 1996. CO-DEN CGWODH. ISSN 0271-4159.

Potter:1996:NAB

[Pot96f] Caren D. Potter. New architectures: a blueprint for CAD’s future. Computer Graphics World, 19(3):??, March 1996. CODEN CGWODH. ISSN 0271-4159.

Potter:1996:STB

[Pot96g] Caren D. Potter. Sorting through the bargain bin of 3D CAD — looking for a 3D mechanical CAD program, but all you’ve got is USD 1500? take heart: a little money goes a long way in the world of low-cost mechanical CAD. Computer Graphics World, 19(1):23– ??, January 1996. CODEN CG-WODH. ISSN 0271-4159.

Potter:1997:ADE

[Pot97a] Caren Potter. Analysis for design engineers — the push is on to move analysis up to the front of the design process, and one way vendors are trying to accommodate that trend is by developing analysis tools that can be put directly into the hands of designers. but can such tools ever really be used safely by anyone other than specialized analysts? Computer Graphics World, 20(3):23– ??, March 1997. CODEN CG-WODH. ISSN 0271-4159.

Potter:1997:CAF

[Pot97b] Caren Potter. Can AutoCAD fight back? it’s been a tough year for Autodesk. its stock price has taken a beating, its last release of Auto-CAD software was soundly criticized, and its competitors are licking their chops. is this the beginning of a gloomy future, or is the once high-flying company poised to make a comeback? Computer Graphics World, 20(4):61– ??, April 1997. CODEN CG-WODH. ISSN 0271-4159.

Potter:1997:CML

[Pot97c] Caren Potter. Concept modelers: The latest trend in rapid prototyping — smaller, faster, cheaper siblings of conventional rapid-prototyping tools, concept modelers. Computer Graphics World, 20(12):45–??, December 1997. CODEN CGWODH. ISSN 0271-4159.

Potter:1997:EMC

[Pot97d] Caren Potter. The engines of mechanical CAD — as modeling kernels such as ACIS, Parasolid, and DesignBase continue their battle for market share, the debate rages. Computer Graphics World, 20(5): 31–??, May 1997. CODEN CG-WODH. ISSN 0271-4159.

Potter:1997:WEE

[Pot97e] Caren Potter. Web-enabled engineering. Computer Graphics World, 20(11):64–??, November 1997. CODEN CGWODH. ISSN 0271-4159.

Potter:1998:BMM

[Pot98a] Caren D. Potter. Battlinq for the midranqe. Computer Graphics World, 21(6):42–??, June 1998. CODEN CGWODH. ISSN 0271-4159.

Potter:1998:CMP

[Pot98b] Caren D. Potter. CAD makes a play in sports. Computer Graphics World, 21(5):24–??, May 1998. CODEN CGWODH. ISSN 0271-4159.

Pyros:1996:OAV

[Pyr96] Gregory Pyros. Outputting animation to video — there are a variety of methods available for outputting animation to video. but to decide which method is best for you and your application, you must understand several concepts related to the. Computer Graphics World, 19(6):41–??, June 1996. CODEN CGWODH. ISSN 0271-4159.

Pyros:1998:DVS

[Pyr98] Gregory Pyros. Designing virtual sets. Computer Graphics World, 21(6):63–??, June 1998. CODEN CGWODH. ISSN 0271-4159.

Reveaux:1994:IG

[Rev94] T. Reveaux. Interacting With the Gods. Computer Graphics World, 17(9):40–??, September 1994. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1990:BPG

[Rob90a] Barbara Robertson. Bridging the Photo-Retouching Gap. Computer Graphics World, 13(10):52– ??, October 1990. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1990:PTF

[Rob90b] Barbara Robertson. Photo-Retouching Technology: Friend or Foe? Computer Graphics World, 13(11):92–??, November 1990. CODEN CGWODH. ISSN 0271-4159.

Robertson:1990:SS

[Rob90c] Barbara Robertson. Sculpting the Scenery. Computer Graphics World, 13(6):48–??, June 1990. CODEN CGWODH. ISSN 0271-4159.

Robertson:1990:CFA

[Rob90d] Barbara Robertson. The Changing Face of AEC Presentations. Computer Graphics World, 13(3):

52–??, March 1990. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1990:TWH

[Rob90e] Barbara Robertson. The Tangled Web of HDTV. Computer Graphics World, 13(1):52–??, January 1990. CODEN CGWODH. ISSN 0271-4159.

Robinson:1990:S

[Rob90f] Phillip Robinson. Stereo 3D. Computer Graphics World, 13(6): 68–??, June 1990. CODEN CG-WODH. ISSN 0271-4159.

Robinson:1990:ESC

[Rob90g] Phillip Robinson. The Emergence of Smart CAD Software. Computer Graphics World, 13(7):55– ??, July 1990. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1991:MP

[Rob91a] Barbara Robertson. 3D Modeling for Print. Computer Graphics World, 14(11):36–??, November 1991. CODEN CGWODH. ISSN 0271-4159.

Robertson:1991:PR

[Rob91b] Barbara Robertson. A Painting Revolution. Computer Graphics World, 14(4):34–??, April 1991. CODEN CGWODH. ISSN 0271-4159.

Robertson:1991:BV

[Rob91c] Barbara Robertson. Biz Viz. Computer Graphics World, 14(9): 45–??, September 1991. CODEN CGWODH. ISSN 0271-4159.

Robertson:1991:P

[Rob91d] Barbara Robertson. Painting in 3D. Computer Graphics World, 14(10):38–??, October 1991. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1991:ET

[Rob91e] Barbara Robertson. The Electronic Theatre. Computer Graphics World, 14(7):46–??, July 1991. CODEN CGWODH. ISSN 0271-4159.

Robertson:1992:AI

[Rob92a] Barbara Robertson. An Art Unto Itself. Computer Graphics World, 15(12):28–??, December 1992. CODEN CGWODH. ISSN 0271-4159.

Robertson:1992:D

[Rob92b] Barbara Robertson. Disc-O-Mania. Computer Graphics World, 15(6):40–??, June 1992. CODEN CGWODH. ISSN 0271-4159.

Robertson:1992:MP

[Rob92c] Barbara Robertson. Moving Pictures. Computer Graphics World, 15(10):38–??, October 1992. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1992:RR

[Rob92d] Barbara Robertson. Roll Over Rodin. Computer Graphics World, 15(3):33–??, March 1992. CODEN CGWODH. ISSN 0271-4159.

Robertson:1992:SM

[Rob92e] Barbara Robertson. Shades of Magic. Computer Graphics World, 15(9):59–??, September 1992. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1992:AFM

[Rob92f] Barbara Robertson. Those Amazing Flying Machines. Computer Graphics World, 15(5):67–??, May 1992. CODEN CGWODH. ISSN 0271-4159.

Robertson:1993:BG

[Rob93a] Barbara Robertson. Board Games. Computer Graphics World, 16(5):39–??, May 1993. CODEN CGWODH. ISSN 0271-4159.

Robertson:1993:DM

[Rob93b] Barbara Robertson. Dinosaur Magic. Computer Graphics World, 16(9):44–??, September 1993. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1993:EM

[Rob93c] Barbara Robertson. Easy Motion. Computer Graphics World, 16 (12):33–??, December 1993. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1993:PG

[Rob93d] Barbara Robertson. Physical Graphics. Computer Graphics World, 16(3):37–??, March 1993. CODEN CGWODH. ISSN 0271-4159.

Robertson:1993:PP

[Rob93e] Barbara Robertson. Powerful Particles. Computer Graphics World, 16(7):40–??, July 1993. CODEN CGWODH. ISSN 0271-4159.

Robertson:1993:PPG

[Rob93f] Barbara Robertson. Prime-Time Proving Ground for 3D Graphics. Computer Graphics World, 16(1):

35–??, January 1993. CODEN CGWODH. ISSN 0271-4159.

Robertson:1993:R

[Rob93g] Barbara Robertson. Rendering in the ’90s. Computer Graphics World, 16(10):28–??, October 1993. CODEN CGWODH. ISSN 0271-4159.

Robertson:1994:CAA

[Rob94a] B. Robertson. Caught in the Act: The ability to capture motion from live beings is finding wide acceptance as a pragmatic and artful solution to many types of animation problems. Computer Graphics World, 17(9):23– ??, September 1994. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1994:DLC

[Rob94b] B. Robertson. Disney lets CAPS out of the bag. Computer Graphics World, 17(7):58–??, July 1994. CODEN CGWODH. ISSN 0271-4159.

Robertson:1994:FPL

[Rob94c] B. Robertson. Fresh Paint: Like word processing, page-layout, and image-editing programs, 3D painting software-which enables artists to paint texture maps directly and interactively onto 3D modelsis one of those technologies that seems impossible to imagine ever having been without. Computer Graphics World, 17(12):28– ??, December 1994. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1994:KST

[Rob94d] B. Robertson. Kid’s Stuff: Today, multimedia technology has become sophisticated enough to support some rather interesting children’s software designed to teach as it entertains. Computer Graphics World, 17(10):32–??, October 1994. CODEN CGWODH. ISSN 0271-4159.

Robertson:1994:SCE

[Rob94e] B. Robertson. Studios on the Cutting Edge. Computer Graphics World, 17(7):29–??, July 1994. CODEN CGWODH. ISSN 0271-4159.

Robertson:1994:URN

[Rob94f] B. Robertson. Unbelievably real: For the new feature film The Mask, the team of award-winning special-effects artists at Industrial Light and Magic got their first chance to work with a comedy. Computer Graphics World, 17(8): 52–??, August 1994. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1994:CP

[Rob94g] Barbara Robertson. 3D Characters on Parade. Computer Graphics World, 17(3):28–??, March 1994. CODEN CGWODH. ISSN 0271-4159.

Robertson:1994:DT

[Rob94h] Barbara Robertson. Digital Toons. Computer Graphics World, 17(6):40–??, June 1994. CODEN CGWODH. ISSN 0271-4159.

Robertson:1994:FF

[Rob94i] Barbara Robertson. Fascinating filters. Computer Graphics World, 17(1):33–??, January 1994. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1994:RWR

[Rob94j] Barbara Robertson. Ready for a wild ride? Computer Graphics World, 17(2):30–??, February 1994. CODEN CGWODH. ISSN 0271-4159.

Robertson:1994:MAD

[Rob94k] Barbara Robertson. The Magical Art of Digital Matte Painting. Computer Graphics World, 17(5): 14–??, May 1994. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1995:BSD

[Rob95a] Barbara Robertson. Behind the scenes of disclosure: How IL and M created the virtual-reality environment for Warner Bros.’ hot hit Disclosure. Computer Graphics World, 18(2):35–38, February 1995. CODEN CGWODH. ISSN 0271-4159.

Robertson:1995:DAI

[Rob95b] Barbara Robertson. Direct Action — By improving on its motioncapture system Boss Film created an animation tool that enabled the director of the new sci-fi adventure film species to work in real time with a textured CG creature that was moving within an environment. Computer Graphics World, 18(7):52–??, July 1995. CODEN CGWODH. ISSN 0271-4159.

Robertson:1995:HET

[Rob95c] Barbara Robertson. Hair-Raising Effects — Techniques for putting hair on humans and fur on ani als are moving out of custom effects houses and into commercial software. Computer Graphics World, 18(10):24–??, October 1995. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1995:PES

[Rob95d] Barbara Robertson. Plug-Ins Electrify 3D Studio. Computer Graphics World, 18(5):36–??, May 1995. CODEN CGWODH. ISSN 0271-4159.

Robertson:1995:MBC

[Rob95e] Barbara Robertson. The Magic Behind Casper. Computer Graphics World, 18(6):38–??, June 1995. CODEN CGWODH. ISSN 0271-4159.

Robertson:1995:TST

[Rob95f] Barbara Robertson. Toy Story: A Triumph of Animation. Computer Graphics World, 18(8):28–??, August 1995. CODEN CGWODH. ISSN 0271-4159.

Robertson:1995:UA

[Rob95g] Barbara Robertson. Untamed animation. Computer Graphics World, 18(4):24–??, April 1995. CODEN CGWODH. ISSN 0271-4159.

Robertson:1995:WSV

[Rob95h] Barbara Robertson. What a set-up!. . .. virtual sets — 3D computer graphics environments synced to real cameras, rendered in real time, and populated with real people and CG characters who interact with each other — are generating excitement in the broadcast world. Computer Graphics World, 18(7):36–??, July 1995. CODEN CGWODH. ISSN 0271-4159.

Robertson:1996:BDV

[Rob96a] Barbara Robertson. Battling for dominance — vendors claiming their workstations are “the priceperformance leaders” or “the honest 3D machines” battle for dominance in the midrange of the market. Computer Graphics World, 19 (3):52–??, March 1996. CODEN CGWODH. ISSN 0271-4159.

Robertson:1996:HBT

[Rob96b] Barbara Robertson. He’s back!

T2 3D, the third installment of the fast-action, science-fiction Terminator saga, combines elements from the theater, virtualreality, theme-park, and feature film worlds to create an immersive, sensory experience. Computer Graphics World, 19(7):36– ??, July 1996. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1996:JAA

[Rob96c] Barbara Robertson. Jumanji’s amazing animals — ILM’s visualeffects team has done it again. the creators of the believable dinosaurs in Jurassic Park have performed their magic once again for Jumanji, TriStar Pictures’ new action-adventure film, which stars hundreds of 3D, animated animals you’d swear are living, breathing creatures. Computer Graphics World, 19(1):28–??, January 1996. CODEN CGWODH. ISSN 0271-4159.

Robertson:1996:TTE

[Rob96d] Barbara Robertson. Texture tactics — even though the quality differs from product to product, hardware acceleration for texture mapping is showing up on everything from . . .. Computer Graphics World, 19(6):33–??, June 1996. CODEN CGWODH. ISSN 0271-4159.

Robertson:1997:PVC

[Rob97a] Barbara Robertson. 3D power. Computer Graphics World, 20 (10):65–??, October 1997. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1997:ASR

[Rob97b] Barbara Robertson. Animation special report: Making motion easy — computer animation programs have not always been easy to use. but that’s starting to change as vendors bring out an ever-increasing number of clever tools that make the art of animation a little faster, a little easier, and maybe even a little better. Computer Graphics World, 20(4): 42–??, April 1997. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1997:DSC

[Rob97c] Barbara Robertson. Different strokes. Computer Graphics World, 20(12):26–??, December 1997. CODEN CGWODH. ISSN 0271-4159.

Robertson:1997:PPP

[Rob97d] Barbara Robertson. Pioneers and performers. Computer Graphics World, 20(11):46–??, November 1997. CODEN CGWODH. ISSN 0271-4159.

Robertson:1997:PE

[Rob97e] Barbara Robertson. Pushing the envelope. Computer Graphics World, 20(8):81–??, August 1997. CODEN CGWODH. ISSN 0271-4159.

Robertson:1997:RML

[Rob97f] Barbara Robertson. Read my lips.

Computer Graphics World, 20(8): 26–??, August 1997. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1997:RTS

[Rob97g] Barbara Robertson. Real troopers. Computer Graphics World, 20 (9):26–??, September 1997. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1997:SSI

[Rob97h] Barbara Robertson. Something has survived. Computer Graphics World, 20(6):34–??, June 1997. CODEN CGWODH. ISSN 0271-4159.

Robertson:1997:VHA

[Rob97i] Barbara Robertson. Virtual humans at work. Computer Graphics World, 20(7):33–??, July 1997. CODEN CGWODH. ISSN 0271-4159.

Robertson:1998:APS

[Rob98a] Barbara Robertson. Antzpiration. Computer Graphics World, 21(10):34–??, October 1998. CODEN CGWODH. ISSN 0271-4159.

Robertson:1998:BPP

[Rob98b] Barbara Robertson. Brushing up. Computer Graphics World, 21(4): 26–??, April 1998. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1998:CCN

[Rob98c] Barbara Robertson. Crowd control. Computer Graphics World, 21(2):30–??, February 1998. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1998:GIO

[Rob98d] Barbara Robertson. The grand illusion. Computer Graphics World, 21(1):22–??, January 1998. CO-DEN CGWODH. ISSN 0271-4159.

Robertson:1998:MGN

[Rob98e] Barbara Robertson. Meet geri: The new face of animation. Computer Graphics World, 21(2):20– ??, February 1998. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1998:NCC

[Rob98f] Barbara Robertson. A new cast of characters: From Antz to ’Zillas. Computer Graphics World, 21(7): 31–??, July 1998. CODEN CG-WODH. ISSN 0271-4159.

Robertson:1998:RTC

[Rob98g] Barbara Robertson. Right on track. Computer Graphics World, 21(5):54–??, May 1998. CODEN CGWODH. ISSN 0271-4159.

Robertson:1998:SAN

[Rob98h] Barbara Robertson. Sav A-h-h-h-h. Computer Graphics World, 21 (3):28–??, March 1998. CODEN CGWODH. ISSN 0271-4159.

Romwell:1997:DBG

[Rom97] Amy Romwell. The design benefits of group VR — what’s the best way for engineers and scientists to share their ideas with one another? for some, the best solution might be to use virtual reality. from VR theaters to shared virtual space via the Internet, virtual reality is becoming an increasingly important medium in project design, presentation, and collaboration. Computer Graphics World, 20(2):21–??, February 1997. CO-DEN CGWODH. ISSN 0271-4159.

Rosenberg:1990:WAP

[Ros90] Debra Rosenberg. What’s Ahead for PC CAD? Computer Graphics World, 13(4):63–??, April 1990. CODEN CGWODH. ISSN 0271-4159.

Rowell:1996:COD

[Row96] Amy Rowell. The 2D CAD option — despite the growing popularity of solid modeling software, the demand for ZD CAD software continues to be strong. and whether you’re a technical illustrator, an engineer, an architect, or a hobbyist, chances are good that you can find a package perfectly suited to your particular needs. Computer Graphics World, 19(7):52–??, July 1996. CODEN CGWODH. ISSN 0271-4159.

Rowe:1997:SMW

[Row97a] Jeffrey Rowe. Surface modeling.

Computer Graphics World, 20(9): 47–??, September 1997. CODEN CGWODH. ISSN 0271-4159.

Rowell:1997:CCI

[Row97b] Amy Rowell. CAD/CAM interoperability. Computer Graphics World, 20(6):57–??, June 1997. CODEN CGWODH. ISSN 0271-4159.

Rowell:1997:ELR

[Row97c] Amy Rowell. Engineered to look real. Computer Graphics World, 20(11):26–??, November 1997. CODEN CGWODH. ISSN 0271-4159.

Rowell:1997:PDW

[Row97d] Amy Rowell. Prototyping in a digital world. Computer Graphics World, 20(9):55–??, September 1997. CODEN CGWODH. ISSN 0271-4159.

Rowell:1997:CMH

[Row97e] Amy A. Rowell. CAD meets hollywood at TFX. Computer Graphics World, 20(8):53–??, August 1997. CODEN CGWODH. ISSN 0271-4159.

Rowe:1998:TDS

[Row98a] Jeffrey Rowe. Tools for the digital stylist. Computer Graphics World, 21(3):73–??, March 1998. CODEN CGWODH. ISSN 0271-4159.

Rowell:1998:VVS

[Row98b] Amy A. Rowell. Virtual vehicles set the pace. Computer Graphics World, 21(3):61–??, March 1998. CODEN CGWODH. ISSN 0271-4159.

Rowell:1998:PDM

[Row98c] Arny A. Rowell. Product data management sports a new look.

Computer Graphics World, 21(2): 42–??, February 1998. CODEN CGWODH. ISSN 0271-4159.

Rowell:1997:CCS

[RR97] Amy Rowell and Jeffrey Rowe.

CAD/CAM special report: Industrial design — industrial design is a complex field of activity that straddles the worlds of both art and engineering. the job of the industrial designer is to create innovative products and help bring them to market quickly. Computer Graphics World, 20(3):S1– ??, March 1997. CODEN CG-WODH. ISSN 0271-4159.

Rosenberg:1990:SP

[RV90] Debra Rosenberg and Audrey Vasilopoulos. Smart Prototypes. Computer Graphics World, 13(6): 99–??, June 1990. CODEN CG-WODH. ISSN 0271-4159.

Savage:1998:AIT

[Sav98] Lorraine Savage. The anime invasion. Computer Graphics World, 21(4):73–??, April 1998. CODEN CGWODH. ISSN 0271-4159.

Schultz:1992:HMM

[Sch92] Steven Schultz. High Marks for Multimedia. Computer Graphics World, 15(7):61–??, July 1992. CODEN CGWODH. ISSN 0271-4159.

Schechter:1995:AG

[Sch95a] Joanne Schechter. Accelerating 3D graphics. Computer Graphics World, 18(10):??, October 1995. CODEN CGWODH. ISSN 0271-

Schechter:1995:AGG

[Sch95b] Joanne Schechter. Accelerating 3D Graphics — 3D graphics accelerators are promising to bring workstation-quality 3D graphics to Pentium PCs. But can you really get workstation-level 3D graphics on the PC at PC prices? Computer Graphics World, 18 (10):35–??, October 1995. CO-DEN CGWODH. ISSN 0271-4159.

Schechter:1995:SCD

[Sch95c] Joanne Schechter. The Spin on CD-ROM Drives. Computer Graphics World, 18(9):23– ??, September 1995. CODEN CG-WODH. ISSN 0271-4159.

Schechter:1997:FMO

[Sch97a] Joanne Schechter. Flat monitors. Computer Graphics World, 20(7): 53–??, July 1997. CODEN CG-WODH. ISSN 0271-4159.

Schechter:1997:SPL

[Sch97b] Joanne Schechter. Selecting a projector. Computer Graphics World, 20(6):51–??, June 1997. CODEN CGWODH. ISSN 0271-4159.

Schechter:1997:UWS

[Sch97c] Joanne Schechter. Unix workstations stand their ground. Computer Graphics World, 20(10):36– ??, October 1997. CODEN CG-WODH. ISSN 0271-4159.

Schechter:1997:UYT

[Sch97d] Joanne Schechter. Update your thinking of stereoscopic displays. Computer Graphics World, 20(9): 37–??, September 1997. CODEN CGWODH. ISSN 0271-4159.

Schechter:1997:WED

[Sch97e] Joanne Schechter. The workstation explosion — dramatic advances in PC and graphics technology have effectively married the full-featured graphics workstation to the ordinary PC. the result? the powerful NT workstation. as the market for these new machines explodes, vendors scramble to claim their slice of the pie. Computer Graphics World, 20(2): 40–??, February 1997. CODEN CGWODH. ISSN 0271-4159.

Schechter:1998:DMT

[Sch98a] Joanne Schechter. 3D digitizers make their mark. Computer Graphics World, 21(3):79– ??, March 1998. CODEN CG-WODH. ISSN 0271-4159.

Schechter:1998:RRF

[Sch98b] Joanne Schechter. Richer ride films. Computer Graphics World, 21(1):44–??, January 1998. CO-DEN CGWODH. ISSN 0271-4159.

Sena:1990:CD

[Sen90a] Michael Sena. Computer-Aided Dispatching. Computer Graphics World, 13(5):34–??, May 1990. CODEN CGWODH. ISSN 0271-4159.

Sena:1990:DM

[Sen90b] Michael L. Sena. Do-it-Yourself Mapping. Computer Graphics World, 13(3):62–??, March 1990. CODEN CGWODH. ISSN 0271-4159.

Sculley:1990:PA

[SH90] Jack Sculley and Inge Hutzel. Playing Author. Computer Graphics World, 13(2):58–??, February 1990. CODEN CGWODH. ISSN 0271-4159.

ShammsMortier:1995:SCA

[Sha95] R. Shamms Mortier. A Second Chance for the Amiga — How long a life the Amiga can sustain depends on the direction in which German computer manufacturer ESCOM AG takes the technology, how quickly it can bring new products to market, and how users and developers respond. Computer Graphics World, 18(10):52– ??, October 1995. CODEN CG-WODH. ISSN 0271-4159.

Shoor:1990:PP

[Sho90a] Rita Shoor. Pouring Plastic. Computer Graphics World, 13(12):60– ??, December 1990. CODEN CG-WODH. ISSN 0271-4159.

Shoor:1990:TT

[Sho90b] Rita Shoor. Targeting Treatment. Computer Graphics World, 13(2): 38–??, February 1990. CODEN CGWODH. ISSN 0271-4159.

Shoor:1991:HS

[Sho91] Rita Shoor. Hollywood Style. Computer Graphics World, 14(3): 34–??, March 1991. CODEN CG-WODH. ISSN 0271-4159.

Shoor:1993:EOO

[Sho93a] Rita Shoor. The Erasable Optical Option. Computer Graphics World, 16(1):59–??, January 1993.

CODEN CGWODH. ISSN 0271-4159.

Shoor:1993:TD

[Sho93b] Rita Shoor. Trapping on the Desktop. Computer Graphics World, 16(6):67–??, June 1993. CODEN CGWODH. ISSN 0271-4159.

Simerman:1995:AAT

[Sim95a] T. Simerman. Anatomy of an Animation: Technical hurdles, time constraints, and high costs are often cited as obstacles to the generation of more sophisticated 3D characters. Computer Graphics World, 18(3):36–??, March 1995. CODEN CGWODH. ISSN 0271-4159.

Simerman:1995:AA

[Sim95b] Tony Simerman. Anatomy of an animation. Computer Graphics World, 18(3):??, March 1995. CO-DEN CGWODH. ISSN 0271-4159.

Smith:1996:CCS

[Smi96a] Susan Smith. The challenges of complex surfaces — while some argue that surfaces have not progressed nearly enough, others point to developments in surface modeling that make working with complex surfaces just a little easier. Computer Graphics World, 19 (11):26–??, November 1996. CO-DEN CGWODH. ISSN 0271-4159.

Smith:1996:HMI

[Smi96b] Susan Smith. Hybrid modeling — if this new approach to solid modeling does everything vendors claim it does, then it’s no wonder that some companies are touting it as the wave of the future.

Computer Graphics World, 19(9): 25–??, September 1996. CODEN CGWODH. ISSN 0271-4159.

Sorensen:1990:AAA

[Sor90a] Peter Sorensen. Animation Across the Airwaves. Computer Graphics World, 13(11):80–??, November 1990. CODEN CGWODH. ISSN 0271-4159.

Sorensen:1990:CSS

[Sor90b] Peter Sorensen. From the Computer Screen to the Silver Screen. Computer Graphics World, 13(9): 57–??, September 1990. CODEN CGWODH. ISSN 0271-4159.

Sorensen:1990:HG

[Sor90c] Peter Sorensen. Hair-Raising Graphics. Computer Graphics World, 13(1):44–??, January 1990. CODEN CGWODH. ISSN 0271-4159.

Sorensen:1990:FF

[Sor90d] Peter Sorensen. The Fruits of Fractals. Computer Graphics World, 13(5):54–??, May 1990. CODEN CGWODH. ISSN 0271-4159.

Sorensen:1991:AA

[Sor91a] Peter Sorensen. Algorithmic Advancements. Computer Graphics World, 14(6):42–??, June 1991. CODEN CGWODH. ISSN 0271-4159.

Sorensen:1991:EC

[Sor91b] Peter Sorensen. Exploring the Cosmos. Computer Graphics World, 14(4):86–??, April 1991. CODEN CGWODH. ISSN 0271-4159.

Sorensen:1991:MW

[Sor91c] Peter Sorensen. Making Waves.

Computer Graphics World, 14(3): 40–??, March 1991. CODEN CG-WODH. ISSN 0271-4159.

Sorensen:1991:TFE

[Sor91d] Peter Sorensen. Terminator 2: A Film Effects Revolution. Computer Graphics World, 14(10):56– ??, October 1991. CODEN CG-WODH. ISSN 0271-4159.

Sorensen:1991:WWE

[Sor91e] Peter Sorensen. Water, Water, Everywhere. Computer Graphics World, 14(3):48–??, March 1991. CODEN CGWODH. ISSN 0271-4159.

Sorensen:1992:MM

[Sor92] Peter Sorensen. Morphing Magic.

Computer Graphics World, 15(1): 36–??, January 1992. CODEN CGWODH. ISSN 0271-4159.

Spohrer:1998:MCM

[Spo98] Richlard Spohrer. Making CAD models shine. Computer Graphics World, 21(1):51–??, January 1998. CODEN CGWODH. ISSN 0271-4159.

Sprince:1994:MIP

[Spr94] R. Sprince. The Making of an Interactive Press Kit: A behindthe-scenes look at how ABC Television used multimedia technology to create an interactive, CD-ROM-based press kit to promote a television miniseries. Computer Graphics World, 17(9):49– ??, September 1994. CODEN CG-WODH. ISSN 0271-4159.

Stock:1990:NW

[SR90] Rodney Stock and Barbara Robertson. The Next Wave. Computer Graphics World, 13(8):82– ??, August 1990. CODEN CG-WODH. ISSN 0271-4159.

Stapleton:1990:VM

[Sta90] Lisa Stapleton. Visualizing Math. Computer Graphics World, 13(6): 59–60, 63–64, June 1990. CODEN CGWODH. ISSN 0271-4159.

Stock:1990:SC

[Sto90] Rodney Stock. The Super Chip. Computer Graphics World, 13(7): 85–??, July 1990. CODEN CG-WODH. ISSN 0271-4159.

Tests:1994:GTI

[Tes94] B. M. Tests. Graphical Treasures on the Internet. Computer Graphics World, 17(11):34–??, November 1994. CODEN CGWODH. ISSN 0271-4159.

Thornton:1995:MM

[Tho95a] Jack Thornton. From model to mold. Computer Graphics World, 18(11):??, November 1995. CO-DEN CGWODH. ISSN 0271-4159.

Thornton:1995:MMD

[Tho95b] Jack Thornton. From Model to Mold — Developments in computer-aided manufacturing software are helping to speed up the design-to-manufacturing process for plastic parts. Computer Graphics World, 18(11):26– ??, November 1995. CODEN CG-WODH. ISSN 0271-4159.

Trueblood:1986:HH

[Tru86] Mark Trueblood. Highlighting Halley’s. Computer Graphics World, 9(9):44–46, 48, September 1986. CODEN CGWODH. ISSN 0271-4159.

Tucker:1995:WWE

[Tuc95a] M. J. Tucker. The Workstation Warriors: Even for people who have been using workstations for years and who are familiar with the issues of workstation/server computing, deciding what to buy today can be a major headache. Computer Graphics World, 18(1): 25–??, January 1995. CODEN CGWODH. ISSN 0271-4159.

Tucker:1995:WW

[Tuc95b] Michael Jay Tucker. Workstation warriors. Computer Graphics World, 18(1):??, January 1995. CODEN CGWODH. ISSN 0271-4159.

Tucker:1996:HMG

[Tuc96] Michael Jay Tucker. Have highend monitors gone vanilla? if you’re interested in purchasing a 1- to 21-inch, high-end color monitor, your choices are broader than ever. Computer Graphics World, 19(8):65–??, August 1996. CO-DEN CGWODH. ISSN 0271-4159.

Vacca:1996:WWT

[Vac96] John R. Vacca. 3D worlds on the Web — thanks to the stabilization of the Virtual Reality Modeling Language (VRML), the tools needed to create and view 3D worlds on the Web are becoming more widely available. Computer Graphics World, 19(5):43–??, May 1996. CODEN CGWODH. ISSN 0271-4159.

Vasilopoulos:1990:AGG

[Vas90a] Audrey Vasilopoulos. Atlanta Gets the Gold. Computer Graphics World, 13(12):83–??, December 1990. CODEN CGWODH. ISSN 0271-4159.

Vasilopoulos:1990:CA

[Vas90b] Audrey Vasilopoulos. Cartoon Animation. Computer Graphics World, 13(1):77–??, January 1990. CODEN CGWODH. ISSN 0271-4159.

Vasilopoulos:1990:GT

[Vas90c] Audrey Vasilopoulos. Graphics on Trial. Computer Graphics World, 13(2):83–??, February 1990. CO-DEN CGWODH. ISSN 0271-4159.

Vasilopoulos:1990:SC

[Vas90d] Audrey Vasilopoulos. Shopping by Computer. Computer Graphics World, 13(4):85–??, April 1990. CODEN CGWODH. ISSN 0271-4159.

Vasilopoulos:1991:MA

[Vas91] Audrey Vasilopoulos. Mix-and-Match Animation. Computer Graphics World, 14(1):85–??, January 1991. CODEN CGWODH. ISSN 0271-4159.

Vincenzi:1997:BMO

[Vin97] Lisa Vincenzi. Buying motion — one way to create animated characters is to use motion-capture technology. and if you don’t have the tools or the resources to do it yourself, there’s a growing number of motion-capture service bureaus who’ll gladly step in to help you out. Computer Graphics World, 20(4):42–??, April 1997. CODEN CGWODH. ISSN 0271-4159.

Warniers:1998:DPK

[War98a] Randall Warniers. Dirty pictures.

Computer Graphics World, 21(6): 50–??, June 1998. CODEN CG-WODH. ISSN 0271-4159.

Warniers:1998:EPT

[War98b] Randall Warniers. Every picture tells a story. Computer Graphics World, 21(10):25–??, October 1998. CODEN CGWODH. ISSN 0271-4159.

Winiarski:1990:XTD

[Win90] Kathryn Winiarski. X Terminal Displays. Computer Graphics World, 13(3):85–??, March 1990. CODEN CGWODH. ISSN 0271-4159.

Wohlers:1990:PP

[Woh90] Terry T. Wohlers. Practical Prototypes. Computer Graphics World, 13(3):73–??, March 1990. CODEN CGWODH. ISSN 0271-4159.

Wohlers:1992:D

[Woh92] Terry Wohlers. 3D Digitizers.

Computer Graphics World, 15(7): 73–??, July 1992. CODEN CG-WODH. ISSN 0271-4159.

Wohlers:1997:TBD

[Woh97] Terry Wohlers. The technology behind 3D digitizing — one of the most common misconceptions is that 3D digitizers somehow magically covert an object to computer model data. in truth, there’s nothing magical about it. making 3D digitizing work is as much an art as it is a science. Computer Graphics World, 20(3):47– ??, March 1997. CODEN CG-WODH. ISSN 0271-4159.

Wolfe:1993:FS

[Wol93] L. Stephen Wolfe. Freeform Surfaces. Computer Graphics World, 16(7):59–??, July 1993. CODEN CGWODH. ISSN 0271-4159.

Wood:1993:BGM

[Woo93] Elizabeth Wood. A Beginner’s Guide to Multimedia. Computer Graphics World, 16(10):71– ??, October 1993. CODEN CG-WODH. ISSN 0271-4159.

Yares:1995:PYC

[Yar95]  Evan Yares.  Planning Your

Choice of Inkjet Plotters — Inkjet plotters not only dominate the CAD output market, they’re also gaining popularity with graphic artists. What makes them so popular? And what should you consider when purchasing such a device? Computer Graphics World, 18(11):41–??, November 1995. CODEN CGWODH. ISSN 0271-4159.

Yares:1997:PPC

[Yar97] Evan Yares. Photo-quality printers. Computer Graphics World, 20(10):50–??, October 1997. CO-DEN CGWODH. ISSN 0271-4159.